### v2

# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

Abstract
Vision Transformers (ViTs) achieve superior performance on various tasks compared to convolutional neural networks (CNNs), but ViTs are also vulnerable to adversarial attacks. Adversarial training is one of the most successful methods to build robust CNN models. Thus, recent works explored new methodologies for adversarial training of ViTs based on the differences between ViTs and CNNs, such as better training strategies, preventing attention from focusing on a single block, or discarding low-attention embeddings. However, these methods still follow the design of traditional supervised adversarial training, limiting the potential of adversarial training on ViTs. This paper proposes a novel defense method, MIMIR, which aims to build a different adversarial training methodology by utilizing Masked Image Modeling at pre-training. We create an autoencoder that accepts adversarial examples as input but takes the clean examples as the modeling target. Then, we create a mutual information (MI) penalty following the idea of the Information Bottleneck. Among the two information source inputs and corresponding adversarial perturbation, the perturbation information is eliminated due to the constraint of the modeling target. Next, we provide a theoretical analysis of MIMIR using the bounds of the MI penalty. We also design two adaptive attacks when the adversary is aware of the MIMIR defense and show that MIMIR still performs well. The experimental results show that MIMIR improves (natural and adversarial) accuracy on average by 4.19% on CIFAR-10 and 4.64% on ImageNet-1K, compared to baselines. On Tiny-ImageNet, we obtained improved natural accuracy of 2.99% on average and comparable adversarial accuracy. Our code and trained models are publicly available [^1] https://github.com/xiaoyunxxy/MIMIR.
Index Terms:
Mask Image Modeling, Adversarial Robustness, Vision Transformer, Mutual Information
1 Introduction
While Vision Transformers [28] and the variants [51, 24] have achieved great advances, these attention-based technologies fail in being more robust than CNNs against evasion attacks in common threat models [6, 1, 57, 87, 35]. Evasion attacks [11, 71] (also known as adversarial attacks) refer to fooling well-trained deep models by adding human-imperceptible perturbations to inputs. To build robust transformer models, adversarial training is considered one of the most successful methods [87, 57, 25], in which the general idea is to use adversarial examples for training. However, adversarial training has limitations due to its expensive computational cost to generate adversarial examples. Significant efforts have been made to decrease the adversarial training cost [67, 86, 80, 69], but the nature of attention-based transformer introduces new challenges for accelerating adversarial training. For example, transformers lack inductive biases [28], including locality, two-dimensional neighborhood structure, and translation equivariance. These biases are inherent to CNNs as prior knowledge but not applicable to ViTs [28]. The lack of such prior knowledge makes ViTs require a large amount of training data to generalize well, which means generating more adversarial perturbations for adversarial training.
Due to the differences between CNN and ViTs, previous works try to design new methodologies for adversarial training when considering the robustness of ViTs. Gu et al. [35] proposed Smoothed Attention to improve the robustness against patch-wise attacks by blocking attention to individual patches. Attention-Guided Adversarial Training (AGAT) [87] can efficiently drop image embeddings with lower attention and keep the higher ones after each layer of self-attention, accelerating adversarial training. Mo et al. [57] found that the robustness of ViTs can be improved by randomly masking perturbations or masking gradients from some attention blocks. In addition to new methodologies, some work has also tried to find better training strategies for the robustness of ViTs. The authors of [57] observed beneficial behaviors for adversarial training of ViTs, including fine-tuning on clean pre-trained models, using gradient clipping, and combining SGD optimizer and piecewise learning rate scheduler. Bai et al. [6] discussed that adversarial examples can be too difficult for ViTs to learn when combined with strong data augmentation, such as Randaugment [21], CutMix [95], and MixUp [97]. Then, they proposed a warmup strategy for augmentations to alleviate this issue. They progressively increase the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix. Debenedetti et al. [25] also found that strong data augmentations harm adversarial training. They proposed to use basic data augmentations (random-resize-and-crop + horizontal-flipping + color-jitter) and high weight decay for better performance.
However, previous research focuses on adversarial training at the fine-tuning stage or lacks rigorous explanations, limiting the potential for improving adversarial training. For example, AGAT [87] accelerates adversarial training by dropping a constant ratio of low-attention image embeddings after every layer. Overall, they reduce 40% computation of the entire model. Designing a method that discards more image embeddings is challenging because the accuracy of clean inputs also decreases when a larger number of embeddings are dropped. We investigate a novel approach where we conduct adversarial training at the pre-training stage. Our motivation is to discard a much larger ratio of image embeddings at the pre-training stage, further accelerating adversarial training. In addition, we show a very intuitive analysis of the training process (i.e., decreasing MI between adversarial inputs and latent representation) with theoretical proof, which ensures a good performance. The above motivations yield a self-supervised learning task suitable for Masked Image Modeling (MIM). MIM refers to learning from unlabeled data through masked inputs to recover the original information, such as recovering the original visual tokens [7] or original image pixels [36] based on the masked image patches, masking out a portion of the input sequence and then predicting the feature of the masked regions [85]. Usually, MIM methods can mask out a much larger ratio of inputs by self-supervised pre-training. For example, MAE [36] masks out 75% of image patches at the beginning of the forward pass. The reason is that MIM intends to learn more discriminative information, so masking out the discriminative part of images and recovering it gives a more difficult reconstructing task [81]. The reconstructing task is more difficult when applying an adversarial perturbation to the original image, as the perturbation corrupts the natural features in clean images. In this way, the model learns more discriminative information, which is also robust to adversarial perturbation. Consequently, adversarial training can be both efficient and effective.
Specifically, we propose Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR) as a pre-training method, where we use adversarial examples as inputs and clean examples as reconstructing objectives. As illustrated in Figure 1, MIMIR [^2] MIMIR is a figure in Norse mythology renowned for his knowledge and wisdom. consists of an encoder and a decoder following the design of MAE [36]. The encoder is a ViT, which has the same architecture as the model to be fine-tuned but without the final classification layer. The decoder consists of a few lightweight ViT blocks that reconstruct the original clean image from the latent features generated by the encoder. At pre-training, adversarial examples are separated into image patches. A large ratio of image patches is masked, so the encoder only processes a small part of adversarial patches. At fine-tuning, the decoder is discarded. The encoder is combined with the initialized classification layer and applied to complete images for classification tasks. As we use latent representation from adversarial examples to reconstruct clean images, this represents a more difficult reconstruction task. In addition, we theoretically analyze the behaviors of our encoder and decoder following the idea of Information Bottleneck [72]. The MI between adversarial inputs and latent representation will decrease concerning the accuracy of adversarial examples while pre-training. This motivates us to directly add an MI penalty (between adversarial inputs and latent representation) into the loss function of MIMIR at pre-training. This new pre-training loss further improves performance at the fine-tuning stage (see Table VII). We perform comprehensive experiments on CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26] and compare to state-of-the-art results. The results demonstrate that MIMIR significantly improves the performance on both natural and adversarial accuracy. Finally, previous works [25, 57, 6] claim that strong data augmentations (Randaugment [21], CutMix [95], and MixUp [97]) are unsuccessful in adversarial training. However, we show in Table VI that strong augmentations actually improve the natural and adversarial accuracy with the longer training schedule. Our key contributions are summarized as follows:
  • We propose a novel and efficient defense against adversarial attacks on ViTs - MIMIR, which discards 75% of image patches and then reconstructs the complete clean image under a self-supervised framework.
  • We provide a theoretical analysis on using adversarial examples and MI penalty in MIMIR. We show that MI between adversarial examples and learned latent representation should be decreased.
  • We systematically evaluate MIMIR using three datasets and three transformer architectures with multiple model scales under various adversarial attacks.
  • We show MIMIR is resistant to two adaptive attacks when the adversary is aware of MIMIR. First, we build a PGD-feature attack inspired by [65], which directly increases the distance between features extracted from clean and adversarial examples. Then, we build a PGD MI attack that uses cross entropy and the MI penalty as learning objectives to find adversarial examples.
Figure 1: The structure of MIMIR. First, we generate adversarial perturbation $\delta$ and apply it to the input image. Then, following the original ViT, we use separated image patches as input for the autoencoder. We use the decoder output, i.e., $x_{re}$, and the clean input image, i.e., $x$, to calculate the loss. Finally, we use MI ($I(x+\delta,z)$) as an additional penalty in the pre-training loss. After pre-training, the encoder plus a classification layer is fine-tuned for classification tasks.
2 Background
2.1 Adversarial Attack
[…]
  1. 1.
    The generated perturbation $\delta$ can successfully mislead the network by maximizing (e.g., PGD attack [55]):
    $\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}),$ (1)
[…]
  2. 2.
    The generated adversarial examples are as similar as possible to the original clean examples by limiting $\delta$ to a relatively small domain:
[…]
    where $S$ is the $l_{\infty}$-ball of radius $r$ at a position $x_{i}$ in $\mathbb{R}^{d_{X}}$. The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{0},l_{2},$ and $l_{\infty}$.
When using the above attacks to generate adversarial examples for training, the learning objective is:
$\mathop{\min}\limits_{\theta}\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}).$ (3)
2.2 Masked Image Modeling - MIM
MIM refers to a self-supervised pre-training framework that aims to reconstruct pre-defined targets, such as discrete tokens [7], raw RGB pixels [36, 49], or features [85]. The final goal is to use the pre-trained model as a starting point for downstream fine-tuning. The downstream tasks include, for instance, classification and object detection. More specifically, to build a high-performance ViT $f_{e}$ without a classification layer, we consider $f_{e}$ as an encoder to extract discriminative input features. Then, we design a lightweight decoder $f_{d}$, which uses the output of $f_{e}$ as its input. The goal of $f_{d}$ is to reconstruct the original inputs (let us consider MAE [36] as an example). The aim is to decrease the distance between the input $x$ and $x_{re}=f_{d}\circ f_{e}(x)$. After the encoder $f_{e}$ and decoder $f_{d}$ are trained, we use $f_{e}$ plus a manually initialized classification layer as the starting point of fine-tuning.
2.3 Mutual Information - MI
MI measures the mutual dependence between two random variables, $X$ and $Y$. It quantifies the amount of information contained in one random variable about another random variable or the reduced uncertainty of a random variable when another random variable is known. It can be written as:
$I(X,Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{(X)}(x)P_{(Y)}(y)}\right),$ (4)
where $P_{(X,Y)}$ is the joint probability density function of $X$ and $Y$. $P_{(X)}$ and $P_{(Y)}$ are the marginal probability density function of $X$ and $Y$. In addition, MI can be equivalently expressed as:
$I(X,Y)$ $=H(X)-H(X|Y).$ (5)
However, quantifying MI is difficult in practice, as it is difficult to precisely estimate $P_{X,Y}$ or $P_{X}$ and $P_{Y}$ in high dimensional space [31]. DIB [94] suggested using the matrix-based Renyi’s $\alpha$-order entropy $I_{\alpha}$ [66, 93] to estimate MI, which avoids density estimation and variational approximation. An alternative way is the Hilbert-Schmidt Independence Criterion (HSIC) [34], which is a kernel-based dependence measure defined in a reproducing kernel Hilbert space (RKHS) and usually used as a surrogate of MI. Details about definitions and empirical estimators of $I_{\alpha}$ and HSIC are provided in Appendix G. In this paper, we use both $I_{\alpha}$ and HSIC as our MI measurements.
2.4 Information Bottleneck - IB
The IB concepts were first proposed in [72] and further developed for deep learning in [73, 68]. It has been widely used to improve adversarial robustness [90, 2, 94, 84]. IB describes the generalization of a deep network in two phases: 1) empirical error minimization (ERM) and 2) representation compression [68]. For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer. The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$. The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I(\cdot)$. During neural network training, in the ERM phase, the model increases shared information between $t_{l}$ with respect to both $x$ and $y$. Afterward, in the compression phase, the model decreases information contained in $t_{l}$ about $x$ but preserves (or even increases) information about $y$. The reduction of $I(x,t_{l})$ can be interpreted as a way of reducing noise or compressing irrelevant or redundant features in $x$. At the end of the training, the model strikes a trade-off that maximizes $I(y,t_{l})$ and minimizes $I(x,t_{l})$. Formally, the IB minimizes the following Lagrangian:
$\mathcal{L}=I(x,t_{l})-\beta I(y,t_{l}),$ (6)
where $\beta$ is a Lagrange multiplier that controls the trade-off between predicting $y$ and compressing $x$.
3 MIMIR
3.1 Threat Model
We assume the adversary intends to create adversarial examples from clean images. The adversary has the strongest ability, i.e., the white box, to access the target model. Then, adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets. In addition, the adversary is also aware of potential defenses. An adversary can design new attacks for a specific model according to the details of our defense method.
From the defender’s perspective, the main goal is to train a robust model against potential adversarial attacks. The defender considers the following four objectives:
  • The performance on clean data should not decrease significantly. We allow a small drop of clean accuracy for a trade-off in robustness.
  • The accuracy of models trained by MIMIR should be significantly higher than for models without defenses under various adversarial attacks.
  • Deploying MIMIR should be efficient and scalable to large datasets such as ImageNet-1K [26].
  • MIMIR should be robust against adaptive attacks.
3.2 Design Intuition
MIM has been proved useful as a pre-training method for ViTs on various tasks [79, 36, 85, 7]. To train a powerful model by MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model. Masking out the foreground instead of the background reduces discriminative information in visible information to the model. Reconstructing foreground parts is harder than background and helps the model learn more discriminative information [81]. Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbation to clean inputs. The adversarial perturbation increases the distance between the input and reconstruction target. If we can reconstruct clean inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks. In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial inputs while enabling the decoder to reconstruct the clean image. Note that simply masking out all foreground does not make sense, as it would not be possible to reconstruct meaningful content. From the perspective of IB, there is a bottleneck between the encoder and decoder. Our adversarial pre-training task contains two information sources: the clean data $x$ and the adversarial perturbation $\delta$. As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated. The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.
3.3 Design Details
Autoencoder. MIMIR follows the general design of MAE [36], which consists of an encoder $f_{e}$ and a decoder $f_{d}$. Like all other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$. The decoder reconstructs original inputs according to the discriminative features. Following ViT [28], the input $x$ is separated into non-overlapping image patches. Then, we randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder. This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it is not possible to find the reconstruction target then. Therefore, we hope to keep a part of the foreground as a hint for reconstruction. The information of masked content is recorded as mask tokens $m$, which are not used by the encoder but will be used later by the decoder. Each token is a learned vector that indicates the presence of a masked patch to be predicted. The mask token is shared by all inputs of the decoder. Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image. We emphasize that mask tokens are not used in the encoder part.
To train a ViT, we use the same transformer blocks as ViT to build the encoder. The encoder only processes the visible patches, making training much more efficient. When converting to other architectures, such as ConVit [24], we use corresponding transformer blocks to build the encoder. The decoder accepts the encoded visible image patches and mask tokens as inputs. The decoder is built using the same transformer blocks as the encoder instead of using ViT [28] transformer blocks for all. Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
Adversarial pre-training target. The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder. Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results. Following the original MAE [36], this distance is measured by Mean Squared Error (MSE). To create a more difficult reconstruction task, we apply adversarial perturbation $\delta$ on the inputs of the encoder. Thus, the adversarial perturbation is also masked along with the image when inputted into the encoder. The decoder reconstructs the original clean inputs by using the latent features $z$ extracted from adversarial examples. The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model. Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets. Formally, the pre-training process (described in Algorithm 1) can be written as follows:
$z$ $=f_{e}(x+\delta),$ (7)
$x_{re}=f_{d}(z),$
$loss_{\text{mse}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re}).$
MI as penalty. Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases with respect to the accuracy on adversarial example, i.e., $I(x+\delta,z)$ is decreasing while training. Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
$loss_{\text{mi}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re})+\lambda I(x+\delta,z),$ (8)
where $\lambda$ is a regularizer for the MI penalty. We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty. This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$. According to Data Processing Inequality (DPI) [8], $I(x,z)\leq I(x+\delta,z)$. Minimizing $I(x,z)$ in loss equals minimizing a lower bound of $I(x+\delta,z)$.
Algorithm 1 Masked Adversarial Pre-Training
0:  training data $D$, number of epochs $E$, encoder $f_{e}$, decoder $f_{d}$, network parameters $\theta$, $\mathcal{L}_{\text{mse}}$, $\lambda$.
0:  optimized weights $\theta$
1:  for $e=0\to E-1$ do
2:     $x\leftarrow$ sample_batch($D$)
3:     $\delta\leftarrow$ random_initialization
4:     $x_{re}\leftarrow f_{d}\circ f_{e}(x+\delta)$
5:     $\delta\leftarrow\mathop{max}\limits_{\delta\in S}\mathcal{L}_{\text{mse}}(x+\delta,x_{re})$
6:     Forward:
7:     $z\leftarrow f_{e}(x+\delta)$
8:     $x_{re}\leftarrow f_{d}(z)$
9:     $loss\leftarrow\mathcal{L}_{\text{mse}}(x,x_{re})+\lambda I(x+\delta,z)$
10:     Backward:
11:     $\theta\leftarrow\theta-\alpha\nabla loss$
12:  end for
Generating adversarial examples. To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbation $\delta$. As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD [55]. Still, we can easily design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq. (7). As the feature $z$ is extracted from only visible image patches, we only attack the visible patches. We do not add any perturbation to mask tokens since the outputs of the autoencoder are only impacted by visible patches. Then, the adversarial pre-training learning objective can be written as:
$\begin{split}\mathcal{L}_{\text{adv}}=\mathop{\max}\limits_{\delta\in S}&\mathcal{L}_{\text{mse}}(f_{d}\circ f_{e}(x+\delta),x),\\ \mathop{\min}\limits_{\theta}\mathcal{L}_{\text{adv}}&+\lambda I(x+\delta,z),\end{split}$ (9)
where $\theta$ are the parameters of the autoencoder.
Adversarial fine-tuning. After the autoencoder is trained, we discard the decoder and initialize a classification layer for the encoder to build a complete model. We use simple adversarial training, PGD [55], to fine-tune our pre-trained models so that we can ensure the improvements come from the MIMIR pre-training. The complete model is trained at the fine-tuning stage.
3.4 Theoretical Justification
In this section, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained. Let $F$ be any classifier trained on clean samples with desirable prediction accuracy, which may suffer from adversarial attacks. We begin our analysis by first presenting Lemma 3.1.
Lemma 3.1
Given an arbitrary classifier $F$, let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$. We have:
$I(F(x+\delta),F(x_{re}))\leq I(F(x+\delta),x_{re})\leq I(x+\delta,z).$ (10)
Proof:
There are two Markov chains:
$x+\delta\rightarrow F(x+\delta),$ (11)
$z\rightarrow x_{re}\rightarrow F(x_{re}),$
which indicate that $F(x+\delta)$ is an indirect observation of $x+\delta$, whereas both $F(x_{re})$ and $x_{re}$ are indirect observations of $z$.
By the data processing inequality (DPI), we have
$I(x+\delta,z)\geq I(F(x+\delta),z),$ (12)
and
$I(F(x+\delta),z)\geq I(F(x+\delta),x_{re})\geq I(F(x+\delta),F(x_{re})).$ (13)
Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$. Intuitively, our autoencoder is trained to only recover clean sample $x$ without any interference from $\delta$. Hence, a relatively large value of $p_{e}$ is expected. In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
Proposition 3.2
Let $H(\cdot)$ denote the information entropy and $H_{b}(p_{e})=-p_{e}\log_{2}p_{e}-(1-p_{e})\log_{2}(1-p_{e})$ be the binary entropy, we have:
$H(F(x+\delta))-H_{b}(p_{e})-p_{e}\log(|F(x+\delta)|-1)\leq I(x+\delta,z),$ (14)
where $|F(x+\delta)|$ is the total number of categories [^3] For instance, for CIFAR-10 [46], $|F(x+\delta)|=10$..
Proof:
By the chain rule of MI, we have
$I(F(x+\delta),F(x_{re}))=H(F(x+\delta))-H(F(x+\delta)|F(x_{re})).$ (15)
By applying Fano’s inequality [30, 58], we obtain:
$H(F(x+\delta)|F(x_{re}))\leq H_{b}(p_{e})+p_{e}\log(|F(x+\delta)|-1).$ (16)
Adding $I(F(x+\delta),F(x_{re}))$ to both sides of Eq. (16):
$H(F(x+\delta))$ $-H_{b}(p_{e})-p_{e}\log(|F(x+\delta)|-1)$ (17)
$\leq I(F(x+\delta),F(x_{re}))$
$\leq I(x+\delta,z).$
The last line of Eq. (17) is by Lemma 3.1. ∎
Therefore, we obtain a lower bound of $I(x+\delta,z)$. If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$.
Figure 2: The example plots for the lower and upper bounds on the MI in Propositions 3.2 and 3.3. The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes. The lower bound reaches its minimum at $p_{e}=0.9$.
In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (11).
Proposition 3.3
If $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, we have:
$I(x+\delta,z)\lesssim H(F(x_{re}))-2p_{e},$ (18)
in which the notation $\lesssim$” refers to less than or similar to.
Proof:
By the Hellman-Raviv inequality [38, 13], we have:
$2p_{e}$ $\leq H(F(x+\delta)|x_{re})$ (19)
$=H(F(x+\delta))-I(F(x+\delta),x_{re})$
$\approx H(F(x+\delta))-I(x+\delta,z).$
Similar to the lower bound, the upper bound also indicates $I(x+\delta,z)$ is inversely proportional to $p_{e}$ as shown in Figure 2.
In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$. For simplicity, let us assume the clean data $x$ and adversarial perturbation $\delta$ are independent, then:
$I(x+\delta,z)=I(x,z)+I(\delta,z).$ (20)
According to [78], minimizing the expected reconstruction error between clean sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input. Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized). In other words, only the information about $\delta$ has been removed from $z$. This also explains the robustness of $z$.
TABLE I: Different learning rates. Fine-tuned for 50 epochs.
[…]
CIFAR-10 ViT-T 5.0e-4 76.30 47.60
1.0e-3 80.69 49.56
1.0e-2 85.62 48.78
5.0e-2 85.12 50.30
1.0e-1 84.51 50.40
TABLE II: Top 1 natural and adversarial accuracy on CIFAR-10 test set. The results of ARD+PRM [57] are taken from the original paper. As there is no ViT-T architecture, we use the same architecture as DeiT-T.
[…]
CIFAR-10 [46] ($l_{\infty},\epsilon=8/255$) ViT-T (DeiT-T) ImageNet-1K PGD 75.46 48.10 47.96 43.62
MAE [36] PGD 79.89 48.64 48.43 44.27
ImageNet-1K ARD+PRM [57] 79.60 50.33 50.15 45.99
MIMIR PGD 84.82 57.51 57.38 52.96
ViT-S ImageNet-1K PGD 79.59 50.86 50.73 46.37
MAE [36] PGD 85.97 54.19 53.94 50.49
ImageNet-1K ARD+PRM [57] 81.86 51.73 51.46 47.33
MIMIR PGD 88.11 56.63 56.34 53.18
ViT-B ImageNet-1K PGD 83.16 52.98 52.71 49.06
MAE [36] PGD 88.86 57.03 56.86 52.42
ImageNet-1K ARD+PRM [57] 84.90 53.80 53.51 50.03
MIMIR PGD 89.30 58.07 57.87 54.55
ConViT-T ImageNet-1K PGD 53.09 33.63 33.61 29.65
MAE [36] PGD 75.49 45.29 45.17 41.12
ImageNet-1K ARD+PRM [57] 80.28 49.86 49.55 45.42
MIMIR PGD 80.74 49.37 49.16 45.04
ConViT-S ImageNet-1K PGD 54.03 34.61 34.60 30.60
MAE [36] PGD 86.43 54.76 54.59 51.13
ImageNet-1K ARD+PRM [57] 84.32 53.10 52.81 48.85
MIMIR PGD 87.49 56.35 56.20 52.54
ConViT-B ImageNet-1K PGD 61.54 38.77 38.71 34.21
MAE [36] PGD 89.03 58.19 57.75 54.25
ImageNet-1K ARD+PRM [57] 85.80 53.36 53.05 49.33
MIMIR PGD 89.30 59.31 59.01 55.64
4 Experimental Evaluation
4.1 Experimental Setup
We use PyTorch [60] 2.1.0 for implementation, and mixed precision [56] to accelerate training. We use three datasets to comprehensively evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26], details of datasets are given in Appendix A. Appendix B provides details about decoder hyperparameters.
Model Architecture: We evaluate MIMIR with three commonly used architectures with multiple scales: ViT [28], ConViT [24], and XCiT [3].
Hardware. Experiments are performed on a single machine with 4 RTX A6000 (48GB) and 4 RTX A5000 (24GB) GPUs, CUDA 12.0.
Training hyperparameters. For all experiments, we train models from scratch (unless otherwise specified). Following MAE [36], we do pre-training by MIMIR for 800 epochs and then fine-tuning for 100 epochs. The number of warmup epochs is 40 and 10 for pre-training and fine-tuning, respectively. We use AdamW [53] as an optimizer for both pre-training and fine-tuning. We use cosine decay as the learning rate scheduler.
At pre-training, one-step FGSM with random initialization achieves better performance than using ten-step PGD in Table VII. Therefore, we use a one-step algorithm to generate adversarial examples for all three datasets. The perturbation budget is $\epsilon=8,\alpha=10$. At fine-tuning stage, we use a 10-step PGD algorithm to generate adversarial examples for training to be aligned with baselines. The perturbation limit is $\epsilon=8,\alpha=2$. When fine-tuning with ImageNet-1K, we use one-step FGSM with random initialization to align with baselines and for better efficiency. Following previous research [86, 87], we fine-tuned twice for two different perturbation limits, $\epsilon=2,\epsilon=4$.
As discussed in Section 4.3, strong data augmentation is harmful to adversarial training at a common training schedule (50 or 100 epochs for fine-tuning). We only use simple augmentations, including random resized crops and random horizontal flips.
Previous research used rather different fine-tuning learning rates, so we chose the learning rate by testing different values as given in Table I. We use the best, i.e., $25.0e-$ 2, as the base learning rate (blr) for CIFAR-10. The learning rate is calculated based on blr, i.e., $lr=blr\times\frac{batchsize}{256}$. For Tiny-ImageNet and ImageNet-1K, we use the same blr as MAE [36], i.e., $0.001$. At the pre-training stage, we use $41.5e-$ 4 as the learning rate for all datasets. In addition, we noticed that it is possible to get NaN (Not a Number) for loss in a few cases while adversarial training, i.e., the training fails. We solve this problem by using a learning rate that is ten times smaller. More details about hyperparameters can be found in the Appendix D.
TABLE III: Top 1 natural and adversarial accuracy on ImageNet-1K.
Dataset Models Pre-train Fine-tune $\epsilon=2/255$ $\epsilon=4/255$
[…]
ImageNet-1K [26] ($l_{\infty}$) ViT-S - FastAT [86] 71.37 49.49 49.41 66.92 34.07 33.82
- AGAT [87] 70.62 49.00 48.85 66.10 33.62 33.40
MIMIR FastAT [86] 74.60 54.56 54.55 71.29 40.98 40.63
ViT-B - FastAT [86] 70.31 50.55 50.06 65.18 33.59 33.39
- AGAT [87] 70.41 51.23 51.11 67.93 34.94 34.78
MIMIR FastAT [86] 75.88 55.42 55.36 73.22 41.26 40.46
CaiT-XXS24 - FastAT [86] 72.84 54.31 54.26 68.77 32.14 32.09
- AGAT [87] 71.15 54.17 54.08 68.22 31.46 31.39
MIMIR FastAT [86] 73.39 53.39 53.37 69.90 40.53 40.22
CaiT-S36 - FastAT [86] 72.51 53.12 52.76 71.20 33.02 32.84
- AGAT [87] 72.69 53.66 53.48 71.06 33.46 33.17
MIMIR FastAT [86] 76.05 56.78 56.75 73.57 40.03 39.16
TABLE IV: Comparison of different architectures under 100-step PGD attack.
Dataset Models $\epsilon=2/255$ $\epsilon=4/255$
Natural PGD Natural PGD
ImageNet-1K ($l_{\infty}$) ViT-S 74.60 54.56 71.29 40.63
ConViT-S 74.72 55.65 72.08 42.62
XCiT-S 74.39 55.21 71.77 40.97
XCiT-M 73.63 53.04 72.01 42.39
[…]
Tiny-ImageNet [47] ($l_{\infty},\epsilon=8/255$) WRN-28-10 Gowal et al. [33] $\times$ 512 400 51.56 N/A N/A 21.56
WRN-28-10 Wang et at. [83] 1m 512 400 53.62 N/A N/A 23.40
[…]
ViT-S MIMIR 1m 512/256 800/100 63.33 26.37 26.17 22.52
ViT-B MIMIR 1m 512/256 800/100 65.06 25.41 25.03 22.82
ConViT-T MIMIR 1m 512/256 800/100 52.61 24.61 24.58 19.42
ConViT-S MIMIR 1m 512/256 800/100 62.29 26.39 26.30 22.88
Evaluation metrics. We use natural accuracy and adversarial accuracy as evaluation metrics. Natural accuracy refers to the accuracy of a model on clean and unmodified inputs. The adversarial accuracy measures the accuracy of a model under a specific adversarial attack. To show that the improvement is not from obfuscated gradient [5], we use PGD [55] and AutoAttack [20] to evaluate the robustness of trained models. PGD is one of the most popular adversarial attacks used for generating adversarial examples. We evaluate MIMIR under 20 and 100 steps of PGD attack. AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks. In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce [20], APGD-t [20], FAB-t [19], and Square [4].
We use the same perturbation budget for our experiments to compare with previous works. We apply the commonly-used $l_{\infty}$ norm for PGD [55] and AutoAttack (AA) [20]. We set the perturbation radius $\epsilon$ to $8/255$ for CIFAR-10 [46] and Tiny-ImageNet [47]. The step length is $\alpha=2/255$. As the image size of ImageNet-1K is much larger, we use different scales of $\epsilon$ for adversarial training and evaluation [26]. Following previous works [57, 87, 50], we use two perturbation radius $2/255,4/255$ for evaluation. The step length is $\alpha=1/255$.
4.2 Experimental Results
Overall performance. We apply MIMIR to CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26]. The results are shown in Tables IIIII, and V, respectively. Overall, MIMIR significantly improves adversarial accuracy under PGD and AA attacks. While training with CIFAR-10 on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.50% compared to the baseline. On ConViTs (ConViT-T, ConViT-S, and ConViT-B), MIMIR improves the accuracy by an average of 2.87% compared to the baseline. While using ImageNet-1K on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.52% compared to the baseline. When fine-tuning with clean data, MIMIR also shows improved natural accuracy in Table XIV. In addition, we also evaluate the time consumption and memory usage of MIMIR in Table X. We compare MIMIR with FastAT [86] and 10-step PGD adversarial training. The results show that using MIMIR for pre-training is much more efficient than traditional adversarial training.
[…]
Comparison with related work. We present the performances of MIMIR compared with recent baseline methods in Tables II and III. In Table II, we perform experiments using CIFAR-10 [46]. The results are compared with three baselines, including pre-training on ImageNet-1K and fine-tuning with PGD, MAE [36], and ARD+PRM [57]. The “ImageNet-1K” under the Pre-train column refers to using models pre-trained on ImageNet-1K, then fine-tuned with PGD for 40 epochs, which is the first baseline in ARD+PRM [57]. The MAE [36] baseline is pre-trained using clean CIFAR-10 images and fine-tuned with 10-step PGD adversarial training. The training schedule of MAE [36] baseline is 800 epochs for pre-training and 100 epochs for fine-tuning. Our MIMIR uses the same training schedule as MAE [36]. We provide results on fine-tuning with clean images in Appendix E. The ARD+PRM [57] baseline refers to the best algorithm in [57], where we use the results from that work. Note that ARD+PRM [57] baseline uses models pre-trained on ImageNet-1K. Using models pre-trained on ImageNet-1K and then fine-tuning for CIFAR-10 can easily get accuracy higher than 95% [75].
As MIMIR only uses CIFAR-10 for training, it is more challenging to achieve good performance. For a more fair comparison, we use elucidating diffusion model (EDM) data as a data augmentation since the baselines also use EDM data (Wang et at. [83] and Rebuffi et al. [62]) or models pre-trained on ImageNet-1K (ARD+PRM [57]). Generative data is usually used to improve adversarial training [83, 61, 63, 33]. Specifically, we use 5 million generated CIFAR-10 data and 1 million Tiny-ImageNet data provided by [83]. Note that the EDM data is generated only based on CIFAR-10 or Tiny-ImageNet, and the MAE baseline uses EDM data as well. The EDM data is applied to experiments with CIFAR-10 and Tiny-ImageNet but not to ImageNet-1K.
According to results in Table II, MIMIR consistently improves adversarial robustness compared to all baselines. It also improves the natural accuracy compared to the MAE baseline. This is because self-supervised MIMIR does not suffer from inconsistency between the ground truth and adversarial example [59]. Let us consider a joint data distribution $p_{d}(x,y)$ (training data) and a discriminative model $p_{\theta}(y|x)$, where $x$ is the data with its label $y$. In standard training, we hope the model learns discriminative features by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x)$. In adversarial training, we hope the model is robust by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x+\delta)$, i.e., Eq. (3). This means $p_{\theta}(y|x+\delta)$, instead of $p_{\theta}(y|x)$, is encouraged to converge to $p_{d}(y|x)$. Supervised adversarial training builds a shortcut from $x+\delta$ to $y$. For the optimal $\theta^{*}$ of minimizing Eq. (3), $p_{\theta^{*}}(y|x)\neq p_{d}(y|x)$ [59]. Therefore, there is always a trade-off between natural and adversarial accuracy in supervised adversarial training. However, such inconsistency does not apply to self-supervised MIMIR, as there is no label for self-supervised training. MIMIR encourages $p_{\theta_{ae}}(x_{re}|x+\delta)$ to converge to $p_{d}(x)$, which does no affect convergence of $p_{\theta_{ae}}(x_{re}|x)$ to converge to $p_{d}(x)$. This implies that supervised pre-training is harmful to natural accuracy, which is also aligned with the findings of previous research [39].
Table III shows the results evaluated on the ImageNet-1K [26] dataset. The performance of MIMIR is compared with two baselines, including FastAT [86] and AGAT [87]. The FastAT [86] is designed for efficient adversarial training so that it can be applied to large models and large datasets, such as ImageNet-1K. The main idea of FastAT is to use a single-step adversarial attack to generate adversarial examples. The authors refined the standard FGSM [32] by applying a random start at the beginning of the algorithm and a much larger step size to reach the perturbation limit. According to empirical results, they used $\alpha=1.25\times\epsilon$ for step size, which provided comparable performance to the best-reported result from free adversarial training [25]. The results of FastAT [86] and AGAT [87] are taken from [87]. The baselines of FastAT [86] and AGAT [87] are trained for 300 epochs with a cosine decay learning rate schedule, and the first 20 epochs for learning rate warmup. As shown in Table III, the accuracy of AGAT [87] may decrease compared to FastAT [86]. This is because AGAT [87] discards low-attention image patches for efficiency. The image patches, even with the low-attention scores, are useful for classification. In Table IV, we compare the performance among three architectures with multiple scales. MIMIR shows similar performance on other architectures as on the original ViTs, which means MIMIR is well-generalized to different architectures.
Figure 3: The loss and natural accuracy for three different augmentation strategies, using CIFAR-10 and ViT-small.
As CIFAR-10 is small compared to ImageNet-1K, we also provide experiments on Tiny-ImageNet for completeness. We use the same training schedule and hyperparameters for experiments on Tiny-ImageNet. As AGAT [87] and ARD+PRM [57] do not provide experiments on Tiny-ImageNet, we compare our results with the most recent adversarial robustness results from Wang et at. [83] and Rebuffi et al. [62]. They use generated data as augmentation for more robust adversarial training. Specifically, Gowal et al. [33] study four methods for generated data as augmentation, including Denoising Diffusion Probabilistic Model (DDPM) [41], Very Deep Variational Auto-Encoder (VDVAE) [18], StyleGAN [45], and BigGAN [12]. Wang et at. [83] demonstrated better diffusion models and further improved adversarial robustness compared to previous works. In Table V, we show our experimental results compared with previous works.
Our method provides competitive performance. Especially when the baselines only use generated data based on Tiny-ImageNet, MIMIR provides the best natural accuracy (65.06%). When the baselines use generated data based on ImageNet-1K, and we only use that based on Tiny-ImageNet, MIMIR still shows comparable performance. In addition, according to Table X, the MIMIR training schedule (800 pre-training + 100 fine-tuning) takes approximately 200 fine-tuning (10-step PGD) epochs of training time. Thus, MIMIR costs much less training time and gets comparable results to baselines.
TABLE VI: Data augmentation ablation with longer fine-tuning schedule.
[…]
ViT-B 800 Weak Augmentation 89.90 60.26
+ CutMix [95],MixUp [97] 91.01 60.62
+ Randaugment [21] 90.19 62.75
4.3 Data Augmentation is Harmful
As mentioned in previous works [6, 25], strong data augmentation makes the training samples too difficult to be learned by ViTs while conducting adversarial training. The strong data augmentation refers to the combination of Randaugment [21], CutMix [95], and MixUp [97]. In this section, we evaluate two different solutions to ease this problem. First, we only use simple data augmentation for adversarial training, including random crop (or random resize crop for ImageNet-1K) and random horizontal flip (“weak aug”). Second, we use a 10-epoch warmup procedure for strong data augmentation. The warmup of Randaugment is implemented by progressively increasing the distortion magnitude from 1 to 9 (“warmup aug”). For CutMix and MixUp, we warm up by increasing the mixup probability from 0.5 to 1.0. As shown in Figure 3, “weak aug” provides the best accuracy. The “warmup aug” shows a slightly improved accuracy compared to fusing strong augmentation. Therefore, we provide a different result from [6] on the smaller dataset CIFAR-10, i.e., we show that weak augmentation is better than warmup augmentation. Even data augmentation with reduced amplitude is still difficult to learn at the beginning of adversarial training. Although strong augmentation is harmful to a normal training schedule, we show in Table VI that CutMix [95], MixUp [97], and Randaugment [21] increase the accuracy of adversarial training when training with a longer schedule, e.g., 800 epochs of fine-tuning. We conjecture that the data with strong augmentation is helpful but difficult for adversarial training to learn. Thus, it needs more epochs to learn meaningful representation. Loss and accuracy curves can be found in Appendix C. Due to the longer schedule, we use patch size 4 to reduce training time.
TABLE VII: Comparison between HSIC and $I_{\alpha}$. Models are pre-trained for 800 epochs. For fine-tuning, all models are adversarially trained with 10-step PGD for 50 epochs.
Pre-train $\lambda$ Estimator Natural PGD
MIMIR 0.001 HSIC [34] 69.63 43.17
MIMIR 0.001 $I_{\alpha}$ [94] 75.00 46.11
MIMIR 1e-05 HSIC [34] 76.30 47.60
MIMIR 1e-05 $I_{\alpha}$ [94] 75.53 46.75
MIMIR 1e-06 HSIC [34] 74.90 46.19
MIMIR 1e-06 $I_{\alpha}$ [94] 74.60 45.66
Adv-MAE (1-step) 0.0 - 74.69 46.28
Adv-MAE (10-step) 0.0 - 73.96 45.77
MAE 0.0 - 69.02 42.31
4.4 MI Measures
In Section 3.4, we provide lower and upper bound (Eq. (14)) of $I(x+\delta,z)$. According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the clean image $x$. This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective. Before achieving the goal, we must choose a suitable and efficient MI estimator and a regularizer ($\lambda$ in Eq. (8)). In this paper, we use $I_{\alpha}$ [94] and HSIC [54] as estimators. In Table VII, we show the performance of $I_{\alpha}$ [94] and HSIC [54] with different values of $\lambda$. According to the results, using HSIC as an estimator with $\lambda=1e-05$ shows the best performance, which is used for other experiments.
We also show that MIMIR outperforms original MAE [36] and Adv-MAE [91]. To make it clearer, the MAE baseline in Table VII refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD. Adv-MAE refers to using adversarial examples but not using $I(x+\delta,z)$ in loss. Adv-MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training. Adv-MAE provides better accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task. This more difficult task further improves the performance of downstream models. We use the default learning rate (i.e., $45.0e-$ 4) of MAE [36] in experiments in Table VII. This is aligned with Adv-MAE [91].
Figure 4: The fine-tuning results with different numbers of pre-train epochs. The left column is trained on CIFAR-10, and the right column is on Tiny-ImageNet. The above is natural accuracy, and the below is adversarial accuracy under a 20-step PGD attack.
4.5 Training Epochs Ablation Study
Our experiments so far are based on 800-epoch pre-training. In Figure 4, we show the influence of different numbers of epochs. We use ViT-T and ViT-S as our baseline models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet. We fine-tune each model 50 epochs after MIMIR pre-training. Both adversarial and natural accuracy are improved with a longer training schedule. On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S. On Tiny-ImageNet, the increase is more obvious.
4.6 Fine-tuning Using Advanced Methods
Our fine-tuning thus far is based on traditional 10-step PGD adversarial training for CIFAR-10. The reason is to compare with previous works that use a 10-step PGD algorithm. Next, we consider using TRADES [96] and MART [82] to study whether MIMIR can achieve better performance with better fine-tuning methods. TRADES [96] identify a trade-off between robustness and accuracy on clean data. The authors decompose the prediction error of adversarial examples as natural and boundary errors. According to these two kinds of errors, the authors find a differentiable upper bound based on the theory of classification-calibrated loss. MART [82] finds that misclassified and correctly classified natural examples have different influences on the final robustness after adversarial training. Therefore, the authors designed MART to explicitly differentiate the misclassified and correctly classified natural examples during adversarial training.
In Table VIII, the MIMIR performance could be further improved with the two advanced adversarial training methods. Specifically, TRADES slightly improves the accuracy on clean examples but shows worse robustness compared to PGD. MART improves the adversarial accuracy when using ViT-B, but we notice a significant drop in natural accuracy.
TABLE VIII: The experimental results with different fine-tuning methods.
[…]
ViT-S PGD 88.11 56.63 56.34 53.18
TRADES 88.19 56.42 56.36 51.70
MART 80.55 50.81 50.75 39.92
ViT-B PGD 89.30 58.07 57.87 54.55
TRADES 90.91 57.73 57.58 52.79
MART 85.78 60.35 60.16 54.27
4.7 Adaptive Attack Evaluation
As the adversary is aware of the defense MIMIR, they can design adaptive attacks [76]. For example, the adversary may attack feature space [52, 65] since MIMIR trains backbone to extract features. Here, the backbone refers to the ViT model without the final classification layer, i.e., the encoder of MIMIR. In this section, we provide two adaptive attacks specifically designed against MIMIR. First, we introduce PGD-MI, which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in loss. Therefore, it is reasonable to attack the model by directly increasing the MI $I(x+\delta,z)$. Formally, we embed the following loss into the PGD algorithm:
$\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(x_{i}+\delta,y_{i})+\lambda I(x+\delta,z).$ (21)
We use the same value of $\lambda$ as in MIMIR pre-training.
Second, we introduce PGD-feature that directly attacks the feature extracted by ViT backbones. As the backbones are trained by self-supervised MIMIR pre-training (the whole model is trained when fine-tuning), we hope to evaluate the feature extracted by the backbones. The PGD-feature increases the distance between features extracted from clean and adversarial examples. We implement it using the $L_{2}$ norm and the PGD algorithm:
$\mathop{\max}\limits_{\delta\in S}l_{2}$ $(f_{e}(x),f_{e}(x+\delta)).$ (22)
Table IX shows the adaptive evaluation results for PGD-MI and PGD-feature. We use 100 steps for both to make sure the attacking algorithm converges. The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$. For ImageNet-1K, we use $\epsilon=4/255$. According to the results in Table IX, MIMIR is still robust against adaptive attacks.
4.8 Evaluation of the Loss Surface
To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot loss surfaces [48] in Figure 5. The loss surface is the visualization of the loss function as parameters change. The basic idea is to plot the loss around the optimal parameters. Formally, we consider in the 2D case:
$f_{l}(\alpha,\beta)=L(\theta^{*}+\alpha\theta_{1}+\beta\theta_{2}),$ (23)
where $\theta_{1}$ and $\theta_{2}$ are two direction vectors, $\alpha$ and $\beta$ are two arguments of $f_{l}$. In practice, we use the parameters of trained models, i.e., $\theta^{*}$. In Figure 5, we provide the loss surfaces for models. The surfaces of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
Figure 5: The loss surfaces of the models. The four figures of each column are the loss information of one model. The first column is ConViT-S pre-trained with MIMIR and fine-tuned with adversarial examples. The second column is ConViT-S, pre-trained with clean data and fine-tuned with adversarial examples. The third is ViT-S, pre-trained with clean data and then fine-tuned with adversarial examples. The fourth one is ViT-S, pre-trained with MIMIR and fine-tuned with adversarial examples. The training dataset is CIFAR-10.
TABLE IX: The experimental results for adaptive attacks.
[…]
CIFAR-10 ConViT-S 56.35 56.16 78.52
ViT-S 56.63 56.31 78.41
ViT-B 58.14 57.85 80.49
Tiny-ImageNet ConViT-S 26.39 26.29 58.50
ViT-S 26.37 26.18 57.36
ViT-B 25.41 25.05 58.90
ImageNet-1K ConViT-S 53.86 53.84 72.10
ViT-S 54.56 54.55 72.27
ViT-B 55.41 55.36 73.51
4.9 Efficiency
Table X provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning. The time consumption and memory usage are evaluated on 4 A6000 GPUs. Compared to 10-step PGD and FAST AT, MIMIR shows much less time consumption. At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$. Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time. We also consider using 10-step PGD at pre-training, Table VII shows that using a one-step algorithm with larger $\alpha=10$ provides better performance. In addition, pre-training with MIMIR also costs less memory. Note in Table X that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning. In Table X, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86]. Therefore, MIMIR is more promising concerning training costs when conducting adversarial training.
4.10 Discussion and Limitations
Across our experiments, we have observed the promising performance of MIMIR. This is because we build a different adversarial training paradigm. With MIMIR, the adversarial pre-training can be more efficient and effective. Following the idea of IB, we can intuitively consider a bottleneck between the encoder and decoder. As the reconstruction output is constrained by clean data $x$, the bottleneck will filter out information from adversarial perturbation $\delta$. We provide a theoretical guarantee of this bottleneck in Section 3.4. In Eq. (8), we embed this bottleneck as a learning object to further improve the performance. This improvement also confirms the correctness of our theoretical guarantee in Section 3.4. Table VII shows that our method works better than baseline even without embedding the bottleneck in Eq. (8). With the two information sources of $x$ and $\delta$, the model is trained to learn the robust features from $x$ and forget the information of $\delta$ under the constraint of the reconstruction target.
While MIMIR shows better performance, there are still certain limitations. MIMIR is a pre-training method. An adversarial fine-tuning is necessary to build the final model. Thus, the shortcomings of traditional adversarial training cannot be completely avoided. In our experiments, we utilize the simple PGD algorithm for fine-tuning, but one can further improve MIMIR pre-trained models with more advanced approaches. In addition, MIMIR follows the design of MAE, and we also utilize the characteristic that ViTs are capable of processing variable-length inputs. Therefore, currently, MIMIR cannot handle pyramid-based ViTs and CNNs. While it is not trivial to apply MIMIR to other pyramid-based ViTs and CNNs, this is feasible [89, 49, 81].
TABLE X: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
CIFAR-10 [46] Tiny-ImageNet [47] ImageNet-1K [26]
Models #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
[…]
FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
[…]
FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
[…]
FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
5 Related Work
5.1 Vision Transformer
The transformers [77] were first proposed in natural language processing (NLP). With the mechanism of global self-attention, transformers can effectively capture the non-local relationships among all text tokens [27, 14, 22]. The pioneering work, ViT [28], demonstrated that the pure transformer architecture can achieve competitive performance on various tasks. ViT also reveals that transformers lack inductive biases [28]. For example, locality, two-dimensional neighborhood structure, and translation equivariance are inherent to CNNs but not applicable to ViTs [28]. Due to this shortcoming, ViTs usually require large-scale training to get competitive performance, such as pre-training on ImageNet-21K [26] and JFT-300M [70]. To alleviate the ViT need for large datasets, DeiT [74] introduced a teacher-student strategy to distill knowledge from a teacher CNN for a student ViT. In the MIM field, MAE [36] uses a masked autoencoder with a lightweight decoder as a visual representation learner. Its learning objective is to reconstruct the original image by the decoder while using masked images as input to the autoencoder. The advantage is that MAE can randomly discard 75% image patches when pre-training under ImageNet-1K [26], which means more efficient training.
5.2 Adversarial Attacks on ViTs
The concept of adversarial attacks first appeared in [23], which proposed a formal framework and algorithms against the adversarial spam detection domain. Then, the adversarial attacks were popularized by Biggio et al. [11] and Szegedy et al. [71] in image classification. The generation of adversarial examples depends on the model’s gradient or estimated gradient in black box situation [43]. Therefore, adversarial attacks can be easily applied to transformers by using the gradient of attention blocks concerning inputs. This also raises a question: are transformers more robust than CNNs? Benz et al. [9] found that CNNs are less robust than ViTs due to their shift-invariant property. Bhojanapalli et al. [10] found that ResNet models are more robust than transformers at the same model size under FGSM attack, but under PGD [55] attack, transformer models show better robustness. As ViTs process the input image as a sequence of patches, Gu et al. [35] found that ViTs are more robust than CNNs to naturally corrupted patches because the attention mechanism is helpful in ignoring naturally corrupted image patches. The later work [6] revealed that CNNs could be as robust as ViTs against adversarial attacks if CNNs are trained with proper hyperparameters.
5.3 Adversarial Defense
Due to the difference between CNNs and ViTs, there have been some recent efforts to explore new adversarial training approaches for ViTs [57, 25, 87]. Mo et al. [57] presented a new adversarial training strategy based on the following observations: 1) pre-training with natural data can provide better robustness after adversarial fine-tuning, 2) gradient clipping is necessary for adversarial training, and 3) using SGD as the optimizer is better than Adam. Debenedetti et al. [25] also presented an improved training strategy for ViTs by evaluating different combinations of data augmentation policies. They found that weak data augmentation and large weight decay are better than previous canonical training approaches. As adversarial training is time-consuming, AGAT [87] leverages the attention score while training to discard non-critical image patches after every layer. Unlike previous works, we provide a different training paradigm by using MIM for adversarial pre-training. Our method is efficient as we discard 75% image patches while pre-training. Our method is effective as we eliminate the information of adversarial perturbation from two information sources of clean and adversarial inputs. We also provide theoretical proof that the information of adversarial perturbation is eliminated.
5.4 Self-Supervised Adversarial Pre-Training
[…]
6 Conclusions and Future Work
This paper investigates MIMIR as a pre-training method to improve adversarial robustness for ViTs under various benchmark datasets. MIMIR uses adversarial examples as inputs and clean data as the reconstruction target. In this way, the information from the adversarial perturbation is removed by the bottleneck, but the information from clean data is maintained due to the constraint while reconstructing the target. We also theoretically prove that MI between adversarial inputs and their latent representations, i.e., $I(x+\delta,z)$, should be decreased while training. Our experimental results show that MIMIR significantly improves adversarial robustness compared to recent state-of-the-art methods. In addition, a single epoch of MIMIR pre-training costs only about $\frac{1}{8}$ of training time compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86] (see Table X). The aforesaid limitations of MIMIR represent a good starting point for future work. For example, we will investigate how to apply MIMIR to CNNs and pyramid-based ViTs and build new self-supervised tasks for adversarial training.
Appendix A Datasets
We use three commonly used datasets to evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26]. CIFAR-10 [46] comprises 50,000 images with size $3\times 32\times 32$ in 10 classes. ImageNet-1K [26] is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes. In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$. For completeness, we also include Tiny-ImageNet [47] as a medium size dataset between CIFAR-10 [46] and ImageNet-1K [26]. Tiny-ImageNet [47] contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
Appendix B Decoder Hyperparameters
We use transformer blocks but fewer layers as the backbone of the decoder. For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Table XI shows the hyperparameters of decoder architectures. For different ViT architectures, we use the transformer blocks of the respective architectures to build the decoder.
TABLE XI: Model architectures of the encoder and decoder.
Model Layers Hidden size MLP ratio Heads
ViT-T (encoder) 12 192 4 3
decoder 2 128 4 16
Appendix C Data Augmentation Ablation Study
Figure 6 demonstrates the loss and accuracy while training with different augmentations. “no mix” refers to using only weak augmentation, including RandomResizedCrop and RandomHorizontalFlip. “+mix” refers to using MixUp (0.8) and CutMix (1.0). “+aug” refers to using MixUp (0.8), CutMix (1.0), and Randaugment (rand-m9-mstd0.5-inc1).
Figure 6: The training results of using different data augmentations with 800 epochs.
Appendix D Details of Training Hyperparameters
In Tables XII and XIII, we provide the default hyperparameters used in our experiments. In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient. In addition, due to the depth and comparatively small embedding size of CaiT, we use different drop path and layer-wise decay at fine-tuning (for ImageNet-1K). For CaiT-XXS24, we use 0.95 and 0.15 as layer-wise decay and dropout, and 0.85 and 0.35 for CaiT-S36. We also apply stochastic depth decay rule [42] to CaiT. CaiT-S36 models are only fine-tuned 50 epochs due to time-consumption and it is sufficient to get superior results. The batch size to fine-tune CaiT is 512 due to limitation of GPU memory. Other hyperparameters are consistent with Tables XII and XIII.
TABLE XII: Pre-training hyperparameters.
Config Value
optimizer AdamW
base learning rate 1.5e-4
weight decay 0.05
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.95$
batch size 512(CIFAR-10, Tiny), 2,048 (ImageNet-1K)
learning rate schedule cosine decay
warmup epochs 40
training epochs 800
augmentation RandomResizedCrop, RandomHorizontalFlip
TABLE XIII: Fine-tuning hyperparameters.
Config Value
optimizer AdamW
base learning rate 0.5e-2 (CIFAR-10), 1e-3 (ImageNet-1K, Tiny)
weight decay 0.05
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
layer-wise lr decay 0.65
batch size 128 (CIFAR-10), 256 (Tiny), 1,024 (ImageNet-1K)
learning rate schedule cosine decay
warmup epochs 10
training epochs 100
augmentation RandomResizedCrop, RandomHorizontalFlip
drop path 0.1
Appendix E Fine-tuning with Clean Images
In Table XIV, we show the results of MIMIR pre-training with adversarial data, and MAE [36] pre-training with clean data. MIMIR shows comparable and even better performance than MAE when fine-tuning with clean data. The results in Table XIV is reported with 800 pre-training epochs. The blr used in Table XIV is 0.001, and we fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K.
TABLE XIV: The experimental results for fine-tuning with clean images.
Model Dataset Pre-train Natural Acc.
ViT-B CIFAR-10 MIMIR 96.91
MAE 96.79
Tiny-ImageNet MIMIR 75.43
MAE 73.38
ImageNet-1K MIMIR 83.20
MAE 82.92
ConViT-S CIFAR-10 MIMIR 95.38
MAE 94.95
Tiny-ImageNet MIMIR 70.40
MAE 69.03
ImageNet-1K MIMIR 79.21
MAE 78.37
Appendix F Dropout is Important for CaiT
We also consider applying our method to deeper ViTs (i.e., CaiT [75]). CaiT is designed for deeper high-capacity transformers that benefit from depth. Simply increasing the depth may cause training failure of ViTs [75]. The CaiT solves this problem by LayerScale and class-attention. In adversarial training, training deeper networks is even more difficult. The dropout has a great impact on robustness. In Table XV, we show the results using MIMIR with different dropout values.
TABLE XV: The experimental results on CaiT-S36. The “sdd” refers to stochastic depth decay rule [42].
Dropout Layer Decay LR $\epsilon=2/255$
Natural PGD
0.1 0.65 72.68 35.60
0.1+SDD 0.65 73.06 38.88
0.2 0.75 75.41 43.99
0.25 0.75 71.46 48.86
0.25 0.85 76.43 53.49
0.25+SDD 0.85 75.92 56.53
0.35+SDD 0.85 76.05 56.78
Appendix G Mutual Information and HSIC
MI measures the mutual dependence between two random variables, $X$ and $Y$. It can be decomposed as:
$I(X,Y)$ $=H(X)-H(X|Y),$ (24)
$=H(Y)-H(Y|X),$
$=H(X)+H(Y)-H(X,Y),$
where $H(X)$ and $H(Y)$ are the information entropies, $H(X|Y)$ and $H(Y|X)$ are the conditional entropies, and $H(X,Y)$ is the joint entropy of $X$ and $Y$.
Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$. To address this issue, the deterministic information bottleneck (DIB) [94] uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ [66, 93], which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
Specifically, given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$ (in our setup, $N$ refers to the mini-batch size), we can obtain two Gram (or kernel) matrices $K_{x}$ and $K_{y}$, for variables $X$ and $Y$, respectively, with $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, in which $\kappa_{x}$ and $\kappa_{y}$ are corresponding kernel functions. The information entropy of $X$ can be expressed as:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log_{2}\left(\text{tr}(\tilde{K_{x}}^{\alpha})\right)=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{N}\lambda_{i}(\tilde{K_{x}})^{\alpha}\right),$ (25)
where $\tilde{K}$ is the normalized version of $K$, i.e., $\tilde{K}=K/{\text{tr}(K)}$, and $\lambda_{i}(\tilde{K})$ denotes the $i$-th eigenvalue of $\tilde{K}$.
Further, the joint entropy for $X$ and $Y$ can be expressed as:
$H_{\alpha}(X,Y)=H_{\alpha}\left(\frac{K_{x}\circ K_{y}}{\text{tr}(K_{x}\circ K_{y})}\right),$ (26)
where $K_{x}\circ K_{y}$ denotes the Hadamard product between the matrices $K_{x}$ and $K_{y}$.
Given Eqs. (25) and (26), the matrix-based Rényi’s $\alpha$-order mutual information $I_{\alpha}(X;Y)$ in analogy of Shannon’s MI is given by:
$I_{\alpha}(X;Y)=H_{\alpha}(X)+H_{\alpha}(Y)-H_{\alpha}(X,Y).$ (27)
Throughout this paper, we use the radial basis function (RBF) kernel $\kappa(x_{i},x_{j})=\exp(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}})$ with kernel width $\sigma$ to obtain the Gram matrices.
The Hilbert–Schmidt Independence Criterion (HSIC) [34] is also a kernel-based dependence measure and is usually used as a surrogate of MI. Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
$\begin{split}&\text{HSIC}_{P_{X,Y}}(X,Y)\\ &=||C_{XY}||^{2}\\ &=\mathbb{E}_{xyx^{\prime}y^{\prime}}[\kappa_{x}(x,x^{\prime})\kappa_{y^{\prime}}(y,y^{\prime})]\\ &+\mathbb{E}_{xx^{\prime}}[\kappa_{x}(x,x^{\prime})]E_{yy^{\prime}}[\kappa_{y}(y,y^{\prime})]\\ &-2\mathbb{E}_{xy}[\mathbb{E}_{x^{\prime}}[\kappa_{x}(x,x^{\prime})]\mathbb{E}_{y^{\prime}}[\kappa_{y}(y,y^{\prime})]],\end{split}$ (28)
where $\kappa_{x}$ and $\kappa_{y}$ are kernel functions, $\mathbb{E}$ is the expectation, $x^{\prime}$ and $y^{\prime}$ are independent copies of $x$ and $y$, respectively.
Given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$, the empirical estimator of HSIC is given by:
$\widehat{\text{HSIC}}_{P_{X,Y}}(X,Y)=\frac{1}{N^{2}}\text{tr}\left(K_{x}HH_{y}H\right),$ (29)
in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathbbm{1}\mathbbm{1}^{T}$ is the centering matrix.

### v3

@@
 Abstract
-Vision Transformers (ViTs) achieve superior performance on various tasks compared to convolutional neural networks (CNNs), but ViTs are also vulnerable to adversarial attacks.
-Adversarial training is one of the most successful methods to build robust CNN models.
-Thus, recent works explored new methodologies for adversarial training of ViTs based on the differences between ViTs and CNNs, such as better training strategies, preventing attention from focusing on a single block, or discarding low-attention embeddings.
-However, these methods still follow the design of traditional supervised adversarial training, limiting the potential of adversarial training on ViTs.
-This paper proposes a novel defense method, MIMIR, which aims to build a different adversarial training methodology by utilizing Masked Image Modeling at pre-training.
-We create an autoencoder that accepts adversarial examples as input but takes the clean examples as the modeling target.
-Then, we create a mutual information (MI) penalty following the idea of the Information Bottleneck.
-Among the two information source inputs and corresponding adversarial perturbation, the perturbation information is eliminated due to the constraint of the modeling target.
-Next, we provide a theoretical analysis of MIMIR using the bounds of the MI penalty.
-We also design two adaptive attacks when the adversary is aware of the MIMIR defense and show that MIMIR still performs well.
-The experimental results show that MIMIR improves (natural and adversarial) accuracy on average by 4.19% on CIFAR-10 and 4.64% on ImageNet-1K, compared to baselines.
-On Tiny-ImageNet, we obtained improved natural accuracy of 2.99% on average and comparable adversarial accuracy.
-Our code and trained models are publicly available [^1] https://github.com/xiaoyunxxy/MIMIR.
-Index Terms:
-Mask Image Modeling, Adversarial Robustness, Vision Transformer, Mutual Information
+Vision Transformers (ViTs) achieve excellent performance in various tasks, but they are also vulnerable to adversarial attacks.
+Building robust ViTs is highly dependent on dedicated Adversarial Training (AT) strategies.
+However, current ViTs’ adversarial training only employs well-established training approaches from convolutional neural network (CNN) training, where pre-training provides the basis for AT fine-tuning with the additional help of tailored data augmentations.
+In this paper, we take a closer look at the adversarial robustness of ViTs by providing a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training.
+Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained by utilizing the MI bounds.
+Based on this finding, we propose a masked autoencoder-based pre-training method, MIMIR, that employs an MI penalty to facilitate the adversarial training of ViTs.
+Extensive experiments show that MIMIR outperforms state-of-the-art adversarially trained ViTs on benchmark datasets with higher natural and robust accuracy, indicating that ViTs can substantially benefit from exploiting MI.
+In addition, we consider two adaptive attacks by assuming that the adversary is aware of the MIMIR design, which further verifies the provided robustness.
+Our code and trained models are publicly available at the anonymous link.
+[^1] https://anonymous.4open.science/r/MIMIR-5444
 1 Introduction
-While Vision Transformers [28] and the variants [51, 24] have achieved great advances, these attention-based technologies fail in being more robust than CNNs against evasion attacks in common threat models [6, 1, 57, 87, 35].
-Evasion attacks [11, 71] (also known as adversarial attacks) refer to fooling well-trained deep models by adding human-imperceptible perturbations to inputs.
-To build robust transformer models, adversarial training is considered one of the most successful methods [87, 57, 25], in which the general idea is to use adversarial examples for training.
-However, adversarial training has limitations due to its expensive computational cost to generate adversarial examples.
-Significant efforts have been made to decrease the adversarial training cost [67, 86, 80, 69], but the nature of attention-based transformer introduces new challenges for accelerating adversarial training.
-For example, transformers lack inductive biases [28], including locality, two-dimensional neighborhood structure, and translation equivariance.
-These biases are inherent to CNNs as prior knowledge but not applicable to ViTs [28].
-The lack of such prior knowledge makes ViTs require a large amount of training data to generalize well, which means generating more adversarial perturbations for adversarial training.
-Due to the differences between CNN and ViTs, previous works try to design new methodologies for adversarial training when considering the robustness of ViTs.
-Gu et al.
-[35] proposed Smoothed Attention to improve the robustness against patch-wise attacks by blocking attention to individual patches.
-Attention-Guided Adversarial Training (AGAT) [87] can efficiently drop image embeddings with lower attention and keep the higher ones after each layer of self-attention, accelerating adversarial training.
-Mo et al.
-[57] found that the robustness of ViTs can be improved by randomly masking perturbations or masking gradients from some attention blocks.
-In addition to new methodologies, some work has also tried to find better training strategies for the robustness of ViTs.
-The authors of [57] observed beneficial behaviors for adversarial training of ViTs, including fine-tuning on clean pre-trained models, using gradient clipping, and combining SGD optimizer and piecewise learning rate scheduler.
-Bai et al.
-[6] discussed that adversarial examples can be too difficult for ViTs to learn when combined with strong data augmentation, such as Randaugment [21], CutMix [95], and MixUp [97].
-Then, they proposed a warmup strategy for augmentations to alleviate this issue.
-They progressively increase the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix.
-Debenedetti et al.
-[25] also found that strong data augmentations harm adversarial training.
-They proposed to use basic data augmentations (random-resize-and-crop + horizontal-flipping + color-jitter) and high weight decay for better performance.
-However, previous research focuses on adversarial training at the fine-tuning stage or lacks rigorous explanations, limiting the potential for improving adversarial training.
-For example, AGAT [87] accelerates adversarial training by dropping a constant ratio of low-attention image embeddings after every layer.
-Overall, they reduce 40% computation of the entire model.
-Designing a method that discards more image embeddings is challenging because the accuracy of clean inputs also decreases when a larger number of embeddings are dropped.
-We investigate a novel approach where we conduct adversarial training at the pre-training stage.
-Our motivation is to discard a much larger ratio of image embeddings at the pre-training stage, further accelerating adversarial training.
-In addition, we show a very intuitive analysis of the training process (i.e., decreasing MI between adversarial inputs and latent representation) with theoretical proof, which ensures a good performance.
-The above motivations yield a self-supervised learning task suitable for Masked Image Modeling (MIM).
-MIM refers to learning from unlabeled data through masked inputs to recover the original information, such as recovering the original visual tokens [7] or original image pixels [36] based on the masked image patches, masking out a portion of the input sequence and then predicting the feature of the masked regions [85].
-Usually, MIM methods can mask out a much larger ratio of inputs by self-supervised pre-training.
-For example, MAE [36] masks out 75% of image patches at the beginning of the forward pass.
-The reason is that MIM intends to learn more discriminative information, so masking out the discriminative part of images and recovering it gives a more difficult reconstructing task [81].
-The reconstructing task is more difficult when applying an adversarial perturbation to the original image, as the perturbation corrupts the natural features in clean images.
-In this way, the model learns more discriminative information, which is also robust to adversarial perturbation.
-Consequently, adversarial training can be both efficient and effective.
-Specifically, we propose Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR) as a pre-training method, where we use adversarial examples as inputs and clean examples as reconstructing objectives.
-As illustrated in Figure 1, MIMIR [^2] MIMIR is a figure in Norse mythology renowned for his knowledge and wisdom.
-consists of an encoder and a decoder following the design of MAE [36].
-The encoder is a ViT, which has the same architecture as the model to be fine-tuned but without the final classification layer.
-The decoder consists of a few lightweight ViT blocks that reconstruct the original clean image from the latent features generated by the encoder.
-At pre-training, adversarial examples are separated into image patches.
-A large ratio of image patches is masked, so the encoder only processes a small part of adversarial patches.
-At fine-tuning, the decoder is discarded.
-The encoder is combined with the initialized classification layer and applied to complete images for classification tasks.
-As we use latent representation from adversarial examples to reconstruct clean images, this represents a more difficult reconstruction task.
-In addition, we theoretically analyze the behaviors of our encoder and decoder following the idea of Information Bottleneck [72].
-The MI between adversarial inputs and latent representation will decrease concerning the accuracy of adversarial examples while pre-training.
-This motivates us to directly add an MI penalty (between adversarial inputs and latent representation) into the loss function of MIMIR at pre-training.
-This new pre-training loss further improves performance at the fine-tuning stage (see Table VII).
-We perform comprehensive experiments on CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26] and compare to state-of-the-art results.
-The results demonstrate that MIMIR significantly improves the performance on both natural and adversarial accuracy.
-Finally, previous works [25, 57, 6] claim that strong data augmentations (Randaugment [21], CutMix [95], and MixUp [97]) are unsuccessful in adversarial training.
-However, we show in Table VI that strong augmentations actually improve the natural and adversarial accuracy with the longer training schedule.
-Our key contributions are summarized as follows:
-  • We propose a novel and efficient defense against adversarial attacks on ViTs - MIMIR, which discards 75% of image patches and then reconstructs the complete clean image under a self-supervised framework.
-  • We provide a theoretical analysis on using adversarial examples and MI penalty in MIMIR. We show that MI between adversarial examples and learned latent representation should be decreased.
-  • We systematically evaluate MIMIR using three datasets and three transformer architectures with multiple model scales under various adversarial attacks.
-  • We show MIMIR is resistant to two adaptive attacks when the adversary is aware of MIMIR.
-First, we build a PGD-feature attack inspired by [65], which directly increases the distance between features extracted from clean and adversarial examples.
-Then, we build a PGD MI attack that uses cross entropy and the MI penalty as learning objectives to find adversarial examples.
-Figure 1: The structure of MIMIR.
-First, we generate adversarial perturbation $\delta$ and apply it to the input image.
-Then, following the original ViT, we use separated image patches as input for the autoencoder.
-We use the decoder output, i.e., $x_{re}$, and the clean input image, i.e., $x$, to calculate the loss.
-Finally, we use MI ($I(x+\delta,z)$) as an additional penalty in the pre-training loss.
-After pre-training, the encoder plus a classification layer is fine-tuned for classification tasks.
-2 Background
-2.1 Adversarial Attack
+Adversarial training (Madry et al.
+2018) is one of the standard techniques to build robust models.
+Regarding adversarial training for ViTs, early works focus on investigating the unique attention mechanism of ViTs.
+There, robustness can be improved by randomly dropping gradient according to attention (Mo et al.
+2022) or improving training efficiency by dropping low-attention image embeddings (Wu et al.
+2022).
+The majority of recent works focus on improving the ViTs’ adversarial training based on experience with CNNs, given AT’s success in building robust CNNs.
+Looking at the adversarial training recipe differences between ViTs and CNNs, strong data augmentation, e.g., Randaugment (Cubuk et al.
+2020), CutMix (Yun et al.
+2019), and MixUp (Zhang et al.
+2018), show positive influence in CNN but are ineffective for ViTs AT and sometimes even reduce the AT performance of ViTs (Bai et al.
+2021; Singh, Croce, and Hein 2023).
+To mitigate this, the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix need to be progressively increased (Bai et al.
+2021).
+Similar problems of strong data augmentations are also observed by (Debenedetti, Sehwag, and Mittal 2023), where basic data augmentations, i.e., random-resize-and-crop, horizontal-flipping, and color-jitter with high weight decay are shown to be promising to improve ViT adversarial training.
+Another common practice in ViT adversarial training is to leverage pre-trained models, which has also been extensively explored (Mo et al.
+2022; Singh, Croce, and Hein 2023).
+Fine-tuning on natural pre-trained models could benefit the adversarial robustness of ViTs.
+For instance, recently, AdvXL (Wang et al.
+2024) introduced an efficient and effective adversarial training strategy for giant models on web-scale datasets.
+However, as the current ViT adversarial training strategies are mainly inherited from CNN AT, the key factor that drives the success of pre-training has not been completely revealed.
+For instance, pre-training with ImageNet-21K and SimMIM (Xie et al.
+2022) performs similarly to the model trained from scratch, and even the CLIP (Radford et al.
+2021) pre-trained model shows a performance drop (Liu et al.
+2024).
+In this work, we aim to provide an in-depth analysis of ViT’s robustness using MI, focusing on self-supervised pre-training.
+Information Bottleneck (IB) has been previously used in supervised learning to analyze adversarial robustness, where a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features is introduced to mitigate the impact of the adversarial noise in the inputs.
+Regarding the task of ViT adversarial training, we provide a novel theoretical analysis for the self-supervised autoencoder, showing that the self-supervised ViT benefits from decreasing the MI between inputs and latent features.
+Based on this finding, we propose a theoretically grounded adversarial pre-training method, Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR).
+Figure 1 provides an illustrative diagram of MIMIR.
+MIMIR substantially improves the adversarial robustness of ViT models as demonstrated by comprehensive experiments on ImageNet-1K (Deng et al.
+2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al.
+2009).
+We also provide an experimental comparison with the state-of-the-art, showing the effectiveness of MIMIR in improving both natural and adversarial prediction accuracy (see Section 4.2).
+Our main contributions are:
+  • By revising the current ViT AT strategies, we point out that ViT adversarial pre-training methods compromise natural accuracy and lack systematic analysis.
+To this end, we provide a theoretical analysis using adversarial examples and the Mutual Information penalty.
+The MI bounds motivate us to decrease MI between adversarial examples and learned latent representation.
+  • Based on the findings of our analysis, we propose an efficient self-supervised defense MIMIR against adversarial attacks on ViTs.
+We systematically evaluate MIMIR using three datasets and multiple model scales under various adversarial attacks, demonstrating its superiority.
+  • We show MIMIR is resistant to two adaptive attacks, where the adversary is aware of MIMIR design.
+We build a PGD-fea attack that increases the distance between features extracted from natural and adversarial examples.
+Moreover, we build a PGD-MI attack that includes the MI penalty as the learning objective to find adversarial examples.
+Figure 1: Diagram illustrating the working mechanism of MIMIR.
+In the pre-training, adversarial perturbations $\delta$ are generated and added to natural images $x$ to create adversarial images $x+\delta$, as shown on the left.
+Patches of generated adversarial images are fed as training data to the ViT-based autoencoder, where the output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss.
+In particular, we propose using MI ($I(x+\delta,z)$) as an additional penalty to compose the pre-training loss of MIMIR, as shown in the bottom right.
+After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final model that is further fine-tuned for classification tasks.
+2 Background and Related Work
 […]
-  1. 1.
-    The generated perturbation $\delta$ can successfully mislead the network by maximizing (e.g., PGD attack [55]):
 $\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}),$ (1)
-[…]
-  2. 2.
-    The generated adversarial examples are as similar as possible to the original clean examples by limiting $\delta$ to a relatively small domain:
+The generated adversarial examples are as similar as possible to the original natural examples by limiting $\delta$ to a relatively small domain:
 […]
@@
 The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{0},l_{2},$ and $l_{\infty}$.
-When using the above attacks to generate adversarial examples for training, the learning objective is:
+When using the above adversarial examples to conduct adversarial training, the learning objective is:
 $\mathop{\min}\limits_{\theta}\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}).$ (3)
-2.2 Masked Image Modeling - MIM
-MIM refers to a self-supervised pre-training framework that aims to reconstruct pre-defined targets, such as discrete tokens [7], raw RGB pixels [36, 49], or features [85].
-The final goal is to use the pre-trained model as a starting point for downstream fine-tuning.
-The downstream tasks include, for instance, classification and object detection.
-More specifically, to build a high-performance ViT $f_{e}$ without a classification layer, we consider $f_{e}$ as an encoder to extract discriminative input features.
-Then, we design a lightweight decoder $f_{d}$, which uses the output of $f_{e}$ as its input.
-The goal of $f_{d}$ is to reconstruct the original inputs (let us consider MAE [36] as an example).
-The aim is to decrease the distance between the input $x$ and $x_{re}=f_{d}\circ f_{e}(x)$.
-After the encoder $f_{e}$ and decoder $f_{d}$ are trained, we use $f_{e}$ plus a manually initialized classification layer as the starting point of fine-tuning.
-2.3 Mutual Information - MI
-MI measures the mutual dependence between two random variables, $X$ and $Y$.
-It quantifies the amount of information contained in one random variable about another random variable or the reduced uncertainty of a random variable when another random variable is known.
-It can be written as:
-$I(X,Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{(X)}(x)P_{(Y)}(y)}\right),$ (4)
-where $P_{(X,Y)}$ is the joint probability density function of $X$ and $Y$.
-$P_{(X)}$ and $P_{(Y)}$ are the marginal probability density function of $X$ and $Y$.
-In addition, MI can be equivalently expressed as:
-$I(X,Y)$ $=H(X)-H(X|Y).$ (5)
-However, quantifying MI is difficult in practice, as it is difficult to precisely estimate $P_{X,Y}$ or $P_{X}$ and $P_{Y}$ in high dimensional space [31].
-DIB [94] suggested using the matrix-based Renyi’s $\alpha$-order entropy $I_{\alpha}$ [66, 93] to estimate MI, which avoids density estimation and variational approximation.
-An alternative way is the Hilbert-Schmidt Independence Criterion (HSIC) [34], which is a kernel-based dependence measure defined in a reproducing kernel Hilbert space (RKHS) and usually used as a surrogate of MI.
-Details about definitions and empirical estimators of $I_{\alpha}$ and HSIC are provided in Appendix G.
-In this paper, we use both $I_{\alpha}$ and HSIC as our MI measurements.
-2.4 Information Bottleneck - IB
-The IB concepts were first proposed in [72] and further developed for deep learning in [73, 68].
-It has been widely used to improve adversarial robustness [90, 2, 94, 84].
-IB describes the generalization of a deep network in two phases: 1) empirical error minimization (ERM) and 2) representation compression [68].
+Information Bottleneck (IB).
+The IB concepts were first proposed in (Tishby, Pereira, and Bialek 2000) and further developed for deep learning in (Tishby and Zaslavsky 2015; Shwartz-Ziv and Tishby 2017).
+It has been widely used to improve adversarial robustness (Xu, Perin, and Picek 2023; Alemi et al.
+2017; Yu, Yu, and Príncipe 2021; Wang et al.
+2021).
+IB describes the generalization of a deep neural network in two phases: 1) empirical error minimization (ERM) and 2) representation compression.
 For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer.
 The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$.
-The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I(\cdot)$.
-During neural network training, in the ERM phase, the model increases shared information between $t_{l}$ with respect to both $x$ and $y$.
-Afterward, in the compression phase, the model decreases information contained in $t_{l}$ about $x$ but preserves (or even increases) information about $y$.
-The reduction of $I(x,t_{l})$ can be interpreted as a way of reducing noise or compressing irrelevant or redundant features in $x$.
-At the end of the training, the model strikes a trade-off that maximizes $I(y,t_{l})$ and minimizes $I(x,t_{l})$.
-Formally, the IB minimizes the following Lagrangian:
-$\mathcal{L}=I(x,t_{l})-\beta I(y,t_{l}),$ (6)
-where $\beta$ is a Lagrange multiplier that controls the trade-off between predicting $y$ and compressing $x$.
+The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I$.
+Adversarial Robustness of ViTs.
+PGD (Madry et al.
+2018) adversarial training is considered one of the most effective defenses that can withstand adaptively designed attacks (Athalye, Carlini, and Wagner 2018).
+Due to the difference between CNNs and ViTs, there have been substantial recent efforts to explore new adversarial training approaches for ViTs.
+EasyRobust (Mao et al.
+2022) evaluates the robustness of ViTs using PGD adversarial training.
+(Mo et al.
+2022; Bai et al.
+2021; Debenedetti, Sehwag, and Mittal 2023) demonstrate the classical training recipe for CNN is not suitable for ViTs and present new adversarial training strategies for ViTs by evaluating different hyperparameters and data augmentation policies.
+(Mo et al.
+2022) also finds that pre-training with natural data provides better robustness after adversarial fine-tuning.
+(Singh, Croce, and Hein 2023) explores the effect of strong pre-training and replacing the patch embedding in ViTs with convolutional blocks.
+AdvXL (Wang et al.
+2024) shows an outstanding performance of giant models and web-scale data with adversarial pre-training and fine-tuning.
+However, (Liu et al.
+2024) provides a different view that pre-training with ImageNet-21K and SimMIM (Xie et al.
+2022) performs similarly to the model trained from scratch, and the CLIP pre-trained model has a performance drop.
+Unlike existing works, we provide a formal and in-depth analysis leveraging MI, based on which we introduce a new adversarial pre-training scheme.
 3 MIMIR
 3.1 Threat Model
-We assume the adversary intends to create adversarial examples from clean images.
-The adversary has the strongest ability, i.e., the white box, to access the target model.
-Then, adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets.
-In addition, the adversary is also aware of potential defenses.
-An adversary can design new attacks for a specific model according to the details of our defense method.
-From the defender’s perspective, the main goal is to train a robust model against potential adversarial attacks. The defender considers the following four objectives:
-  • The performance on clean data should not decrease significantly. We allow a small drop of clean accuracy for a trade-off in robustness.
-  • The accuracy of models trained by MIMIR should be significantly higher than for models without defenses under various adversarial attacks.
-  • Deploying MIMIR should be efficient and scalable to large datasets such as ImageNet-1K [26].
-  • MIMIR should be robust against adaptive attacks.
+Adversary’s Goal.
+The attacker aims to fool the trained model by decreasing the overall classification accuracy (non-targeted) or forcing the model to recognize an input as a specific target (targeted).
+The image perturbations are limited by the $l_{\infty}$ norm for invisibility.
+Adversary Knowledge.
+The attacker has white-box access to the model, including training data, architectures, hyperparameters, and model weights.
+The attacker can implement iterative attacks and unlimited queries to update adversarial examples multiple times in white and black-box settings.
+Adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets.
+The adversary can adaptively design new attacks for a specific model according to the design details of the defense method.
+Defender’s Goal.
+The defender’s main goal is to train a robust model against potential adversarial attacks and prevent the performance of natural data from decreasing significantly.
+In addition, the defense method should be robust against adaptive attacks specifically designed for it.
 3.2 Design Intuition
-MIM has been proved useful as a pre-training method for ViTs on various tasks [79, 36, 85, 7].
-To train a powerful model by MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model.
+MIM has been proved useful as a pre-training method for ViTs on various tasks (Vincent et al.
+2010; He et al.
+2022; Wei et al.
+2022; Bao et al.
+2022).
+To train a powerful model using MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model.
 Masking out the foreground instead of the background reduces discriminative information in visible information to the model.
-Reconstructing foreground parts is harder than background and helps the model learn more discriminative information [81].
-Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbation to clean inputs.
-The adversarial perturbation increases the distance between the input and reconstruction target.
-If we can reconstruct clean inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks.
-In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial inputs while enabling the decoder to reconstruct the clean image.
-Note that simply masking out all foreground does not make sense, as it would not be possible to reconstruct meaningful content.
+Reconstructing foreground parts is harder than background and helps the model learn more discriminative information (Wang et al.
+2023).
+Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbations to natural inputs.
+The adversarial perturbations increase the distance between the input and reconstruction target.
+If we can reconstruct natural inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks.
+In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial perturbations while enabling the decoder to reconstruct the natural image.
+Note that simply masking out all foreground does not make sense, as it would render the reconstruction of meaningful content infeasible.
 From the perspective of IB, there is a bottleneck between the encoder and decoder.
-Our adversarial pre-training task contains two information sources: the clean data $x$ and the adversarial perturbation $\delta$.
+Our adversarial pre-training task contains two information sources: the natural data $x$ and the adversarial perturbations $\delta$.
 As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated.
 The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.
-3.3 Design Details
-Autoencoder.
-MIMIR follows the general design of MAE [36], which consists of an encoder $f_{e}$ and a decoder $f_{d}$.
-Like all other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$.
-The decoder reconstructs original inputs according to the discriminative features.
-Following ViT [28], the input $x$ is separated into non-overlapping image patches.
-Then, we randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder.
-This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it is not possible to find the reconstruction target then.
-Therefore, we hope to keep a part of the foreground as a hint for reconstruction.
-The information of masked content is recorded as mask tokens $m$, which are not used by the encoder but will be used later by the decoder.
-Each token is a learned vector that indicates the presence of a masked patch to be predicted.
-The mask token is shared by all inputs of the decoder.
-Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image.
-We emphasize that mask tokens are not used in the encoder part.
-To train a ViT, we use the same transformer blocks as ViT to build the encoder.
-The encoder only processes the visible patches, making training much more efficient.
-When converting to other architectures, such as ConVit [24], we use corresponding transformer blocks to build the encoder.
-The decoder accepts the encoded visible image patches and mask tokens as inputs.
-The decoder is built using the same transformer blocks as the encoder instead of using ViT [28] transformer blocks for all.
-Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
-Adversarial pre-training target.
-The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder.
-Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results.
-Following the original MAE [36], this distance is measured by Mean Squared Error (MSE).
-To create a more difficult reconstruction task, we apply adversarial perturbation $\delta$ on the inputs of the encoder.
-Thus, the adversarial perturbation is also masked along with the image when inputted into the encoder.
-The decoder reconstructs the original clean inputs by using the latent features $z$ extracted from adversarial examples.
-The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model.
-Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets.
-Formally, the pre-training process (described in Algorithm 1) can be written as follows:
-$z$ $=f_{e}(x+\delta),$ (7)
-$x_{re}=f_{d}(z),$
-$loss_{\text{mse}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re}).$
-MI as penalty.
-Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases with respect to the accuracy on adversarial example, i.e., $I(x+\delta,z)$ is decreasing while training.
-Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
-$loss_{\text{mi}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re})+\lambda I(x+\delta,z),$ (8)
-where $\lambda$ is a regularizer for the MI penalty.
-We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty.
-This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$.
-According to Data Processing Inequality (DPI) [8], $I(x,z)\leq I(x+\delta,z)$.
-Minimizing $I(x,z)$ in loss equals minimizing a lower bound of $I(x+\delta,z)$.
-Algorithm 1 Masked Adversarial Pre-Training
+Algorithm 1 MIMIR Pre-training
 0:  training data $D$, number of epochs $E$, encoder $f_{e}$, decoder $f_{d}$, network parameters $\theta$, $\mathcal{L}_{\text{mse}}$, $\lambda$.
@@
 12:  end for
+3.3 Design Details
+Autoencoder.
+MIMIR follows the general design of MAE (He et al.
+2022), consisting of an encoder $f_{e}$ and a decoder $f_{d}$.
+As with other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$.
+The decoder reconstructs original inputs according to the discriminative features.
+Following ViT (Dosovitskiy et al.
+2021), the input $x$ is separated into non-overlapping image patches.
+We randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder.
+This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it becomes infeasible to find the reconstruction target.
+Thus, we aim to keep a part of the foreground as a hint for reconstruction.
+The information of masked content is recorded as mask tokens $m$, not used by the encoder but reserved for later use by the decoder.
+Each token is a learned vector indicating the presence of a masked patch to be predicted.
+The mask token is shared by all inputs of the decoder.
+Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image.
+We emphasize that mask tokens are not used in the encoder part.
+To train a ViT, we use the same transformer blocks as ViT to build the encoder.
+The encoder only processes the visible patches, making training much more efficient.
+When converting to other architectures, such as ConViT (D’Ascoli et al.
+2021), we use corresponding transformer blocks to build the encoder.
+The decoder accepts the encoded visible image patches and mask tokens as inputs.
+The decoder is built using the same transformer blocks as the encoder instead of using ViT (Dosovitskiy et al.
+2021) transformer blocks for all.
+Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
+Adversarial Pre-training Target.
+The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder.
+Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results.
+Following the original MAE (He et al.
+2022), this distance is measured by Mean Squared Error (MSE).
+To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ on the inputs of the encoder.
+Thus, the adversarial perturbations are also masked along with the image upon input into the encoder.
+The decoder reconstructs the original natural inputs by using the latent features $z$ extracted from adversarial examples.
+The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model.
+Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets.
+Formally, the pre-training process (described in Algorithm 1) can be written as follows:
+$z=f_{e}(x+\delta),~{}x_{re}=f_{d}(z),$ (4)
+$loss_{\text{mse}}=\mathcal{L}_{\text{mse}}(x,x_{re}).$
+MI as Penalty.
+Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$ is decreasing while training.
+Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
+$loss_{\text{mi}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re})+\lambda I(x+\delta,z),$ (5)
+where $\lambda$ is a regularizer for the MI penalty.
+We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty.
+This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$.
+According to Data Processing Inequality (DPI) (Beaudry and Renner 2011), $I(x,z)\leq I(x+\delta,z)$.
+$I(x+\delta,z)$ is closer to $z$ on the Markov chain.
 Generating Adversarial Examples.
-To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbation $\delta$.
-As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD [55].
-Still, we can easily design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq.
+To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbations $\delta$.
+As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD (Madry et al.
+2018).
+Nevertheless, it is feasible to design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq.
 (4).
 As the feature $z$ is extracted from only visible image patches, we only attack the visible patches.
-We do not add any perturbation to mask tokens since the outputs of the autoencoder are only impacted by visible patches.
+We do not add any perturbations to mask tokens since the outputs of the autoencoder are only impacted by visible patches.
 Then, the adversarial pre-training learning objective can be written as:
 $\begin{split}\mathcal{L}_{\text{adv}}=\mathop{\max}\limits_{\delta\in S}&\mathcal{L}_{\text{mse}}(f_{d}\circ f_{e}(x+\delta),x),\\ \mathop{\min}\limits_{\theta}\mathcal{L}_{\text{adv}}&+\lambda I(x+\delta,z),\end{split}$ (6)
-where $\theta$ are the parameters of the autoencoder.
-Adversarial fine-tuning.
+where $\theta$ are the autoencoder parameters.
 After the autoencoder is trained, we discard the decoder and initialize a classification layer for the encoder to build a complete model.
-We use simple adversarial training, PGD [55], to fine-tune our pre-trained models so that we can ensure the improvements come from the MIMIR pre-training.
-The complete model is trained at the fine-tuning stage.
+Finally, the complete model is fine-tuned by AT methods.
+Figure 2: The example plots for the lower and upper bounds on the MI in Propositions 3.2 and 3.3.
+The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes.
+The lower bound reaches its minimum at $p_{e}=0.9$.
 3.4 Theoretical Justification
-In this section, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained.
-Let $F$ be any classifier trained on clean samples with desirable prediction accuracy, which may suffer from adversarial attacks.
+Next, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained.
+Let $F$ denote any classifier trained on natural samples with desirable prediction accuracy, which may suffer from adversarial attacks.
 We begin our analysis by first presenting Lemma 3.1.
-Lemma 3.1
-Given an arbitrary classifier $F$, let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$. We have:
+Lemma 3.1.
+Let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$, we have:
 $I(F(x+\delta),F(x_{re}))\leq I(F(x+\delta),x_{re})\leq I(x+\delta,z).$ (7)
-Proof:
+Proof.
 There are two Markov chains:
@@
 $z\rightarrow x_{re}\rightarrow F(x_{re}),$
-which indicate that $F(x+\delta)$ is an indirect observation of $x+\delta$, whereas both $F(x_{re})$ and $x_{re}$ are indirect observations of $z$.
+which implies that $F(x+\delta)$ is an indirect observation of $x+\delta$, whereas both $F(x_{re})$ and $x_{re}$ are indirect observations of $z$.
 By the data processing inequality (DPI), we have
@@
 Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$.
-Intuitively, our autoencoder is trained to only recover clean sample $x$ without any interference from $\delta$.
+Intuitively, our autoencoder is trained to recover only natural sample $x$ without any interference from $\delta$.
 Hence, a relatively large value of $p_{e}$ is expected.
 In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
-Proposition 3.2
+[…]
+Architecture Params (M) FT FT Epoch Adv. Steps Source Natural Robust
+ResNet-50 25.0 PGD 100 1 (Bai et al. 2021) 67.44 35.54
+DeiT-S 22.1 PGD 100 1 (Bai et al. 2021) 66.62 36.56
+DeiT-S 22.1 PGD 110 1 (Debenedetti, Sehwag, and Mittal 2023) 66.80 37.90
+ViT-S 22.1 PGD 120 3 (Mao et al. 2022) 66.43 39.20
+ViT-S 22.1 PGD 300 3 (Liu et al. 2024) 70.7 43.7
+RobArch-S 26.1 PGD 110 3 (Peng et al. 2023) 70.17 44.14
+ViT-S 22.1 APGD 300 2 (Singh, Croce, and Hein 2023) 69.22 44.04
+ViT-S 22.1 APGD 300 2 Ours 71.00 46.10
+ViT-B 86.6 ARD+PRM 10 5 (Mo et al. 2022) 69.10 34.62
+ViT-B 86.6 PGD 120 3 (Mao et al. 2022) 70.64 43.04
+RobArch-L 104 PGD 100 3 (Peng et al. 2023) 73.44 48.94
+ViT-B 86.6 PGD 300 3 (Liu et al. 2024) 74.7 49.7
+[…]
+ViT-B 86.6 APGD 300 2 (Singh, Croce, and Hein 2023) 74.10 50.30
+ViT-B 86.6 APGD 300 2 Ours 76.32 54.28
+Proposition 3.2.
 Let $H(\cdot)$ denote the information entropy and $H_{b}(p_{e})=-p_{e}\log_{2}p_{e}-(1-p_{e})\log_{2}(1-p_{e})$ be the binary entropy, we have:
 $H(F(x+\delta))-H_{b}(p_{e})-p_{e}\log(|F(x+\delta)|-1)\leq I(x+\delta,z),$ (11)
-where $|F(x+\delta)|$ is the total number of categories [^3] For instance, for CIFAR-10 [46], $|F(x+\delta)|=10$..
-Proof:
+where $|F(x+\delta)|$ is the total number of categories. [^2] For instance, for CIFAR-10, $|F(x+\delta)|=10$.
+Proof.
 By the chain rule of MI, we have
 $I(F(x+\delta),F(x_{re}))=H(F(x+\delta))-H(F(x+\delta)|F(x_{re})).$ (12)
-By applying Fano’s inequality [30, 58], we obtain:
+By applying Fano’s inequality (Fano 1949; Ocal et al. 2019), we obtain:
 $H(F(x+\delta)|F(x_{re}))\leq H_{b}(p_{e})+p_{e}\log(|F(x+\delta)|-1).$ (13)
@@
 If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$.
-Figure 2: The example plots for the lower and upper bounds on the MI in Propositions 3.2 and 3.3.
-The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes.
-The lower bound reaches its minimum at $p_{e}=0.9$.
-In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (11).
-Proposition 3.3
+In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq.
+(8).
+Proposition 3.3.
 If $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, we have:
@@
 in which the notation $\lesssim$” refers to less than or similar to.
-Proof:
-By the Hellman-Raviv inequality [38, 13], we have:
+Proof.
+By the Hellman-Raviv inequality (Hellman and Raviv 1970; Brown 2009), we have:
 $2p_{e}$ $\leq H(F(x+\delta)|x_{re})$ (16)
@@
 In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$.
-For simplicity, let us assume the clean data $x$ and adversarial perturbation $\delta$ are independent, then:
+For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^3] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations (Moosavi-Dezfooli et al.
+2017)., then:
 $I(x+\delta,z)=I(x,z)+I(\delta,z).$ (17)
-According to [78], minimizing the expected reconstruction error between clean sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input.
+According to (Vincent et al.
+2008), minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input.
 Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized).
-In other words, only the information about $\delta$ has been removed from $z$.
+In other words, only the adversarial information about $\delta$ has been removed from $z$ when minimizing $I(x+\delta,z)$.
 This also explains the robustness of $z$.
-TABLE I: Different learning rates. Fine-tuned for 50 epochs.
+4 Experiments
+4.1 Experimental Setup
+We evaluate MIMIR on three datasets: ImageNet-1K (Deng et al.
+2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al.
+2009).
+Details of datasets are provided in Appendix A.
+Hyperparameters of the decoder are included in Appendix B.
+Training Setup.
+We train models from scratch for all experiments.
+Following MAE (He et al.
+2022), we do pre-training by MIMIR for 800 epochs.
+The number of warmup epochs is 40 for pre-training.
+We use AdamW (Loshchilov and Hutter 2019) as an optimizer for both pre-training and fine-tuning.
+The cosine decay as the learning rate scheduler is applied.
+At pre-training, MIMIR uses the 1-step PGD to generate adversarial examples for all three datasets.
+The perturbation budget is $\epsilon=8,\alpha=10$.
+At fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet.
+For ImageNet-1K, we use 1-step PGD with random initialization for better efficiency.
+The perturbation bound is $\epsilon=4$ for ImageNet-1K.
+We also use a 2-step APGD with a longer fine-tuning scheduler to compare with SOTA works in Table 1.
+Details on training hyperparameters are provided in Appendix C.
+Evaluation Metrics.
+We use natural accuracy and robust accuracy as evaluation metrics.
+Natural accuracy refers to the accuracy of natural and unmodified inputs.
+The robust accuracy measures the accuracy under the AutoAttack (Francesco and Hein 2020).
+AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks.
+In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce (Francesco and Hein 2020), APGD-t (Francesco and Hein 2020), FAB-t (Croce and Hein 2020), and Square (Andriushchenko et al.
+2020).
+The perturbation budgets are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K.
+4.2 Experimental Results
+Comparison with SOTA Works.
+Table 1 compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$).
+We report the number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy.
+The robust accuracy is evaluated by AutoAttack on the RobustBench validation set (5000 images).
+We divide the models into two groups: small ($\approx 22$M) and large ($\approx 86$M) models.
+In addition, (Singh, Croce, and Hein 2023) shows that the Convolutional stem (ConvStem) improves the robustness of ViTs.
+We also include an evaluation of ViTs+ConvStem in Table 2.
+ConvStem (Xiao et al.
+2021) refers to replacing the patch embedding in ViTs with a convolutional block.
+MIMIR outperforms all SOTA works.
+Table 2: Influence of the ConvStem block.
+The Standard pre-trained weights refer to pre-trained weights from timm.
+The configurations are deit_small_patch16_224 and vit_base_patch16_224.augreg_in1k.
+The weights are fine-tuned 300 epochs by 2-step APGD.
+†The result on ViT-S is taken from (Singh, Croce, and Hein 2023).
+Architecture Pre-trained weights Natural Robust
+ViT-S+ConvStem Standard 72.56 48.08†
+MIMIR 72.72 48.44
+ViT-B+ConvStem Standard 75.26 52.70
+MIMIR 76.72 54.04
+Ablation Study.
+Table 3 provides an ablation study to verify the design choices of MIMIR.
+The ablation uses 100 epochs of 1-step PGD (PGD1) AT as the baseline.
+Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library) as initialization of AT.
+After that, we replace the clean pre-training with MAE, adv MAE, and MIMIR step by step.
+The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in loss.
+The pre-training scheduler is 800 epochs.
+We also use stronger adversarial fine-tuning for better performance, including 2-step PGD (PGD2), APGD (APGD2) FT, and a longer fine-tuning scheduler (300 epochs).
+Please note that a longer training scheduler does not guarantee better results (Mo et al.
+2022).
+Our results indicate that MIMIR outperforms baselines and can be further improved under the long training scheduler.
+In addition, MIMIR is also efficient and increases less than 5% of time consumption on ImageNet-1K compared to MAE, which is shown in Table 12 in the Appendix E.
 […]
-CIFAR-10 ViT-T 5.0e-4 76.30 47.60
-1.0e-3 80.69 49.56
-1.0e-2 85.62 48.78
-5.0e-2 85.12 50.30
-1.0e-1 84.51 50.40
-TABLE II: Top 1 natural and adversarial accuracy on CIFAR-10 test set.
-The results of ARD+PRM [57] are taken from the original paper.
-As there is no ViT-T architecture, we use the same architecture as DeiT-T.
+Architecture Training Recipe Natural Robust
+ViT-S PGD1 FT w/o PT 66.02 31.40
+clean PT + PGD1 FT 67.04 33.70
+MAE PT + PGD1 FT 69.98 35.64
 […]
-CIFAR-10 [46] ($l_{\infty},\epsilon=8/255$) ViT-T (DeiT-T) ImageNet-1K PGD 75.46 48.10 47.96 43.62
-MAE [36] PGD 79.89 48.64 48.43 44.27
-ImageNet-1K ARD+PRM [57] 79.60 50.33 50.15 45.99
-MIMIR PGD 84.82 57.51 57.38 52.96
-ViT-S ImageNet-1K PGD 79.59 50.86 50.73 46.37
-MAE [36] PGD 85.97 54.19 53.94 50.49
-ImageNet-1K ARD+PRM [57] 81.86 51.73 51.46 47.33
-MIMIR PGD 88.11 56.63 56.34 53.18
-ViT-B ImageNet-1K PGD 83.16 52.98 52.71 49.06
-MAE [36] PGD 88.86 57.03 56.86 52.42
-ImageNet-1K ARD+PRM [57] 84.90 53.80 53.51 50.03
-MIMIR PGD 89.30 58.07 57.87 54.55
-ConViT-T ImageNet-1K PGD 53.09 33.63 33.61 29.65
-MAE [36] PGD 75.49 45.29 45.17 41.12
-ImageNet-1K ARD+PRM [57] 80.28 49.86 49.55 45.42
-MIMIR PGD 80.74 49.37 49.16 45.04
-ConViT-S ImageNet-1K PGD 54.03 34.61 34.60 30.60
-MAE [36] PGD 86.43 54.76 54.59 51.13
-ImageNet-1K ARD+PRM [57] 84.32 53.10 52.81 48.85
-MIMIR PGD 87.49 56.35 56.20 52.54
-ConViT-B ImageNet-1K PGD 61.54 38.77 38.71 34.21
-MAE [36] PGD 89.03 58.19 57.75 54.25
-ImageNet-1K ARD+PRM [57] 85.80 53.36 53.05 49.33
-MIMIR PGD 89.30 59.31 59.01 55.64
-4 Experimental Evaluation
-4.1 Experimental Setup
-We use PyTorch [60] 2.1.0 for implementation, and mixed precision [56] to accelerate training.
-We use three datasets to comprehensively evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26], details of datasets are given in Appendix A.
-Appendix B provides details about decoder hyperparameters.
-Model Architecture: We evaluate MIMIR with three commonly used architectures with multiple scales: ViT [28], ConViT [24], and XCiT [3].
-Hardware. Experiments are performed on a single machine with 4 RTX A6000 (48GB) and 4 RTX A5000 (24GB) GPUs, CUDA 12.0.
-Training hyperparameters.
-For all experiments, we train models from scratch (unless otherwise specified).
-Following MAE [36], we do pre-training by MIMIR for 800 epochs and then fine-tuning for 100 epochs.
-The number of warmup epochs is 40 and 10 for pre-training and fine-tuning, respectively.
-We use AdamW [53] as an optimizer for both pre-training and fine-tuning.
-We use cosine decay as the learning rate scheduler.
-At pre-training, one-step FGSM with random initialization achieves better performance than using ten-step PGD in Table VII.
-Therefore, we use a one-step algorithm to generate adversarial examples for all three datasets.
-The perturbation budget is $\epsilon=8,\alpha=10$.
-At fine-tuning stage, we use a 10-step PGD algorithm to generate adversarial examples for training to be aligned with baselines.
-The perturbation limit is $\epsilon=8,\alpha=2$.
-When fine-tuning with ImageNet-1K, we use one-step FGSM with random initialization to align with baselines and for better efficiency.
-Following previous research [86, 87], we fine-tuned twice for two different perturbation limits, $\epsilon=2,\epsilon=4$.
-As discussed in Section 4.3, strong data augmentation is harmful to adversarial training at a common training schedule (50 or 100 epochs for fine-tuning).
-We only use simple augmentations, including random resized crops and random horizontal flips.
-Previous research used rather different fine-tuning learning rates, so we chose the learning rate by testing different values as given in Table I.
-We use the best, i.e., $25.0e-$ 2, as the base learning rate (blr) for CIFAR-10.
-The learning rate is calculated based on blr, i.e., $lr=blr\times\frac{batchsize}{256}$.
-For Tiny-ImageNet and ImageNet-1K, we use the same blr as MAE [36], i.e., $0.001$.
-At the pre-training stage, we use $41.5e-$ 4 as the learning rate for all datasets.
-In addition, we noticed that it is possible to get NaN (Not a Number) for loss in a few cases while adversarial training, i.e., the training fails.
-We solve this problem by using a learning rate that is ten times smaller.
-More details about hyperparameters can be found in the Appendix D.
-TABLE III: Top 1 natural and adversarial accuracy on ImageNet-1K.
-Dataset Models Pre-train Fine-tune $\epsilon=2/255$ $\epsilon=4/255$
-[…]
-ImageNet-1K [26] ($l_{\infty}$) ViT-S - FastAT [86] 71.37 49.49 49.41 66.92 34.07 33.82
-- AGAT [87] 70.62 49.00 48.85 66.10 33.62 33.40
-MIMIR FastAT [86] 74.60 54.56 54.55 71.29 40.98 40.63
-ViT-B - FastAT [86] 70.31 50.55 50.06 65.18 33.59 33.39
-- AGAT [87] 70.41 51.23 51.11 67.93 34.94 34.78
-MIMIR FastAT [86] 75.88 55.42 55.36 73.22 41.26 40.46
-CaiT-XXS24 - FastAT [86] 72.84 54.31 54.26 68.77 32.14 32.09
-- AGAT [87] 71.15 54.17 54.08 68.22 31.46 31.39
-MIMIR FastAT [86] 73.39 53.39 53.37 69.90 40.53 40.22
-CaiT-S36 - FastAT [86] 72.51 53.12 52.76 71.20 33.02 32.84
-- AGAT [87] 72.69 53.66 53.48 71.06 33.46 33.17
-MIMIR FastAT [86] 76.05 56.78 56.75 73.57 40.03 39.16
-TABLE IV: Comparison of different architectures under 100-step PGD attack.
-Dataset Models $\epsilon=2/255$ $\epsilon=4/255$
-Natural PGD Natural PGD
-ImageNet-1K ($l_{\infty}$) ViT-S 74.60 54.56 71.29 40.63
-ConViT-S 74.72 55.65 72.08 42.62
-XCiT-S 74.39 55.21 71.77 40.97
-XCiT-M 73.63 53.04 72.01 42.39
-[…]
-[…]
-Tiny-ImageNet [47] ($l_{\infty},\epsilon=8/255$) WRN-28-10 Gowal et al. [33] $\times$ 512 400 51.56 N/A N/A 21.56
-WRN-28-10 Wang et at. [83] 1m 512 400 53.62 N/A N/A 23.40
-[…]
-[…]
-ViT-S MIMIR 1m 512/256 800/100 63.33 26.37 26.17 22.52
-ViT-B MIMIR 1m 512/256 800/100 65.06 25.41 25.03 22.82
-ConViT-T MIMIR 1m 512/256 800/100 52.61 24.61 24.58 19.42
-ConViT-S MIMIR 1m 512/256 800/100 62.29 26.39 26.30 22.88
-Evaluation metrics.
-We use natural accuracy and adversarial accuracy as evaluation metrics.
-Natural accuracy refers to the accuracy of a model on clean and unmodified inputs.
-The adversarial accuracy measures the accuracy of a model under a specific adversarial attack.
-To show that the improvement is not from obfuscated gradient [5], we use PGD [55] and AutoAttack [20] to evaluate the robustness of trained models.
-PGD is one of the most popular adversarial attacks used for generating adversarial examples.
-We evaluate MIMIR under 20 and 100 steps of PGD attack.
-AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks.
-In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce [20], APGD-t [20], FAB-t [19], and Square [4].
-We use the same perturbation budget for our experiments to compare with previous works.
-We apply the commonly-used $l_{\infty}$ norm for PGD [55] and AutoAttack (AA) [20].
-We set the perturbation radius $\epsilon$ to $8/255$ for CIFAR-10 [46] and Tiny-ImageNet [47].
-The step length is $\alpha=2/255$.
-As the image size of ImageNet-1K is much larger, we use different scales of $\epsilon$ for adversarial training and evaluation [26].
-Following previous works [57, 87, 50], we use two perturbation radius $2/255,4/255$ for evaluation.
-The step length is $\alpha=1/255$.
-4.2 Experimental Results
-Overall performance.
-We apply MIMIR to CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26].
-The results are shown in Tables IIIII, and V, respectively.
-Overall, MIMIR significantly improves adversarial accuracy under PGD and AA attacks.
-While training with CIFAR-10 on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.50% compared to the baseline.
-On ConViTs (ConViT-T, ConViT-S, and ConViT-B), MIMIR improves the accuracy by an average of 2.87% compared to the baseline.
-While using ImageNet-1K on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.52% compared to the baseline.
-When fine-tuning with clean data, MIMIR also shows improved natural accuracy in Table XIV.
-In addition, we also evaluate the time consumption and memory usage of MIMIR in Table X.
-We compare MIMIR with FastAT [86] and 10-step PGD adversarial training.
-The results show that using MIMIR for pre-training is much more efficient than traditional adversarial training.
-[…]
-Comparison with related work.
-We present the performances of MIMIR compared with recent baseline methods in Tables II and III.
-In Table II, we perform experiments using CIFAR-10 [46].
-The results are compared with three baselines, including pre-training on ImageNet-1K and fine-tuning with PGD, MAE [36], and ARD+PRM [57].
-The “ImageNet-1K” under the Pre-train column refers to using models pre-trained on ImageNet-1K, then fine-tuned with PGD for 40 epochs, which is the first baseline in ARD+PRM [57].
-The MAE [36] baseline is pre-trained using clean CIFAR-10 images and fine-tuned with 10-step PGD adversarial training.
-The training schedule of MAE [36] baseline is 800 epochs for pre-training and 100 epochs for fine-tuning.
-Our MIMIR uses the same training schedule as MAE [36].
-We provide results on fine-tuning with clean images in Appendix E.
-The ARD+PRM [57] baseline refers to the best algorithm in [57], where we use the results from that work.
-Note that ARD+PRM [57] baseline uses models pre-trained on ImageNet-1K.
-Using models pre-trained on ImageNet-1K and then fine-tuning for CIFAR-10 can easily get accuracy higher than 95% [75].
-As MIMIR only uses CIFAR-10 for training, it is more challenging to achieve good performance.
-For a more fair comparison, we use elucidating diffusion model (EDM) data as a data augmentation since the baselines also use EDM data (Wang et at.
-[83] and Rebuffi et al.
-[62]) or models pre-trained on ImageNet-1K (ARD+PRM [57]).
-Generative data is usually used to improve adversarial training [83, 61, 63, 33].
-Specifically, we use 5 million generated CIFAR-10 data and 1 million Tiny-ImageNet data provided by [83].
-Note that the EDM data is generated only based on CIFAR-10 or Tiny-ImageNet, and the MAE baseline uses EDM data as well.
-The EDM data is applied to experiments with CIFAR-10 and Tiny-ImageNet but not to ImageNet-1K.
-According to results in Table II, MIMIR consistently improves adversarial robustness compared to all baselines.
-It also improves the natural accuracy compared to the MAE baseline.
-This is because self-supervised MIMIR does not suffer from inconsistency between the ground truth and adversarial example [59].
-Let us consider a joint data distribution $p_{d}(x,y)$ (training data) and a discriminative model $p_{\theta}(y|x)$, where $x$ is the data with its label $y$.
-In standard training, we hope the model learns discriminative features by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x)$.
-In adversarial training, we hope the model is robust by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x+\delta)$, i.e., Eq.
-(3).
-This means $p_{\theta}(y|x+\delta)$, instead of $p_{\theta}(y|x)$, is encouraged to converge to $p_{d}(y|x)$.
-Supervised adversarial training builds a shortcut from $x+\delta$ to $y$.
-For the optimal $\theta^{*}$ of minimizing Eq.
-(3), $p_{\theta^{*}}(y|x)\neq p_{d}(y|x)$ [59].
-Therefore, there is always a trade-off between natural and adversarial accuracy in supervised adversarial training.
-However, such inconsistency does not apply to self-supervised MIMIR, as there is no label for self-supervised training.
-MIMIR encourages $p_{\theta_{ae}}(x_{re}|x+\delta)$ to converge to $p_{d}(x)$, which does no affect convergence of $p_{\theta_{ae}}(x_{re}|x)$ to converge to $p_{d}(x)$.
-This implies that supervised pre-training is harmful to natural accuracy, which is also aligned with the findings of previous research [39].
-Table III shows the results evaluated on the ImageNet-1K [26] dataset.
-The performance of MIMIR is compared with two baselines, including FastAT [86] and AGAT [87].
-The FastAT [86] is designed for efficient adversarial training so that it can be applied to large models and large datasets, such as ImageNet-1K.
-The main idea of FastAT is to use a single-step adversarial attack to generate adversarial examples.
-The authors refined the standard FGSM [32] by applying a random start at the beginning of the algorithm and a much larger step size to reach the perturbation limit.
-According to empirical results, they used $\alpha=1.25\times\epsilon$ for step size, which provided comparable performance to the best-reported result from free adversarial training [25].
-The results of FastAT [86] and AGAT [87] are taken from [87].
-The baselines of FastAT [86] and AGAT [87] are trained for 300 epochs with a cosine decay learning rate schedule, and the first 20 epochs for learning rate warmup.
-As shown in Table III, the accuracy of AGAT [87] may decrease compared to FastAT [86].
-This is because AGAT [87] discards low-attention image patches for efficiency.
-The image patches, even with the low-attention scores, are useful for classification.
-In Table IV, we compare the performance among three architectures with multiple scales.
-MIMIR shows similar performance on other architectures as on the original ViTs, which means MIMIR is well-generalized to different architectures.
-Figure 3: The loss and natural accuracy for three different augmentation strategies, using CIFAR-10 and ViT-small.
-As CIFAR-10 is small compared to ImageNet-1K, we also provide experiments on Tiny-ImageNet for completeness.
-We use the same training schedule and hyperparameters for experiments on Tiny-ImageNet.
-As AGAT [87] and ARD+PRM [57] do not provide experiments on Tiny-ImageNet, we compare our results with the most recent adversarial robustness results from Wang et at.
-[83] and Rebuffi et al.
-[62].
-They use generated data as augmentation for more robust adversarial training.
-Specifically, Gowal et al.
-[33] study four methods for generated data as augmentation, including Denoising Diffusion Probabilistic Model (DDPM) [41], Very Deep Variational Auto-Encoder (VDVAE) [18], StyleGAN [45], and BigGAN [12].
-Wang et at.
-[83] demonstrated better diffusion models and further improved adversarial robustness compared to previous works.
-In Table V, we show our experimental results compared with previous works.
-Our method provides competitive performance.
-Especially when the baselines only use generated data based on Tiny-ImageNet, MIMIR provides the best natural accuracy (65.06%).
-When the baselines use generated data based on ImageNet-1K, and we only use that based on Tiny-ImageNet, MIMIR still shows comparable performance.
-In addition, according to Table X, the MIMIR training schedule (800 pre-training + 100 fine-tuning) takes approximately 200 fine-tuning (10-step PGD) epochs of training time.
-Thus, MIMIR costs much less training time and gets comparable results to baselines.
-TABLE VI: Data augmentation ablation with longer fine-tuning schedule.
-[…]
-ViT-B 800 Weak Augmentation 89.90 60.26
-+ CutMix [95],MixUp [97] 91.01 60.62
-+ Randaugment [21] 90.19 62.75
-4.3 Data Augmentation is Harmful
-As mentioned in previous works [6, 25], strong data augmentation makes the training samples too difficult to be learned by ViTs while conducting adversarial training.
-The strong data augmentation refers to the combination of Randaugment [21], CutMix [95], and MixUp [97].
-In this section, we evaluate two different solutions to ease this problem.
-First, we only use simple data augmentation for adversarial training, including random crop (or random resize crop for ImageNet-1K) and random horizontal flip (“weak aug”).
-Second, we use a 10-epoch warmup procedure for strong data augmentation.
-The warmup of Randaugment is implemented by progressively increasing the distortion magnitude from 1 to 9 (“warmup aug”).
-For CutMix and MixUp, we warm up by increasing the mixup probability from 0.5 to 1.0.
-As shown in Figure 3, “weak aug” provides the best accuracy.
-The “warmup aug” shows a slightly improved accuracy compared to fusing strong augmentation.
-Therefore, we provide a different result from [6] on the smaller dataset CIFAR-10, i.e., we show that weak augmentation is better than warmup augmentation.
-Even data augmentation with reduced amplitude is still difficult to learn at the beginning of adversarial training.
-Although strong augmentation is harmful to a normal training schedule, we show in Table VI that CutMix [95], MixUp [97], and Randaugment [21] increase the accuracy of adversarial training when training with a longer schedule, e.g., 800 epochs of fine-tuning.
-We conjecture that the data with strong augmentation is helpful but difficult for adversarial training to learn.
-Thus, it needs more epochs to learn meaningful representation.
-Loss and accuracy curves can be found in Appendix C.
-Due to the longer schedule, we use patch size 4 to reduce training time.
-TABLE VII: Comparison between HSIC and $I_{\alpha}$. Models are pre-trained for 800 epochs. For fine-tuning, all models are adversarially trained with 10-step PGD for 50 epochs.
+MIMIR PT + PGD1 FT 71.02 37.22
+MIMIR PT + PGD2 FT 70.78 38.16
+MIMIR PT + APGD2 FT 68.78 42.86
+100 $\rightarrow$ 300 epochs FT 71.00 46.10
+Table 4: Comparison between HSIC (Gretton et al.
+2005) and $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) using ViT-T on CIFAR-10.
+Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
 Pre-train $\lambda$ Estimator Natural PGD
-MIMIR 0.001 HSIC [34] 69.63 43.17
-MIMIR 0.001 $I_{\alpha}$ [94] 75.00 46.11
-MIMIR 1e-05 HSIC [34] 76.30 47.60
-MIMIR 1e-05 $I_{\alpha}$ [94] 75.53 46.75
-MIMIR 1e-06 HSIC [34] 74.90 46.19
-MIMIR 1e-06 $I_{\alpha}$ [94] 74.60 45.66
-Adv-MAE (1-step) 0.0 - 74.69 46.28
-Adv-MAE (10-step) 0.0 - 73.96 45.77
+MIMIR 0.001 HSIC 69.63 43.17
+MIMIR 0.001 $I_{\alpha}$ 75.00 46.11
+MIMIR 1e-05 HSIC 76.30 47.60
+MIMIR 1e-05 $I_{\alpha}$ 75.53 46.75
+MIMIR 1e-06 HSIC 74.90 46.19
+MIMIR 1e-06 $I_{\alpha}$ 74.60 45.66
+Table 5: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
+Pre-train $\lambda$ Estimator Natural PGD
 MAE 0.0 - 69.02 42.31
-4.4 MI Measures
+adv MAE (1-step) 0.0 - 74.69 46.28
+adv MAE (10-step) 0.0 - 73.96 45.77
+MIMIR 1e-05 HSIC 76.30 47.60
+MI Measure.
 In Section 3.4, we provide lower and upper bound (Eq.
 (11)) of $I(x+\delta,z)$.
-According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the clean image $x$.
+According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$.
 This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective.
-Before achieving the goal, we must choose a suitable and efficient MI estimator and a regularizer ($\lambda$ in Eq.
-(8)).
-In this paper, we use $I_{\alpha}$ [94] and HSIC [54] as estimators.
-In Table VII, we show the performance of $I_{\alpha}$ [94] and HSIC [54] with different values of $\lambda$.
-According to the results, using HSIC as an estimator with $\lambda=1e-05$ shows the best performance, which is used for other experiments.
-We also show that MIMIR outperforms original MAE [36] and Adv-MAE [91].
-To make it clearer, the MAE baseline in Table VII refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD.
-Adv-MAE refers to using adversarial examples but not using $I(x+\delta,z)$ in loss.
-Adv-MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training.
-Adv-MAE provides better accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task.
-This more difficult task further improves the performance of downstream models.
-We use the default learning rate (i.e., $45.0e-$ 4) of MAE [36] in experiments in Table VII.
-This is aligned with Adv-MAE [91].
-Figure 4: The fine-tuning results with different numbers of pre-train epochs.
-The left column is trained on CIFAR-10, and the right column is on Tiny-ImageNet.
-The above is natural accuracy, and the below is adversarial accuracy under a 20-step PGD attack.
-4.5 Training Epochs Ablation Study
-Our experiments so far are based on 800-epoch pre-training.
-In Figure 4, we show the influence of different numbers of epochs.
-We use ViT-T and ViT-S as our baseline models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet.
-We fine-tune each model 50 epochs after MIMIR pre-training.
-Both adversarial and natural accuracy are improved with a longer training schedule.
-On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S.
-On Tiny-ImageNet, the increase is more obvious.
-4.6 Fine-tuning Using Advanced Methods
-Our fine-tuning thus far is based on traditional 10-step PGD adversarial training for CIFAR-10.
-The reason is to compare with previous works that use a 10-step PGD algorithm.
-Next, we consider using TRADES [96] and MART [82] to study whether MIMIR can achieve better performance with better fine-tuning methods.
-TRADES [96] identify a trade-off between robustness and accuracy on clean data.
-The authors decompose the prediction error of adversarial examples as natural and boundary errors.
-According to these two kinds of errors, the authors find a differentiable upper bound based on the theory of classification-calibrated loss.
-MART [82] finds that misclassified and correctly classified natural examples have different influences on the final robustness after adversarial training.
-Therefore, the authors designed MART to explicitly differentiate the misclassified and correctly classified natural examples during adversarial training.
-In Table VIII, the MIMIR performance could be further improved with the two advanced adversarial training methods.
-Specifically, TRADES slightly improves the accuracy on clean examples but shows worse robustness compared to PGD.
-MART improves the adversarial accuracy when using ViT-B, but we notice a significant drop in natural accuracy.
-TABLE VIII: The experimental results with different fine-tuning methods.
-[…]
-ViT-S PGD 88.11 56.63 56.34 53.18
-TRADES 88.19 56.42 56.36 51.70
-MART 80.55 50.81 50.75 39.92
-ViT-B PGD 89.30 58.07 57.87 54.55
-TRADES 90.91 57.73 57.58 52.79
-MART 85.78 60.35 60.16 54.27
-4.7 Adaptive Attack Evaluation
-As the adversary is aware of the defense MIMIR, they can design adaptive attacks [76].
-For example, the adversary may attack feature space [52, 65] since MIMIR trains backbone to extract features.
-Here, the backbone refers to the ViT model without the final classification layer, i.e., the encoder of MIMIR.
-In this section, we provide two adaptive attacks specifically designed against MIMIR.
-First, we introduce PGD-MI, which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in loss.
-Therefore, it is reasonable to attack the model by directly increasing the MI $I(x+\delta,z)$.
-Formally, we embed the following loss into the PGD algorithm:
-$\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(x_{i}+\delta,y_{i})+\lambda I(x+\delta,z).$ (21)
-We use the same value of $\lambda$ as in MIMIR pre-training.
-Second, we introduce PGD-feature that directly attacks the feature extracted by ViT backbones.
-As the backbones are trained by self-supervised MIMIR pre-training (the whole model is trained when fine-tuning), we hope to evaluate the feature extracted by the backbones.
-The PGD-feature increases the distance between features extracted from clean and adversarial examples.
-We implement it using the $L_{2}$ norm and the PGD algorithm:
-$\mathop{\max}\limits_{\delta\in S}l_{2}$ $(f_{e}(x),f_{e}(x+\delta)).$ (22)
-Table IX shows the adaptive evaluation results for PGD-MI and PGD-feature.
-We use 100 steps for both to make sure the attacking algorithm converges.
-The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$.
-For ImageNet-1K, we use $\epsilon=4/255$.
-According to the results in Table IX, MIMIR is still robust against adaptive attacks.
-4.8 Evaluation of the Loss Surface
-To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot loss surfaces [48] in Figure 5.
-The loss surface is the visualization of the loss function as parameters change.
-The basic idea is to plot the loss around the optimal parameters.
-Formally, we consider in the 2D case:
-$f_{l}(\alpha,\beta)=L(\theta^{*}+\alpha\theta_{1}+\beta\theta_{2}),$ (23)
-where $\theta_{1}$ and $\theta_{2}$ are two direction vectors, $\alpha$ and $\beta$ are two arguments of $f_{l}$.
-In practice, we use the parameters of trained models, i.e., $\theta^{*}$.
-In Figure 5, we provide the loss surfaces for models.
-The surfaces of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
-Figure 5: The loss surfaces of the models.
-The four figures of each column are the loss information of one model.
-The first column is ConViT-S pre-trained with MIMIR and fine-tuned with adversarial examples.
-The second column is ConViT-S, pre-trained with clean data and fine-tuned with adversarial examples.
-The third is ViT-S, pre-trained with clean data and then fine-tuned with adversarial examples.
-The fourth one is ViT-S, pre-trained with MIMIR and fine-tuned with adversarial examples.
-The training dataset is CIFAR-10.
-TABLE IX: The experimental results for adaptive attacks.
-[…]
+In this paper, we use $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) and HSIC (Ma, Lewis, and Kleijn 2020) as estimators (detailed definitions in Appendix F).
+Table 4 demonstrates the performance of different values of $\lambda$.
+According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.
+1-step is better than 10-step of AT in pre-training.
+We also show that MIMIR outperforms original MAE (He et al.
+2022) and adv MAE with different PGD steps (to generate adversarial examples for training).
+MAE in Table 5 refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD.
+The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in loss.
+The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training.
+The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task.
+This more difficult task further improves the performance of downstream models (See also Table 7).
+We use the default learning rate (i.e., $45.0e-$ 4) of MAE, so there is a performance drop in experiments in Table 5 and 4 since AT prefers larger learning rates on CIFAR-10 as shown in Table 11 in the Appendix.
+Stability.
+Due to the high computation cost, we cannot report standard deviation for all experiments.
+To show that our method MIMIR has low variances, we train ViT-S on CIFAR-10 three times with different random seeds, running pre-training 400 epochs and fine-tuning 50 epochs.
+The natural accuracy is 86.07 $\pm$ 0.16 %.
+The adversarial accuracy under PGD20 ($l_{\infty},\epsilon=8/255$) is 47.24 $\pm$ 0.12 %.
+Table 6: Adversarial accuracy by adaptive attacks. The models are pre-trained 800 epochs by MIMIR and fine-tuned 100 epochs by PGD AT.
+Dataset Model PGD20 PGD-MI100 PGD-fea100
 CIFAR-10 ConViT-S 56.35 56.16 78.52
@@
 ViT-B 55.41 55.36 73.51
-4.9 Efficiency
-Table X provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning.
-The time consumption and memory usage are evaluated on 4 A6000 GPUs.
-Compared to 10-step PGD and FAST AT, MIMIR shows much less time consumption.
-At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$.
-Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time.
-We also consider using 10-step PGD at pre-training, Table VII shows that using a one-step algorithm with larger $\alpha=10$ provides better performance.
-In addition, pre-training with MIMIR also costs less memory.
-Note in Table X that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning.
-In Table X, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86].
-Therefore, MIMIR is more promising concerning training costs when conducting adversarial training.
-4.10 Discussion and Limitations
-Across our experiments, we have observed the promising performance of MIMIR.
-This is because we build a different adversarial training paradigm.
-With MIMIR, the adversarial pre-training can be more efficient and effective.
-Following the idea of IB, we can intuitively consider a bottleneck between the encoder and decoder.
-As the reconstruction output is constrained by clean data $x$, the bottleneck will filter out information from adversarial perturbation $\delta$.
-We provide a theoretical guarantee of this bottleneck in Section 3.4.
-In Eq.
-(8), we embed this bottleneck as a learning object to further improve the performance.
-This improvement also confirms the correctness of our theoretical guarantee in Section 3.4.
-Table VII shows that our method works better than baseline even without embedding the bottleneck in Eq.
-(8).
-With the two information sources of $x$ and $\delta$, the model is trained to learn the robust features from $x$ and forget the information of $\delta$ under the constraint of the reconstruction target.
-While MIMIR shows better performance, there are still certain limitations.
-MIMIR is a pre-training method.
-An adversarial fine-tuning is necessary to build the final model.
-Thus, the shortcomings of traditional adversarial training cannot be completely avoided.
-In our experiments, we utilize the simple PGD algorithm for fine-tuning, but one can further improve MIMIR pre-trained models with more advanced approaches.
-In addition, MIMIR follows the design of MAE, and we also utilize the characteristic that ViTs are capable of processing variable-length inputs.
-Therefore, currently, MIMIR cannot handle pyramid-based ViTs and CNNs.
-While it is not trivial to apply MIMIR to other pyramid-based ViTs and CNNs, this is feasible [89, 49, 81].
-TABLE X: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
-CIFAR-10 [46] Tiny-ImageNet [47] ImageNet-1K [26]
-Models #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
-[…]
-FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
-MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
-MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
-[…]
-FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
-MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
-MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
-[…]
-FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
-MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
-MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
-5 Related Work
-5.1 Vision Transformer
-The transformers [77] were first proposed in natural language processing (NLP).
-With the mechanism of global self-attention, transformers can effectively capture the non-local relationships among all text tokens [27, 14, 22].
-The pioneering work, ViT [28], demonstrated that the pure transformer architecture can achieve competitive performance on various tasks.
-ViT also reveals that transformers lack inductive biases [28].
-For example, locality, two-dimensional neighborhood structure, and translation equivariance are inherent to CNNs but not applicable to ViTs [28].
-Due to this shortcoming, ViTs usually require large-scale training to get competitive performance, such as pre-training on ImageNet-21K [26] and JFT-300M [70].
-To alleviate the ViT need for large datasets, DeiT [74] introduced a teacher-student strategy to distill knowledge from a teacher CNN for a student ViT.
-In the MIM field, MAE [36] uses a masked autoencoder with a lightweight decoder as a visual representation learner.
-Its learning objective is to reconstruct the original image by the decoder while using masked images as input to the autoencoder.
-The advantage is that MAE can randomly discard 75% image patches when pre-training under ImageNet-1K [26], which means more efficient training.
-5.2 Adversarial Attacks on ViTs
-The concept of adversarial attacks first appeared in [23], which proposed a formal framework and algorithms against the adversarial spam detection domain.
-Then, the adversarial attacks were popularized by Biggio et al.
-[11] and Szegedy et al.
-[71] in image classification.
-The generation of adversarial examples depends on the model’s gradient or estimated gradient in black box situation [43].
-Therefore, adversarial attacks can be easily applied to transformers by using the gradient of attention blocks concerning inputs.
-This also raises a question: are transformers more robust than CNNs?
-Benz et al.
-[9] found that CNNs are less robust than ViTs due to their shift-invariant property.
-Bhojanapalli et al.
-[10] found that ResNet models are more robust than transformers at the same model size under FGSM attack, but under PGD [55] attack, transformer models show better robustness.
-As ViTs process the input image as a sequence of patches, Gu et al.
-[35] found that ViTs are more robust than CNNs to naturally corrupted patches because the attention mechanism is helpful in ignoring naturally corrupted image patches.
-The later work [6] revealed that CNNs could be as robust as ViTs against adversarial attacks if CNNs are trained with proper hyperparameters.
-5.3 Adversarial Defense
-Due to the difference between CNNs and ViTs, there have been some recent efforts to explore new adversarial training approaches for ViTs [57, 25, 87].
-Mo et al.
-[57] presented a new adversarial training strategy based on the following observations: 1) pre-training with natural data can provide better robustness after adversarial fine-tuning, 2) gradient clipping is necessary for adversarial training, and 3) using SGD as the optimizer is better than Adam.
-Debenedetti et al.
-[25] also presented an improved training strategy for ViTs by evaluating different combinations of data augmentation policies.
-They found that weak data augmentation and large weight decay are better than previous canonical training approaches.
-As adversarial training is time-consuming, AGAT [87] leverages the attention score while training to discard non-critical image patches after every layer.
-Unlike previous works, we provide a different training paradigm by using MIM for adversarial pre-training.
-Our method is efficient as we discard 75% image patches while pre-training.
-Our method is effective as we eliminate the information of adversarial perturbation from two information sources of clean and adversarial inputs.
-We also provide theoretical proof that the information of adversarial perturbation is eliminated.
-5.4 Self-Supervised Adversarial Pre-Training
-[…]
+Adaptive Attacks.
+We evaluate MIMIR against adaptive adversaries following common practices (Tramer et al.
+2020).
+We provide two adaptive attacks specifically designed against MIMIR.
+First, we introduce the PGD Mutual Information attack (PGD-MI), which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in the loss.
+PGD-MI attacks the model by directly increasing the MI $I(x+\delta,z)$.
+Specifically, we add the MI loss into the PGD algorithm:
+$\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(x_{i}+\delta,y_{i})+\lambda I(x+\delta,z).$ (18)
+where the value of $\lambda$ in MIMIR pre-training is available to adversaries.
+Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following (Sabour et al.
+2016).
+In particular, we attack the feature extractor from the backbones after the adversarial fine-tuning.
+The PGD-fea attack increases the Euclidean distance between features extracted from natural and adversarial examples.
+We implement it using the PGD algorithm:
+$\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{\text{mse}}$ $(f_{e}(x),f_{e}(x+\delta)).$ (19)
+Both PGD-MI and PGD-fea are optimized for 100 steps to make sure the attacking algorithm converges.
+The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$ for CIFAR-10 and Tiny-ImageNet, and $\epsilon=4/255$ for ImageNet-1K.
+Table 6 demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks, showing that MIMIR is robust against adaptive attacks on benchmark datasets.
+Table 7: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
+Architecture Pre-train CIFAR-10 Tiny ImageNet
+ViT-B MAE 96.79 73.38 82.92
+MIMIR 96.91 75.43 83.20
+ConViT-S MAE 94.95 69.03 78.37
+MIMIR 95.38 70.40 79.21
+ViT-S MAE 95.95 70.00 77.45
+MIMIR 95.95 71.14 78.69
+Figure 3: The loss landscapes of MIMIR pre-trained models.
+Visualization of the Loss Landscape.
+To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape (Li et al.
+2018) in Figure 3.
+The loss landscape is the visualization of the loss function as parameters change.
+The basic idea is to plot the loss around the optimal parameters.
+The landscapes of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
+Fine-tuning with Natural Images.
+In Table 7, we show the results of MIMIR and the original MAE (He et al.
+2022) with the same hyperparameter.
+We fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K.
+Surprisingly, MIMIR outperforms MAE when fine-tuning with natural data.
+We leave further exploration of this finding to future work.
 5 Conclusions and Future Work
-This paper investigates MIMIR as a pre-training method to improve adversarial robustness for ViTs under various benchmark datasets.
-MIMIR uses adversarial examples as inputs and clean data as the reconstruction target.
-In this way, the information from the adversarial perturbation is removed by the bottleneck, but the information from clean data is maintained due to the constraint while reconstructing the target.
-We also theoretically prove that MI between adversarial inputs and their latent representations, i.e., $I(x+\delta,z)$, should be decreased while training.
-Our experimental results show that MIMIR significantly improves adversarial robustness compared to recent state-of-the-art methods.
-In addition, a single epoch of MIMIR pre-training costs only about $\frac{1}{8}$ of training time compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86] (see Table X).
-The aforesaid limitations of MIMIR represent a good starting point for future work.
-For example, we will investigate how to apply MIMIR to CNNs and pyramid-based ViTs and build new self-supervised tasks for adversarial training.
+This paper provides a novel theoretical MI analysis on ViT adversarial training, showing that MI between the adversarial example and its latent representation in ViT-based autoencoders, i.e., $I(x+\delta,z)$, should be constrained by utilizing the MI bounds.
+Based on this finding, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs.
+MIMIR uses adversarial examples as inputs and natural data as the reconstruction target.
+In this way, the information from the adversarial perturbations is decreased by the bottleneck, while the information from natural data is preserved while reconstructing the target.
+Our experimental results show that MIMIR substantially improves adversarial robustness compared to recent related works on various benchmark datasets.
 Appendix A Datasets
-We use three commonly used datasets to evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26].
-CIFAR-10 [46] comprises 50,000 images with size $3\times 32\times 32$ in 10 classes.
-ImageNet-1K [26] is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes.
+We use three commonly used datasets to evaluate MIMIR: CIFAR-10 (Krizhevsky, Hinton et al.
+2009), Tiny-ImageNet (Le and Yang 2015), and ImageNet-1K (Deng et al.
+2009).
+CIFAR-10 (Krizhevsky, Hinton et al.
+2009) comprises 50,000 images with size $3\times 32\times 32$ in 10 classes.
+ImageNet-1K (Deng et al.
+2009) is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes.
 In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$.
-For completeness, we also include Tiny-ImageNet [47] as a medium size dataset between CIFAR-10 [46] and ImageNet-1K [26].
-Tiny-ImageNet [47] contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
+For completeness, we also include Tiny-ImageNet (Le and Yang 2015) as a medium size dataset between CIFAR-10 (Krizhevsky, Hinton et al.
+2009) and ImageNet-1K (Deng et al.
+2009).
+Tiny-ImageNet (Le and Yang 2015) contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
 Appendix B Decoder Hyperparameters
@@
 For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K.
-Table XI shows the hyperparameters of decoder architectures.
-For different ViT architectures, we use the transformer blocks of the respective architectures to build the decoder.
-TABLE XI: Model architectures of the encoder and decoder.
+Table 8 shows the hyperparameters of decoder architectures.
+For different ViT architectures, we use the transformer blocks of the respective architectures to build the encoder.
+Table 8: Model architectures of the encoder and decoder.
 Model Layers Hidden size MLP ratio Heads
@@
 decoder 2 128 4 16
-Appendix C Data Augmentation Ablation Study
-Figure 6 demonstrates the loss and accuracy while training with different augmentations.
-“no mix” refers to using only weak augmentation, including RandomResizedCrop and RandomHorizontalFlip.
-“+mix” refers to using MixUp (0.8) and CutMix (1.0).
-“+aug” refers to using MixUp (0.8), CutMix (1.0), and Randaugment (rand-m9-mstd0.5-inc1).
-Figure 6: The training results of using different data augmentations with 800 epochs.
-Appendix D Details of Training Hyperparameters
-In Tables XII and XIII, we provide the default hyperparameters used in our experiments.
+Appendix C Details of Training Hyperparameters
+In Tables 9 and 10, we provide the default hyperparameters used in our experiments.
 In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K.
 Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient.
-In addition, due to the depth and comparatively small embedding size of CaiT, we use different drop path and layer-wise decay at fine-tuning (for ImageNet-1K).
-For CaiT-XXS24, we use 0.95 and 0.15 as layer-wise decay and dropout, and 0.85 and 0.35 for CaiT-S36.
-We also apply stochastic depth decay rule [42] to CaiT.
-CaiT-S36 models are only fine-tuned 50 epochs due to time-consumption and it is sufficient to get superior results.
-The batch size to fine-tune CaiT is 512 due to limitation of GPU memory.
-Other hyperparameters are consistent with Tables XII and XIII.
-TABLE XII: Pre-training hyperparameters.
+Table 9: Pre-training hyperparameters.
 Config Value
@@
 augmentation RandomResizedCrop, RandomHorizontalFlip
-TABLE XIII: Fine-tuning hyperparameters.
+Table 10: Fine-tuning hyperparameters.
 Config Value
@@
 optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
-layer-wise lr decay 0.65
+layer-wise lr decay 0.65, 1.0
 batch size 128 (CIFAR-10), 256 (Tiny), 1,024 (ImageNet-1K)
 learning rate schedule cosine decay
-warmup epochs 10
-training epochs 100
+warmup epochs 10, 20
+training epochs 100, 300
 augmentation RandomResizedCrop, RandomHorizontalFlip
 drop path 0.1
-Appendix E Fine-tuning with Clean Images
-In Table XIV, we show the results of MIMIR pre-training with adversarial data, and MAE [36] pre-training with clean data.
-MIMIR shows comparable and even better performance than MAE when fine-tuning with clean data.
-The results in Table XIV is reported with 800 pre-training epochs.
-The blr used in Table XIV is 0.001, and we fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K.
-TABLE XIV: The experimental results for fine-tuning with clean images.
-Model Dataset Pre-train Natural Acc.
-ViT-B CIFAR-10 MIMIR 96.91
-MAE 96.79
-Tiny-ImageNet MIMIR 75.43
-MAE 73.38
-ImageNet-1K MIMIR 83.20
-MAE 82.92
-ConViT-S CIFAR-10 MIMIR 95.38
-MAE 94.95
-Tiny-ImageNet MIMIR 70.40
-MAE 69.03
-ImageNet-1K MIMIR 79.21
-MAE 78.37
-Appendix F Dropout is Important for CaiT
-We also consider applying our method to deeper ViTs (i.e., CaiT [75]).
-CaiT is designed for deeper high-capacity transformers that benefit from depth.
-Simply increasing the depth may cause training failure of ViTs [75].
-The CaiT solves this problem by LayerScale and class-attention.
-In adversarial training, training deeper networks is even more difficult.
-The dropout has a great impact on robustness.
-In Table XV, we show the results using MIMIR with different dropout values.
-TABLE XV: The experimental results on CaiT-S36. The “sdd” refers to stochastic depth decay rule [42].
-Dropout Layer Decay LR $\epsilon=2/255$
-Natural PGD
-0.1 0.65 72.68 35.60
-0.1+SDD 0.65 73.06 38.88
-0.2 0.75 75.41 43.99
-0.25 0.75 71.46 48.86
-0.25 0.85 76.43 53.49
-0.25+SDD 0.85 75.92 56.53
-0.35+SDD 0.85 76.05 56.78
-Appendix G Mutual Information and HSIC
+Table 11: Different learning rates. Fine-tuned for 50 epochs.
+Dataset Models LR Natural PGD10
+CIFAR-10 ViT-T 5.0e-4 76.30 47.60
+1.0e-3 80.69 49.56
+1.0e-2 85.62 48.78
+5.0e-2 85.12 50.30
+1.0e-1 84.51 50.40
+Hardware. Experiments are performed on a single machine with 4 RTX A6000 (48GB) and 4 RTX A5000 (24GB) GPUs, CUDA 12.0.
+Figure 4: Natural and adversarial accuracy of ViT-S and ViT-B that are fine-tuned with different numbers of pre-train epochs under a 20-step PGD attack.
+Appendix D Training Epochs Evaluation Study
+Pre-training epoch.
+Our experiments so far are based on 800-epoch pre-training.
+In Figure 4, we show the influence of different numbers of epochs.
+We use ViT-T and ViT-S as the models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet.
+We fine-tune each model 50 epochs after MIMIR pre-training.
+Both adversarial and natural accuracy are improved with a longer training schedule.
+On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S.
+On Tiny-ImageNet, the increase is more obvious.
+Therefore, a longer pre-training schedule obviously increases performance.
+Fine-tuning epoch.
+To show that the performance of the final model is largely attributed to MIMIR pre-training, we do a very short fine-tuning for pre-trained models.
+This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training.
+Therefore, we evaluate the pre-training performance.
+In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning.
+Figure 5: Natural and adversarial accuracy of ViT-S adversarially fine-tuned for 5 or 10 epochs on CIFAR-10, Tiny-ImageNet, and ImageNet-1K.
+Table 12: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
+CIFAR-10 Tiny-ImageNet ImageNet-1K
+Architecture #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
+ViT-S 21.34 PGD10 AT 149.3 2.54$\times$4 306.0 3.99$\times$4 2251.7 12.5$\times$4
+FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
+MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
+MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
+ViT-B 85.27 PGD10 AT 361.2 5.39$\times$4 1022.2 8.30$\times$4 5416.7 22.1$\times$4
+FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
+MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
+MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
+ConViT-S 27.05 PGD10 AT 442.5 6.64$\times$4 897.0 12.19$\times$4 6626.5 32.5$\times$4
+FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
+MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
+MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
+Appendix E Efficiency
+Table 12 provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning.
+The time consumption and memory usage are evaluated on 4 A6000 GPUs.
+Compared to 10-step PGD and FastAT, MIMIR shows much less time consumption.
+At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$.
+Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time.
+We also consider using a 10-step PGD at pre-training.
+Table 5 shows that using a one-step algorithm with larger $\alpha=10$ provides better performance.
+In addition, pre-training with MIMIR also costs less memory.
+Note in Table 12 that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning.
+In Table 12, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training (Wong, Rice, and Kolter 2020).
+Therefore, MIMIR is more promising regarding training costs when conducting adversarial training.
+Appendix F Mutual Information and HSIC
 MI measures the mutual dependence between two random variables, $X$ and $Y$. It can be decomposed as:
@@
 Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$.
-To address this issue, the deterministic information bottleneck (DIB) [94] uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ [66, 93], which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
+To address this issue, the deterministic information bottleneck (DIB) (Yu, Yu, and Príncipe 2021) uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ (Sanchez Giraldo, Rao, and Principe 2015; Yu et al.
+2019), which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
 Specifically, given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$ (in our setup, $N$ refers to the mini-batch size), we can obtain two Gram (or kernel) matrices $K_{x}$ and $K_{y}$, for variables $X$ and $Y$, respectively, with $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, in which $\kappa_{x}$ and $\kappa_{y}$ are corresponding kernel functions.
@@
 Throughout this paper, we use the radial basis function (RBF) kernel $\kappa(x_{i},x_{j})=\exp(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}})$ with kernel width $\sigma$ to obtain the Gram matrices.
-The Hilbert–Schmidt Independence Criterion (HSIC) [34] is also a kernel-based dependence measure and is usually used as a surrogate of MI.
+The Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al.
+2005) is also a kernel-based dependence measure and is usually used as a surrogate of MI.
 Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
@@
 $\widehat{\text{HSIC}}_{P_{X,Y}}(X,Y)=\frac{1}{N^{2}}\text{tr}\left(K_{x}HH_{y}H\right),$ (25)
-in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathbbm{1}\mathbbm{1}^{T}$ is the centering matrix.
+in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathds{1}\mathds{1}^{T}$ is the centering matrix.

### v5

@@
 Abstract
-Vision Transformers (ViTs) achieve excellent performance in various tasks, but they are also vulnerable to adversarial attacks.
-Building robust ViTs is highly dependent on dedicated Adversarial Training (AT) strategies.
-However, current ViTs’ adversarial training only employs well-established training approaches from convolutional neural network (CNN) training, where pre-training provides the basis for AT fine-tuning with the additional help of tailored data augmentations.
-In this paper, we take a closer look at the adversarial robustness of ViTs by providing a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training.
-Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained by utilizing the MI bounds.
-Based on this finding, we propose a masked autoencoder-based pre-training method, MIMIR, that employs an MI penalty to facilitate the adversarial training of ViTs.
-Extensive experiments show that MIMIR outperforms state-of-the-art adversarially trained ViTs on benchmark datasets with higher natural and robust accuracy, indicating that ViTs can substantially benefit from exploiting MI.
-In addition, we consider two adaptive attacks by assuming that the adversary is aware of the MIMIR design, which further verifies the provided robustness.
-Our code and trained models are publicly available at the anonymous link.
-[^1] https://anonymous.4open.science/r/MIMIR-5444
-1 Introduction
-Adversarial training (Madry et al.
-2018) is one of the standard techniques to build robust models.
-Regarding adversarial training for ViTs, early works focus on investigating the unique attention mechanism of ViTs.
-There, robustness can be improved by randomly dropping gradient according to attention (Mo et al.
-2022) or improving training efficiency by dropping low-attention image embeddings (Wu et al.
-2022).
-The majority of recent works focus on improving the ViTs’ adversarial training based on experience with CNNs, given AT’s success in building robust CNNs.
-Looking at the adversarial training recipe differences between ViTs and CNNs, strong data augmentation, e.g., Randaugment (Cubuk et al.
-2020), CutMix (Yun et al.
-2019), and MixUp (Zhang et al.
-2018), show positive influence in CNN but are ineffective for ViTs AT and sometimes even reduce the AT performance of ViTs (Bai et al.
-2021; Singh, Croce, and Hein 2023).
-To mitigate this, the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix need to be progressively increased (Bai et al.
-2021).
-Similar problems of strong data augmentations are also observed by (Debenedetti, Sehwag, and Mittal 2023), where basic data augmentations, i.e., random-resize-and-crop, horizontal-flipping, and color-jitter with high weight decay are shown to be promising to improve ViT adversarial training.
-Another common practice in ViT adversarial training is to leverage pre-trained models, which has also been extensively explored (Mo et al.
-2022; Singh, Croce, and Hein 2023).
-Fine-tuning on natural pre-trained models could benefit the adversarial robustness of ViTs.
-For instance, recently, AdvXL (Wang et al.
-2024) introduced an efficient and effective adversarial training strategy for giant models on web-scale datasets.
-However, as the current ViT adversarial training strategies are mainly inherited from CNN AT, the key factor that drives the success of pre-training has not been completely revealed.
-For instance, pre-training with ImageNet-21K and SimMIM (Xie et al.
-2022) performs similarly to the model trained from scratch, and even the CLIP (Radford et al.
-2021) pre-trained model shows a performance drop (Liu et al.
-2024).
-In this work, we aim to provide an in-depth analysis of ViT’s robustness using MI, focusing on self-supervised pre-training.
-Information Bottleneck (IB) has been previously used in supervised learning to analyze adversarial robustness, where a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features is introduced to mitigate the impact of the adversarial noise in the inputs.
-Regarding the task of ViT adversarial training, we provide a novel theoretical analysis for the self-supervised autoencoder, showing that the self-supervised ViT benefits from decreasing the MI between inputs and latent features.
+Vision Transformers (ViTs) have emerged as a fundamental architecture and serve as the backbone of modern vision-language models.
+Despite their impressive performance, ViTs exhibit notable vulnerability to evasion attacks, necessitating the development of specialized Adversarial Training (AT) strategies tailored to their unique architecture.
+While a direct solution might involve applying existing AT methods to ViTs, our analysis reveals significant incompatibilities, particularly with state-of-the-art (SOTA) approaches such as Generalist [1] (CVPR 2023) and DBAT [2] (USENIX Security 2024).
+This paper presents a systematic investigation of adversarial robustness in ViTs and provides a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training.
+Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained via derived MI bounds.
+Building on this insight, we propose a self-supervised AT method, MIMIR, that employs an MI penalty to facilitate adversarial pre-training by masked image modeling with autoencoders.
+Extensive experiments on CIFAR-10, Tiny-ImageNet, and ImageNet-1K show that MIMIR can consistently provide improved natural and robust accuracy, where MIMIR outperforms SOTA AT results on ImageNet-1K.
+Notably, MIMIR demonstrates superior robustness against unforeseen attacks and common corruption data and can also withstand adaptive attacks where the adversary possesses full knowledge of the defense mechanism.
+Our code and trained models are publicly available at: https://github.com/xiaoyunxxy/MIMIR.
+publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23 - 27 February 2026 , San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.241813 www.ndss-symposium.org
+I Introduction
+ViTs [3] and their variants [4, 5] have achieved substantial progress and serve as foundational components in modern vision-language models.
+Prominent multimodal frameworks, including CLIP [6], BLIP [7], and Mini-GPT4 [8], typically employ ViTs as their image encoders.
+However, similar to convolutional neural networks (CNNs), attention-based models provide limited robustness against evasion attacks [9, 10, 11, 12, 13].
+Evasion attacks [14, 15] (also known as adversarial attacks), where well-trained deep models are fooled by introducing human-imperceptible perturbations to inputs, remain a persistent challenge in deep learning security.
+In 2024, the National Institute of Standards and Technology (NIST) explicitly listed adversarial attacks as a significant threat to AI systems, and pointed out the importance of conducting robustness testing and mitigation, such as adversarial training and formal verification, when deploying AI tools [16].
+Nevertheless, improving adversarial robustness remains a difficult task, where even SOTA methods, such as [17, 18, 19], achieve only marginal robustness gains, commonly below 2% compared with those before them.
+So far, Adversarial Training (AT) is widely recognized as the most practically effective defense [12, 11, 20] against evasion attacks.
+AT operates by augmenting the training dataset with adversarially perturbed samples [21], yet introduces two key limitations: (1) substantial computational overhead due to the generation of adversarial examples during training [21], and (2) a potential degradation in natural accuracy [22].
+Numerous methods have been proposed to mitigate these challenges.
+Techniques such as FreeAT [23] optimize efficiency by reusing gradient information during adversarial example generation, while FastAT [24] employs an improved Fast Gradient Sign Method (FGSM) to accelerate training.
+TRADES [25], SCORE [26], Generalist [1], and DBAT [2] explore how to achieve the best trade-off between natural and robust accuracy.
+Additionally, pre-training strategies have also been leveraged to enhance the performance of AT [27, 17].
+Applying existing AT methods to ViTs presents unique challenges due to the fundamental differences between attention-based architectures and CNNs.
+Unlike CNNs, ViTs lack inductive biases [3], including locality, two-dimensional neighborhood structure, and translation equivariance.
+These biases are inherent to CNNs as prior knowledge, enabling efficient learning with limited data, whereas ViTs typically require larger training datasets to achieve comparable generalization performance [3].
+Consequently, AT for ViTs entails significantly higher computational costs.
+Initial research on AT for ViTs explored their unique attention mechanism.
+For instance, robustness can be improved by randomly dropping gradients according to attention [11] or improving training efficiency by dropping low-attention image embeddings [12].
+The majority of recent works have focused on adapting CNN-based AT methodologies to ViTs, given AT’s success in building robust CNNs.
+Unfortunately, standard CNN AT techniques are not fully transferable to ViTs.
+Empirical studies [9, 20] reveal that strong data augmentations (such as Randaugment [28], CutMix [29], and MixUp [30], which improve robustness in CNNs) often degrade AT performance for ViTs.
+To mitigate this, recent work [9] suggests progressively increasing augmentation intensity (e.g., distortion magnitudes in RandAugment or the sampling probability of MixUp/CutMix) during training.
+Furthermore, SOTA AT strategies, such as Generalist [1] and DBAT [2], are less effective when applied to ViTs (see Table I), and there is a lack of evaluation on large datasets, such as ImageNet-1K, which further limits their generalizability.
+Pre-training has emerged as a complementary approach to ViT AT, with studies showing that adversarial fine-tuning of naturally pre-trained models can enhance robustness [11, 17].
+AdvXL [31] notably advanced this paradigm by developing efficient AT for web-scale datasets.
+However, the mechanisms underlying pre-training’s effectiveness are not fully understood, and results are inconsistent across implementations.
+For instance, models pre-trained on ImageNet-21K using SimMIM [32] demonstrate comparable performance to scratch-trained counterparts, while CLIP [6] pre-training has been shown to degrade performance in some configurations [33].
+While previous methods of ViT AT, such as [17, 20, 9], focus on searching for better combinations of training hyperparameters, they suffer from performance drops across different architectures and datasets.
+In contrast, we aim for a generalizable method via pre-training.
+Specifically, this work presents a systematic investigation of self-supervised pre-training for ViT robustness through the lens of Mutual Information (MI) and Information Bottleneck (IB) theory.
+IB introduces a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features to mitigate the impact of the adversarial noise in the inputs.
+Regarding the ViT AT, we develop a novel theoretical justification for self-supervised autoencoders, demonstrating that reducing MI between inputs and latent features enhances ViT robustness.
 Based on this finding, we propose a theoretically grounded adversarial pre-training method, Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR).
+[^1] Mimir is a figure in Norse mythology, renowned for his knowledge and wisdom.
+Specifically, we convert Masked Image Modeling (MIM) into an effective and efficient adversarial pre-training method.
+The basic idea is to predict the masked content of inputs, which is a self-supervised learning task.
+The effectiveness of MIMIR is analyzed and guaranteed by our theoretical justification.
+The efficiency comes from discarding the masked content (75% of image patches are discarded in our experiments), which greatly reduces the computing requirements.
 Figure 1 provides an illustrative diagram of MIMIR.
-MIMIR substantially improves the adversarial robustness of ViT models as demonstrated by comprehensive experiments on ImageNet-1K (Deng et al.
-2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al.
-2009).
-We also provide an experimental comparison with the state-of-the-art, showing the effectiveness of MIMIR in improving both natural and adversarial prediction accuracy (see Section 4.2).
+We validate MIMIR’s effectiveness through extensive experiments on CIFAR-10 [34], Tiny-ImageNet [35], and ImageNet-1K [36], showing consistent improvements in both natural and adversarial accuracy.
+In addition, we test the generalizability of MIMIR by combining MIMIR with three MIM methods for three representative architectures, including MAE [37] for ViTs, Group Window Attention [38] for hierarchical transformer (Swin [39]), and SparK [40] for CNN (ConvNext [41]), where MIMIR outperforms SOTA AT methods on ImageNet-1K.
 Our main contributions are:
-  • By revising the current ViT AT strategies, we point out that ViT adversarial pre-training methods compromise natural accuracy and lack systematic analysis.
-To this end, we provide a theoretical analysis using adversarial examples and the Mutual Information penalty.
-The MI bounds motivate us to decrease MI between adversarial examples and learned latent representation.
-  • Based on the findings of our analysis, we propose an efficient self-supervised defense MIMIR against adversarial attacks on ViTs.
-We systematically evaluate MIMIR using three datasets and multiple model scales under various adversarial attacks, demonstrating its superiority.
-  • We show MIMIR is resistant to two adaptive attacks, where the adversary is aware of MIMIR design.
-We build a PGD-fea attack that increases the distance between features extracted from natural and adversarial examples.
-Moreover, we build a PGD-MI attack that includes the MI penalty as the learning objective to find adversarial examples.
+  • By revisiting the current ViT AT strategies, we point out that ViT AT methods compromise natural accuracy and lack systematic study.
+To this end, we provide a theoretical analysis using adversarial examples and the Information Bottleneck.
+We theoretically show that the Mutual Information between adversarial examples and the learned latent representation should be decreased for better robustness.
+  • Based on our analysis, we propose a self-supervised defense, MIMIR, against adversarial attacks on ViTs.
+We evaluate MIMIR using multiple architectures on three datasets under various adversarial attacks, demonstrating its effectiveness.
+MIMIR achieves SOTA adversarial robustness on ImageNet-1K following the standardized evaluation by RobustBench.
+[^2] https://robustbench.github.io/
+  • We show that MIMIR is robust against unforeseen attacks and common corrupted data (ImageNet-C [42]) and can resist adaptive attacks where the adversary is aware of MIMIR’s design.
 Figure 1: Diagram illustrating the working mechanism of MIMIR.
-In the pre-training, adversarial perturbations $\delta$ are generated and added to natural images $x$ to create adversarial images $x+\delta$, as shown on the left.
-Patches of generated adversarial images are fed as training data to the ViT-based autoencoder, where the output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss.
-In particular, we propose using MI ($I(x+\delta,z)$) as an additional penalty to compose the pre-training loss of MIMIR, as shown in the bottom right.
-After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final model that is further fine-tuned for classification tasks.
-2 Background and Related Work
-[…]
+In the pre-training, adversarial images $x+\delta$ are generated and separated into image patches as the inputs of the encoder.
+The output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss.
+After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final complete model.
+Then the complete model is further fine-tuned for classification tasks.
+II Background
+II-A Evasion Attack
+Evasion attacks [43, 14, 15], also known as adversarial attacks, refer to applying imperceptible perturbations to the original input of the machine learning model (in this work, neural network), which generates adversarial examples to fool the victim model.
+Given an $L$-layer neural network $F_{\theta}$ for classification in $d_{Y}$-dimensional space and a training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ in $d_{X}$-dimensional space, the two primary goals of adversarial attacks are:
+  1. 1.
+    The generated perturbation $\delta$ can successfully mislead the network by maximizing (e.g., PGD attack [21]):
     $\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}),$ (1)
-The generated adversarial examples are as similar as possible to the original natural examples by limiting $\delta$ to a relatively small domain:
-[…]
+    where $x_{i}\in\mathbb{R}^{d_{X}}$ and $y_{i}\in\{0,1\}^{d_{Y}}$, $\theta$ are the parameters of the current network and $\mathcal{L}_{CE}$ is the standard CE (Cross-Entropy) loss.
+The perturbation aims to decrease confidence in the ground truth labels while increasing confidence in wrong labels.
+Thus, the loss between the misleading outputs and the ground truth labels increases.
+  2. 2.
+    The generated adversarial examples are as similar as possible to the original clean examples by limiting $\delta$ to a relatively small domain:
+    $S=B(x_{i},r)=\{\delta\in\mathbb{R}^{d_{X}}:\left\|\delta\right\|_{\infty}\leq r\},$ (2)
     where $S$ is the $l_{\infty}$-ball of radius $r$ at a position $x_{i}$ in $\mathbb{R}^{d_{X}}$.
-The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{0},l_{2},$ and $l_{\infty}$.
-When using the above adversarial examples to conduct adversarial training, the learning objective is:
+The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{2}$ and $l_{\infty}$.
+When using the above attacks to generate adversarial examples for AT, the learning objective is:
 $\mathop{\min}\limits_{\theta}\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{CE}(\theta,x_{i}+\delta,y_{i}).$ (3)
-Information Bottleneck (IB).
-The IB concepts were first proposed in (Tishby, Pereira, and Bialek 2000) and further developed for deep learning in (Tishby and Zaslavsky 2015; Shwartz-Ziv and Tishby 2017).
-It has been widely used to improve adversarial robustness (Xu, Perin, and Picek 2023; Alemi et al.
-2017; Yu, Yu, and Príncipe 2021; Wang et al.
-2021).
-IB describes the generalization of a deep neural network in two phases: 1) empirical error minimization (ERM) and 2) representation compression.
-For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer.
-The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$.
-The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I$.
-Adversarial Robustness of ViTs.
-PGD (Madry et al.
-2018) adversarial training is considered one of the most effective defenses that can withstand adaptively designed attacks (Athalye, Carlini, and Wagner 2018).
-Due to the difference between CNNs and ViTs, there have been substantial recent efforts to explore new adversarial training approaches for ViTs.
-EasyRobust (Mao et al.
-2022) evaluates the robustness of ViTs using PGD adversarial training.
-(Mo et al.
-2022; Bai et al.
-2021; Debenedetti, Sehwag, and Mittal 2023) demonstrate the classical training recipe for CNN is not suitable for ViTs and present new adversarial training strategies for ViTs by evaluating different hyperparameters and data augmentation policies.
-(Mo et al.
-2022) also finds that pre-training with natural data provides better robustness after adversarial fine-tuning.
-(Singh, Croce, and Hein 2023) explores the effect of strong pre-training and replacing the patch embedding in ViTs with convolutional blocks.
-AdvXL (Wang et al.
-2024) shows an outstanding performance of giant models and web-scale data with adversarial pre-training and fine-tuning.
-However, (Liu et al.
-2024) provides a different view that pre-training with ImageNet-21K and SimMIM (Xie et al.
-2022) performs similarly to the model trained from scratch, and the CLIP pre-trained model has a performance drop.
-Unlike existing works, we provide a formal and in-depth analysis leveraging MI, based on which we introduce a new adversarial pre-training scheme.
-3 MIMIR
-3.1 Threat Model
+II-B Masked Image Modeling - MIM
+MIM refers to a self-supervised pre-training framework that aims to reconstruct pre-defined targets, such as discrete tokens [44], raw RGB pixels [37, 45], or features [46].
+The final goal is to use the pre-trained model as a starting point for downstream fine-tuning.
+The downstream tasks include, for instance, classification and object detection.
+More specifically, to build a high-performance ViT $f_{e}$ without a classification layer, we consider $f_{e}$ as an encoder to extract discriminative input features.
+Then, we design a lightweight decoder $f_{d}$, which uses the output of $f_{e}$ as its input.
+The goal of $f_{d}$ is to reconstruct the original inputs (let us consider MAE [37] as an example).
+The aim is to decrease the distance between the input $x$ and $x_{re}=f_{d}\circ f_{e}(x)$.
+After the encoder $f_{e}$ and decoder $f_{d}$ are trained, we use $f_{e}$ plus a manually initialized classification layer as the starting point of fine-tuning.
+II-C Mutual Information - MI
+MI measures the mutual dependence between two random variables, $X$ and $Y$.
+It quantifies the amount of information contained in one random variable about another random variable or the reduced uncertainty of a random variable when another random variable is known.
+It can be written as:
+$I(X,Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{(X)}(x)P_{(Y)}(y)}\right),$ (4)
+where $P_{(X,Y)}$ is the joint probability density function of $X$ and $Y$. $P_{(X)}$ and $P_{(Y)}$ are the marginal probability density function of $X$ and $Y$. MI can be equivalently expressed as:
+$I(X,Y)$ $=H(X)-H(X|Y).$ (5)
+Estimating the exact MI is not easy, as it is difficult to precisely estimate $P_{X,Y}$ or $P_{X}$ and $P_{Y}$ in high-dimensional space [47].
+In practice, Deep Deterministic Information Bottleneck (DIB) [48] suggested using the matrix-based Renyi’s $\alpha$-order entropy $I_{\alpha}$ [49, 50] to estimate MI, which avoids density estimation and variational approximation.
+An alternative way is the Hilbert-Schmidt Independence Criterion (HSIC) [51], which is a kernel-based dependence measure defined in a reproducing kernel Hilbert space (RKHS) and is usually used as a surrogate of MI.
+Details about definitions and empirical estimators of $I_{\alpha}$ and HSIC are provided in Appendix -J.
+In this paper, we evaluate both $I_{\alpha}$ and HSIC as our MI measurements.
+II-D Information Bottleneck - IB
+The IB concepts were first proposed in [52] and further developed for deep learning in [53, 54].
+IB describes the generalization of a deep network in two phases: 1) empirical error minimization (ERM) and 2) representation compression [54].
+For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of the $l$-th layer.
+The IB principle aims to keep more relevant information in $t_{l}$ about the target $y$ while decreasing the irrelevant information about the input $x$.
+The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I(\cdot)$.
+During neural network training, in the ERM phase, the model increases shared information between $t_{l}$ with respect to both $x$ and $y$.
+Afterward, in the compression phase, the model decreases information contained in $t_{l}$ about $x$ but preserves (or even increases) information about $y$.
+The reduction of $I(x,t_{l})$ can be interpreted as a way of reducing noise or compressing irrelevant or redundant features in $x$.
+At the end of the training, the model strikes a trade-off that maximizes $I(y,t_{l})$ and minimizes $I(x,t_{l})$.
+Formally, the IB minimizes the following Lagrangian:
+$\mathcal{L}=I(x,t_{l})-\beta I(y,t_{l}),$ (6)
+where $\beta$ is a Lagrange multiplier that controls the trade-off between predicting $y$ and compressing $x$.
+III MIMIR
+III-A Threat Model
 Adversary’s goal.
-The attacker aims to fool the trained model by decreasing the overall classification accuracy (non-targeted) or forcing the model to recognize an input as a specific target (targeted).
-The image perturbations are limited by the $l_{\infty}$ norm for invisibility.
+The attacker aims to fool the trained model with both non-targeted and targeted attacks.
+The goal is to decrease the overall classification accuracy (non-targeted) or compel the model to recognize any inputs as a specific target (targeted).
+Meanwhile, the adversarial perturbations applied to the input should be invisible so that they will not be easily detected.
+During the training phase, the model optimizes its parameters to minimize the loss between predicted outputs and true labels, thereby enhancing classification accuracy.
+In contrast, the adversary’s objective is to develop algorithms that generate perturbations capable of maximizing this loss.
+For a targeted attack, the attacker decreases the loss between the output and the specified target label.
+To maintain the imperceptibility of the perturbations, the magnitude of the adversarial modifications is constrained by distance measures (such as $l2$ and $l_{\infty}$), ensuring that the alterations to the input data remain within a visually indistinguishable range.
 Adversary knowledge.
@@
 The attacker can implement iterative attacks and unlimited queries to update adversarial examples multiple times in white and black-box settings.
-Adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets.
-The adversary can adaptively design new attacks for a specific model according to the design details of the defense method.
-Defender’s Goal.
-The defender’s main goal is to train a robust model against potential adversarial attacks and prevent the performance of natural data from decreasing significantly.
-In addition, the defense method should be robust against adaptive attacks specifically designed for it.
-3.2 Design Intuition
-MIM has been proved useful as a pre-training method for ViTs on various tasks (Vincent et al.
-2010; He et al.
-2022; Wei et al.
-2022; Bao et al.
-2022).
-To train a powerful model using MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model.
-Masking out the foreground instead of the background reduces discriminative information in visible information to the model.
-Reconstructing foreground parts is harder than background and helps the model learn more discriminative information (Wang et al.
-2023).
-Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbations to natural inputs.
-The adversarial perturbations increase the distance between the input and reconstruction target.
-If we can reconstruct natural inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks.
-In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial perturbations while enabling the decoder to reconstruct the natural image.
-Note that simply masking out all foreground does not make sense, as it would render the reconstruction of meaningful content infeasible.
-From the perspective of IB, there is a bottleneck between the encoder and decoder.
-Our adversarial pre-training task contains two information sources: the natural data $x$ and the adversarial perturbations $\delta$.
-As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated.
-The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.
+Adversarial examples can be created according to model architectures, parameters, the gradients of the loss function, and datasets.
+In addition, we also consider adaptive adversaries who are also aware of potential defenses.
+The adversary can design new attacks for a specific model according to the design details of the defense method.
+Defender’s goal. From the defender’s perspective, the main goal is to train a robust model against potential adversarial attacks. The defender considers the following objectives:
+  • The defender aims to prevent the performance of natural data from decreasing significantly, but allows a slight drop of natural accuracy for a trade-off in exchange for robustness.
+  • The defense method should provide a notable improvement compared to models without defenses when subjected to various adversarial attacks, especially to adaptive attacks that are aware of the details of the defense method.
+  • The defense method should be efficient and scalable to large datasets such as ImageNet-1K [36].
+III-B Design Intuition
+MIM has been established as an effective pre-training paradigm for Vision Transformers (ViTs), demonstrating strong performance across diverse downstream tasks [55, 37, 46, 44].
+The core methodology involves masking foreground regions of input images and tasking the model with their reconstruction.
+Masking foreground (as opposed to background) regions removes high-information content and results in a harder task (than reconstructing background content), forcing the model to develop stronger feature representations to reconstruct semantically meaningful patterns [56].
+Building upon these principles, we introduce an adversarial extension to MIM by incorporating adversarial perturbations into the input space.
+Our formulation is grounded in three interconnected hypotheses:
+  • Adversarial Robustness through Reconstruction: If a model can reconstruct natural images from adversarially perturbed inputs, its latent representations must inherently discard perturbation-specific information while preserving natural data semantics.
+  • IB Perspective: The encoder-decoder architecture naturally imposes an information bottleneck.
+When processing adversarial examples $x+\delta$, the system must suppress perturbation-derived information $(\delta)$ while retaining natural data information $(x)$ to achieve accurate reconstruction (see Figure 1).
+  • Optimal Masking Strategy: Complete foreground masking (to build a difficult task) is suboptimal, as it eliminates essential reconstruction signals.
+Instead, our method employs partial masking to maintain a tractable yet challenging learning objective.
 Algorithm 1 MIMIR Pre-training
@@
 12:end for
-3.3 Design Details
+III-C Design Details
 Autoencoder.
-MIMIR follows the general design of MAE (He et al.
-2022), consisting of an encoder $f_{e}$ and a decoder $f_{d}$.
+MIMIR consists of an encoder $f_{e}$ and a decoder $f_{d}$ aligned with the general design of MAE [37].
 As with other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$.
-The decoder reconstructs original inputs according to the discriminative features.
-Following ViT (Dosovitskiy et al.
-2021), the input $x$ is separated into non-overlapping image patches.
+The decoder reconstructs the original inputs according to the discriminative features.
+Following the design of ViT [3], the input $x$ is separated into non-overlapping image patches.
 We randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder.
@@
 Thus, we aim to keep a part of the foreground as a hint for reconstruction.
-The information of masked content is recorded as mask tokens $m$, not used by the encoder but reserved for later use by the decoder.
+The information of masked content is recorded as mask tokens $m$, which are not used by the encoder but reserved for later use by the decoder.
 Each token is a learned vector indicating the presence of a masked patch to be predicted.
@@
 The encoder only processes the visible patches, making training much more efficient.
-When converting to other architectures, such as ConViT (D’Ascoli et al.
-2021), we use corresponding transformer blocks to build the encoder.
+When converting to other architectures, such as ConViT [5], we use corresponding transformer blocks to build the encoder.
 The decoder accepts the encoded visible image patches and mask tokens as inputs.
-The decoder is built using the same transformer blocks as the encoder instead of using ViT (Dosovitskiy et al.
-2021) transformer blocks for all.
+The decoder is built using the same transformer blocks as the encoder instead of using ViT [3] transformer blocks for all.
 Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
@@
 Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results.
-Following the original MAE (He et al.
-2022), this distance is measured by Mean Squared Error (MSE).
-To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ on the inputs of the encoder.
+Following the original MAE [37], this distance is measured by Mean Squared Error (MSE).
+To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ to the inputs of the encoder.
 Thus, the adversarial perturbations are also masked along with the image upon input into the encoder.
@@
 MI as Penalty.
-Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$ is decreasing while training.
+Inspired by IB, we show in Section III-D that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$, decreases while training.
 Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
@@
 This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$.
-According to Data Processing Inequality (DPI) (Beaudry and Renner 2011), $I(x,z)\leq I(x+\delta,z)$.
+According to Data Processing Inequality (DPI) [57], $I(x,z)\leq I(x+\delta,z)$.
 $I(x+\delta,z)$ is closer to $z$ on the Markov chain.
@@
 To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbations $\delta$.
-As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD (Madry et al.
-2018).
+As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD [21].
 Nevertheless, it is feasible to design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq.
@@
 Finally, the complete model is fine-tuned by AT methods.
-Figure 2: The example plots for the lower and upper bounds on the MI in Propositions 3.2 and 3.3.
+Figure 2: The example plots for the lower and upper bounds on the MI in Propositions III.2 and III.3.
 The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes.
 The lower bound reaches its minimum at $p_{e}=0.9$.
-3.4 Theoretical Justification
-Next, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained.
+III-D Theoretical Justification
+Next, we provide theoretical justification showing that MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained.
 Let $F$ denote any classifier trained on natural samples with desirable prediction accuracy, which may suffer from adversarial attacks.
-We begin our analysis by first presenting Lemma 3.1.
-Lemma 3.1.
+We begin our analysis by first presenting Lemma III.1.
+Lemma III.1.
 Let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$, we have:
@@
 Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$.
-Intuitively, our autoencoder is trained to recover only natural sample $x$ without any interference from $\delta$.
+Intuitively, our autoencoder is trained to recover only the natural sample $x$ without any interference from $\delta$.
 Hence, a relatively large value of $p_{e}$ is expected.
 In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
-[…]
-Architecture Params (M) FT FT Epoch Adv. Steps Source Natural Robust
-ResNet-50 25.0 PGD 100 1 (Bai et al. 2021) 67.44 35.54
-DeiT-S 22.1 PGD 100 1 (Bai et al. 2021) 66.62 36.56
-DeiT-S 22.1 PGD 110 1 (Debenedetti, Sehwag, and Mittal 2023) 66.80 37.90
-ViT-S 22.1 PGD 120 3 (Mao et al. 2022) 66.43 39.20
-ViT-S 22.1 PGD 300 3 (Liu et al. 2024) 70.7 43.7
-RobArch-S 26.1 PGD 110 3 (Peng et al. 2023) 70.17 44.14
-ViT-S 22.1 APGD 300 2 (Singh, Croce, and Hein 2023) 69.22 44.04
-ViT-S 22.1 APGD 300 2 Ours 71.00 46.10
-ViT-B 86.6 ARD+PRM 10 5 (Mo et al. 2022) 69.10 34.62
-ViT-B 86.6 PGD 120 3 (Mao et al. 2022) 70.64 43.04
-RobArch-L 104 PGD 100 3 (Peng et al. 2023) 73.44 48.94
-ViT-B 86.6 PGD 300 3 (Liu et al. 2024) 74.7 49.7
-[…]
-ViT-B 86.6 APGD 300 2 (Singh, Croce, and Hein 2023) 74.10 50.30
-ViT-B 86.6 APGD 300 2 Ours 76.32 54.28
-Proposition 3.2.
+Proposition III.2.
 Let $H(\cdot)$ denote the information entropy and $H_{b}(p_{e})=-p_{e}\log_{2}p_{e}-(1-p_{e})\log_{2}(1-p_{e})$ be the binary entropy, we have:
@@
 $I(F(x+\delta),F(x_{re}))=H(F(x+\delta))-H(F(x+\delta)|F(x_{re})).$ (15)
-By applying Fano’s inequality (Fano 1949; Ocal et al. 2019), we obtain:
+By applying Fano’s inequality [58, 59], we obtain:
 $H(F(x+\delta)|F(x_{re}))\leq H_{b}(p_{e})+p_{e}\log(|F(x+\delta)|-1).$ (16)
@@
 $\leq I(x+\delta,z).$
-The last line of Eq. (14) is by Lemma 3.1. ∎
+The last line of Eq. (17) is by Lemma III.1. ∎
 Therefore, we obtain a lower bound of $I(x+\delta,z)$.
 If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$.
-In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq.
+In fact, we can also obtain an upper bound under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq.
 (11).
-Proposition 3.3.
+Proposition III.3.
 If $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, we have:
-$I(x+\delta,z)\lesssim H(F(x_{re}))-2p_{e},$ (15)
+$I(x+\delta,z)\lesssim H(F(x+\delta))-2p_{e},$ (18)
 in which the notation $\lesssim$” refers to less than or similar to.
 Proof.
-By the Hellman-Raviv inequality (Hellman and Raviv 1970; Brown 2009), we have:
+By the Hellman-Raviv inequality [60, 61], we have:
 $2p_{e}$ $\leq H(F(x+\delta)|x_{re})$ (19)
@@
 In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$.
-For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^3] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations (Moosavi-Dezfooli et al.
-2017)., then:
+For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^4] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations [62]., then:
 $I(x+\delta,z)=I(x,z)+I(\delta,z).$ (20)
-According to (Vincent et al.
-2008), minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input.
+According to [63], minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input.
 Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized).
@@
 This also explains the robustness of $z$.
-4 Experiments
-4.1 Experimental Setup
-We evaluate MIMIR on three datasets: ImageNet-1K (Deng et al.
-2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al.
-2009).
-Details of datasets are provided in Appendix A.
-Hyperparameters of the decoder are included in Appendix B.
+IV Experimental Evaluation
+IV-A Experimental Setup
+We evaluate MIMIR on three datasets: ImageNet-1K [36], Tiny-ImageNet [35], and CIFAR-10 [34], with diverse and commonly used architectures with multiple scales: ViT [3], ConViT [5], Swin Transformer [39], and ConvNext [41].
+Details of datasets are provided in Appendix -A.
+Hyperparameters of the decoder are included in Appendix -B.
 Training Setup.
 We train models from scratch for all experiments.
-Following MAE (He et al.
-2022), we do pre-training by MIMIR for 800 epochs.
+Following MAE [37], we do pre-training by MIMIR for 800 epochs.
+Please note that we also compare our MIMIR + fine-tuning paradigm with the End2End paradigm.
+The End2End paradigm refers to the supervised training of a model from scratch without self-supervised pre-training.
+To compare between End2End and pre-training + fine-tuning, the common training schedule practice is pre-training 800 epochs + fine-tuning 100 epochs versus End2End training 300 epochs in the existing works [37, 32, 64].
 The number of warmup epochs is 40 for pre-training.
-We use AdamW (Loshchilov and Hutter 2019) as an optimizer for both pre-training and fine-tuning.
-The cosine decay as the learning rate scheduler is applied.
+We use AdamW [65] as the optimizer for both pre-training and fine-tuning.
+We apply the cosine decay as the learning rate scheduler.
 At pre-training, MIMIR uses the 1-step PGD to generate adversarial examples for all three datasets.
 The perturbation budget is $\epsilon=8,\alpha=10$.
-At fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet.
-For ImageNet-1K, we use 1-step PGD with random initialization for better efficiency.
-The perturbation bound is $\epsilon=4$ for ImageNet-1K.
-We also use a 2-step APGD with a longer fine-tuning scheduler to compare with SOTA works in Table 1.
-Details on training hyperparameters are provided in Appendix C.
+For the fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet.
+For ImageNet-1K, the perturbation bound is $\epsilon=4$.
+We use different steps of PGD or APGD to generate the adversarial perturbation for training.
+Details on training hyperparameters are provided in Appendix -C.
+TABLE I: Comparison between End2End AT and Pre-training (MIMIR) + Fine-tuning using ViT-S on CIFAR-10.
+Training Method Natural PGD AutoAttack
+End2End AT [21] 75.36 32.84 26.17
+Fast AT [24] 76.81 32.57 21.41
+TRADES [25] 74.96 32.12 24.90
+MART [66] 72.42 24.47 23.45
+Generalist [1] 60.88 14.44 11.20
+DBAT [2] 68.32 18.83 5.25
+MIMIR
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+TABLE II: Standard deviation of 3 runs on 3 datasets.
+Performance CIFAR-10 Tiny-ImageNet ImageNet-1K
+[…]
+[…]
+TABLE III: Comparison with SOTA results on ImageNet-1K under $\epsilon=4/255$.
+The “Adv.
+Steps” refers to attack steps for generating adversarial examples for AT.
+The results of [31] are evaluated using 20-step PGD (AutoAttack is designed as a more powerful alternative to PGD), which is marked as $\dagger$} in the table.
+Although AdvXL only uses 20 epochs, it takes more time due to huge datasets for pre-training and fine-tuning.
+Architecture Params (M) FT FT Epoch Adv. Steps Source Natural AutoAttack
+DeiT-S 22.1 PGD 100 1 Augmentation warm-up [9] 66.62 36.56
+DeiT-S 22.1 PGD 110 1 Light Recipe [20] 66.80 37.90
+ViT-S 22.1 PGD 120 3 EasyRobust [67] 66.43 39.20
+ViT-S 22.1 PGD 300 3 Adversarially Trained [33] 70.7 43.7
+RobArch-S 26.1 PGD 110 3 RobArch [18] 70.17 44.14
+ViT-S 22.1 APGD 300 2 Pre-training+AT [17] 69.22 44.04
+ViT-S 22.1 PGD 300 3 MIMIR 71.52 45.90
+ViT-S 22.1 APGD 300 2 MIMIR 71.00 46.10
+ViT-S 22.1 APGD 300 3 MIMIR 70.96 46.16
+ViT-B 86.6 ARD+PRM 10 5 ARD+PRM [11] 69.10 34.62
+Swin-B 87.7 ARD+PRM 10 5 ARD+PRM [11] 74.36 38.61
+ViT-B 86.6 PGD 120 3 EasyRobust [67] 70.64 43.04
+Swin-B 87.7 PGD 120 3 EasyRobust [67] 75.05 47.42
+RobArch-L 104 PGD 100 3 RobArch [18] 73.44 48.94
+ViT-B 86.6 PGD 300 3 Adversarially Trained [33] 74.7 49.7
+[…]
+ViT-B 86.6 APGD 300 2 Pre-training+AT [17] 74.10 50.30
+ViT-B 86.6 APGD 100 2 MIMIR 74.40 51.92
+ViT-B 86.6 PGD 100 3 MIMIR 75.68 52.96
+ViT-B 86.6 PGD 300 3 MIMIR 76.98 53.84
+ViT-B 86.6 APGD 300 2 MIMIR 76.32 54.28
 Evaluation Metrics.
@@
 Natural accuracy refers to the accuracy of natural and unmodified inputs.
-The robust accuracy measures the accuracy under the AutoAttack (Francesco and Hein 2020).
+The robust accuracy measures the accuracy under the AutoAttack (AA) [68].
 AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks.
-In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce (Francesco and Hein 2020), APGD-t (Francesco and Hein 2020), FAB-t (Croce and Hein 2020), and Square (Andriushchenko et al.
-2020).
-The perturbation budgets are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K.
-4.2 Experimental Results
-Comparison with SOTA Works.
-Table 1 compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$).
-We report the number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy.
-The robust accuracy is evaluated by AutoAttack on the RobustBench validation set (5000 images).
-We divide the models into two groups: small ($\approx 22$M) and large ($\approx 86$M) models.
-In addition, (Singh, Croce, and Hein 2023) shows that the Convolutional stem (ConvStem) improves the robustness of ViTs.
-We also include an evaluation of ViTs+ConvStem in Table 2.
-ConvStem (Xiao et al.
-2021) refers to replacing the patch embedding in ViTs with a convolutional block.
-MIMIR outperforms all SOTA works.
-Table 2: Influence of the ConvStem block.
-The Standard pre-trained weights refer to pre-trained weights from timm.
-The configurations are deit_small_patch16_224 and vit_base_patch16_224.augreg_in1k.
-The weights are fine-tuned 300 epochs by 2-step APGD.
-†The result on ViT-S is taken from (Singh, Croce, and Hein 2023).
-Architecture Pre-trained weights Natural Robust
-ViT-S+ConvStem Standard 72.56 48.08†
-MIMIR 72.72 48.44
-ViT-B+ConvStem Standard 75.26 52.70
-MIMIR 76.72 54.04
-Ablation Study.
-Table 3 provides an ablation study to verify the design choices of MIMIR.
+In our experiments, we use the standard version of AutoAttack that contains four attacks, including APGD-ce [68], APGD-t [68], FAB-t [69], and Square [70].
+The perturbation budgets for evaluation are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K.
+In addition, we also evaluate the MIMIR-trained models with unforeseen attacks, such as CW attacks, attacks with $l_{2}$ norm, and out-of-distribution data (ImageNet-Corruption [42]).
+Training stability.
+We also show that the natural accuracy and robustness of MIMIR are stable evaluation metrics.
+Due to the high computation cost, we cannot report the standard deviation for all experiments.
+To show that our method MIMIR has low variances, we train ViT-S on three datasets three times (1-step PGD AT for ImageNet-1K) and report the standard deviation and average performance in Table II.
+IV-B Main Results
+We first explore different AT methods and MIMIR for ViT on CIFAR-10, demonstrating the fundamental incompatibility between conventional AT approaches and ViT architectures.
+Following this baseline evaluation, we scale our investigation to the more challenging ImageNet-1K dataset, demonstrating the generalizability and scalability of our proposed MIMIR framework.
+The subsequent sections present comprehensive experimental results across three benchmark datasets: CIFAR-10, Tiny-ImageNet, and ImageNet-1K.
+This multi-scale evaluation strategy allows a thorough analysis of MIMIR’s effectiveness under varying conditions, from smaller to large-scale visual recognition tasks.
+In addition, we also explore the effectiveness of elucidating diffusion model (EDM) data on AT in Appendix -D.
+This generated data is commonly used in AT to improve robustness [71, 72, 73, 74].
+Specifically, we use 5 million generated CIFAR-10 data and 1 million Tiny-ImageNet data provided by [71].
+Table XVIII in Appendix -D shows that EDM data significantly improves the robustness.
+CIFAR-10.
+Table I shows the performance of End2End adversarial training from scratch and Pre-training (MIMIR) + Fine-tuning on ViT-S trained on CIFAR-10.
+We provide the performance of 6 established or SOTA AT methods on CIFAR-10, indicating that traditional AT training strategies are not applicable to ViTs.
+Importantly, our experimental results also demonstrate that MIMIR can substantially improve all AT methods.
+The reason is that training ViTs from scratch is known to be difficult [3, 75] and even more difficult for adversarial training [11].
+For example, robust accuracy is lower than 30% on ViT-B without pre-training [11].
+In contrast, MIMIR provides a more straightforward methodology and avoids this difficulty by switching to pre-training with a theoretically grounded MIM learning task.
+TABLE IV: Time consumption of End2End vs. Pre-training+Fine-tuning.
+Arch GPU AT Method Epochs Hours
+ViT-S 4 A6000 PGD10 300 187.64
+4 A6000 MIMIR(PT)+PGD10(FT) 800(PT)+100(FT) 123.76
+ViT-B 4 A6000 PGD10 300 451.39
+4 A6000 MIMIR(PT)+PGD10(FT) 800(PT)+100(FT) 263.77
+Swin-L 4 H100 PGD3 300 363.42
+4 H100 MIMIR(PT)+PGD3(FT) 800(PT)+100(FT) 231.14
+TABLE V: The performance of MIMIR on ImageNet-1K against unforeseen threats. The $l_{2}$ version of the CW [76] attack is limited by $c$. Higher $c$ allows a more powerful perturbation.
+Architecture Method Natural CW ($l_{2},c=1$) CW ($l_{\infty},\epsilon=4/255$) AA ($l_{2},\epsilon=2.0$) AA ($l_{\infty},\epsilon=4/255$) ImageNet-C [42]
+ViT-S [17] 69.22 44.28 45.98 37.52 44.04 45.62
+MIMIR 70.96 66.34 49.22 49.12 46.16 48.78
+ViT-B [17] 74.10 59.42 52.76 52.12 50.30 53.69
+MIMIR 76.32 65.30 56.82 56.72 54.28 57.07
+Time Consumption (End2End vs.
+Pre-training(PT)+Fine-tuning(FT)).
+Note that we follow the standard way to compare End2End and MIMIR (Pre-training + Fine-tuning) training methods by fixing the training schedule, following existing works [37, 32, 64, 77, 40], where we include pre-training 800 epochs + fine-tuning 100 epochs versus supervised End2End training of 300 epochs.
+The reason for having a larger number of pre-training epochs is that self-supervised pre-training is much more efficient (see Table XXII in Appendix -I) than End2End training, and the pre-trained backbone can be used multiple times for various fine-tuning tasks.
+For example, in Table I, the 6 End2End AT methods cost $300\times 6=1800$ fine-tuning epochs.
+MIMIR costs 800 pre-training epochs + $100\times 6$ fine-tuning epochs, i.e., we only conduct the pre-training once for results in Table I.
+MIMIR pre-training epoch is more efficient than a fine-tuning epoch by discarding 75% of image patches.
+More specifically, Table IV shows the total time consumption of End2End vs.
+Pre-training+Fine-tuning on three architectures with ImageNet-1K.
+Although MIMIR takes more training epochs, its pre-training+fine-tuning paradigm still costs less time than End2End adversarial training and gains much better performance on both natural and adversarial examples.
+Additional time consumption results with different datasets can be found in Table XXII in Appendix -I.
+ImageNet-1K.
+Table III compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$), which follows the evaluation of common standardized RobustBench [78].
+Similar to other works [20, 33, 17], we consider simpler AT methods (i.e., PGD and APGD AT) instead of the latest AT methods, such as Generalist [1] and DBAT [2].
+Indeed, since the latest methods introduce tailored components for CNNs to improve their adversarial robustness, they might not be effective for ViTs.
+The number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy are reported to provide a more detailed understanding of the performance.
+The robust accuracy is evaluated by AutoAttack on the RobustBench [78] validation set (5,000 images).
+We divide the models into: small ($\approx 22$M) and large ($\approx 86$M) models, corresponding to ViT-S and ViT-B.
+Experimental results demonstrate that MIMIR outperforms all previous works across various training setups.
+MIMIR with Various Architectures.
+In Table VI, we show that MIMIR can be applied to diverse architectures.
+Specifically, we use three representative options, including ViT+convolutional blocks (CVST) [17], the latest CNN architecture (ConvNext [41]), and a hierarchical vision transformer (Swin [39]).
+The ViT+CVST refers to using ConvStem [79] to replace the patch embedding in ViTs with a convolutional block.
+The ViT+CVST shows improved robustness compared to pure ViT models according to experiments in [17].
+As CNN and hierarchical architecture cannot accept variable-length inputs, MIMIR is not directly compatible with ConvNext and Swin.
+To adapt MIMIR to the hierarchical Swin Transformer, we implemented Masked Image Modeling using Group Window Attention [38], which groups image patches within each local window of arbitrary size and performs masked self-attention in each group.
+To apply MIMIR to ConvNext, we use SparK [40] for CNN to handle irregular and randomly masked input images, which is achieved by sparse convolution.
+MIMIR achieves better or comparable results compared to SOTA results on RobustBench [78].
+MIMIR against Unforeseen Attacks.
+Except for adversaries who are limited by the adversarial budget, e.g., $l_{\infty}=8\text{ or }4$ or from PGD-family in our main experiments, MIMIR also shows the potential to provide robustness against unforeseen attacks and naturally corrupted data (ImageNet-C [42]).
+Table V demonstrates the robustness of MIMIR against practical unforeseen attacks, including non-PGD attacks ($l_{2}$ and $l_{\infty}$ CW attack [76]), $l_{2}$ AutoAttack, and ImageNet-C [42].
+It is clear that MIMIR still performs well against these unforeseen attacks.
+In particular, MIMIR achieved top accuracy compared to the results of the ImageNet-C Leaderboard on the RobustBench.
+IV-C Ablation Study and Further Analysis
+Step by step ablation.
+Table VII provides an ablation study to verify the design choices of MIMIR.
 The ablation uses 100 epochs of 1-step PGD (PGD1) AT as the baseline.
-Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library) as initialization of AT.
+Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library [^5] https://github.com/huggingface/pytorch-image-models/blob/main/timm/models/vision_transformer.py) as initialization of AT.
 After that, we replace the clean pre-training with MAE, adv MAE, and MIMIR step by step.
-The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in loss.
-The pre-training scheduler is 800 epochs.
+The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in the loss.
+The pre-training schedule is 800 epochs.
 We also use stronger adversarial fine-tuning for better performance, including 2-step PGD (PGD2), APGD (APGD2) FT, and a longer fine-tuning scheduler (300 epochs).
-Please note that a longer training scheduler does not guarantee better results (Mo et al.
-2022).
-Our results indicate that MIMIR outperforms baselines and can be further improved under the long training scheduler.
-In addition, MIMIR is also efficient and increases less than 5% of time consumption on ImageNet-1K compared to MAE, which is shown in Table 12 in the Appendix E.
+Our results indicate that MIMIR outperforms baselines and can be further improved under the long training schedule.
+Figure 3: Natural, adversarial accuracy, and MIMIR pre-training loss of ViT-S with different numbers of pre-training epochs under a 20-step PGD attack.
+The performance increases as the number of pre-training epochs increases (loss decreases).
+TABLE VI: Comparsion with SOTA ImageNet-1K results on RobustBench [78] with different architectures. $\dagger$}: The CVST modules are also pre-trained with MIMIR.
+Architecture Method FT Epoch Natural AutoAttack
+ViT-S+CVST [17] 300 72.56 48.08
+MIMIR 300 72.72 48.44
 […]
-Architecture Training Recipe Natural Robust
+ViT-B+CVST [17] 250 76.30 54.66
+MIMIR 300 76.72 54.04
+[…]
+ConvNext-T [17] 300 72.40 48.60
+MIMIR 300 72.50 48.76
+Swin-B [33] 300 76.16 56.16
+MIMIR 150 76.62 55.90
+Swin-L [33] 300 78.92 59.56
+MIMIR 100 78.62 59.68
+Longer epochs (lower loss) provide better performance.
+Pre-training epoch.
+Our experimental framework employs an 800-epoch pre-training as the baseline configuration.
+To systematically evaluate the impact of training duration, we conduct a comprehensive ablation study using ViT-S architectures, varying the number of pre-training epochs while maintaining a fixed 50-epoch fine-tuning for all models.
+As demonstrated in Figure 3, we observe several key phenomena.
+Extended pre-training schedules consistently yield lower MIMIR loss values, and this reduction in loss results in measurable improvements in both adversarial and natural accuracy.
+In addition, the improved performance with loss MIMIR loss also indicates that the model capacity is not saturated within the tested epoch range and could be further improved with a larger number of epochs.
+Fine-tuning epoch.
+To isolate and quantify the contribution of MIMIR pre-training to model performance, we employ a short fine-tuning for the pre-trained models.
+This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training.
+This approach allows us to evaluate the quality of the representations learned during pre-training.
+In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning.
+These results suggest that the majority of the model’s final performance is attributable to the MIMIR pre-training phase.
+TABLE VII: Ablation of pre-training (PT) and fine-tuning (FT) methods on ImageNet-1K.
+($\dagger$}: catastrophic over-fitting [24] due to 1-step AT when fine-tuning, which can be fixed by 2-step AT.
+The fixed natural and robust accuracy are 69.96 and 36.90, respectively.)
+Architecture Training Recipe Natural AutoAttack
 ViT-S PGD1 FT w/o PT 66.02 31.40
@@
 MIMIR PT + APGD2 FT 68.78 42.86
-100 $\rightarrow$ 300 epochs FT 71.00 46.10
-Table 4: Comparison between HSIC (Gretton et al.
-2005) and $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) using ViT-T on CIFAR-10.
-Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
+100 $\rightarrow$ 300 epochs of FT 71.00 46.10
+MI measure.
+In Section III-D, we provide lower and upper bound (Eq.
+(14)) of $I(x+\delta,z)$.
+According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$.
+This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective.
+In this paper, we use $I_{\alpha}$ [48] and HSIC [80] as estimators (detailed definitions in Appendix -J).
+Table VIII demonstrates the performance with different values of $\lambda$.
+According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.
+In addition, Figure 4 provides the quantities of HSIC values while pre-training with or without using MIMIR.
+It is clear that MIMIR can help to decrease the mutual information between adversarial perturbation and the learned features, i.e., $I(x+\delta,z)$.
+Figure 4: The HSIC values (we use HSIC as an alternative to MI) while pre-training ViT-S with and without using MIMIR.
+TABLE VIII: Comparison between HSIC [51] and $I_{\alpha}$ [48] using ViT-T on CIFAR-10. Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
 Pre-train $\lambda$ Estimator Natural PGD
@@
 MIMIR 1e-06 $I_{\alpha}$ 74.60 45.66
-Table 5: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
+1-step is better than the 10-step of AT in pre-training.
+We also show that MIMIR outperforms original MAE [37] and adv MAE with different PGD steps (to generate adversarial examples for training).
+MAE in Table IX refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD.
+The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in the loss.
+The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training.
+The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task.
+This more difficult task further improves the performance of downstream models (see also Table XII).
+We use the default learning rate (i.e., 5.0e45.0e-4) of MAE, so there is a performance drop in experiments in Tables VIII and IX since AT prefers larger learning rates for CIFAR-10, as shown in Table XVII in the appendix.
+TABLE IX: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
 Pre-train $\lambda$ Estimator Natural PGD
@@
 MIMIR 1e-05 HSIC 76.30 47.60
-MI Measure.
-In Section 3.4, we provide lower and upper bound (Eq.
-(11)) of $I(x+\delta,z)$.
-According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$.
-This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective.
-In this paper, we use $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) and HSIC (Ma, Lewis, and Kleijn 2020) as estimators (detailed definitions in Appendix F).
-Table 4 demonstrates the performance of different values of $\lambda$.
-According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.
-1-step is better than 10-step of AT in pre-training.
-We also show that MIMIR outperforms original MAE (He et al.
-2022) and adv MAE with different PGD steps (to generate adversarial examples for training).
-MAE in Table 5 refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD.
-The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in loss.
-The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training.
-The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task.
-This more difficult task further improves the performance of downstream models (See also Table 7).
-We use the default learning rate (i.e., $45.0e-$ 4) of MAE, so there is a performance drop in experiments in Table 5 and 4 since AT prefers larger learning rates on CIFAR-10 as shown in Table 11 in the Appendix.
-Stability.
-Due to the high computation cost, we cannot report standard deviation for all experiments.
-To show that our method MIMIR has low variances, we train ViT-S on CIFAR-10 three times with different random seeds, running pre-training 400 epochs and fine-tuning 50 epochs.
-The natural accuracy is 86.07 $\pm$ 0.16 %.
-The adversarial accuracy under PGD20 ($l_{\infty},\epsilon=8/255$) is 47.24 $\pm$ 0.12 %.
-Table 6: Adversarial accuracy by adaptive attacks. The models are pre-trained 800 epochs by MIMIR and fine-tuned 100 epochs by PGD AT.
+Figure 5: Natural and adversarial accuracy of ViT-S adversarially fine-tuned for 5 or 10 epochs on CIFAR-10, Tiny-ImageNet, and ImageNet-1K.
+Data augmentation is not always harmful.
+Prior research [9, 20] has established that strong data augmentation techniques can adversely affect ViTs during adversarial training, as they may make training samples challenging to learn.
+However, we observe that strong data augmentation does not impair model performance when combined with longer fine-tuning periods.
+The strong data augmentation refers to the combination of Randaugment [28], CutMix [29], and MixUp [30].
+In this section, we evaluate two different solutions to ease this problem.
+First, we only use simple data augmentation for adversarial training, including random crop (or random resize crop for ImageNet-1K) and random horizontal flip (“weak aug”).
+Second, we use a 10-epoch warmup procedure for strong data augmentation.
+The warmup of Randaugment is implemented by progressively increasing the distortion magnitude from 1 to 9 (“warmup aug”).
+For CutMix and MixUp, we warm up by increasing the mixup probability from 0.5 to 1.0.
+As shown in Figure 6, “weak aug” provides the best accuracy.
+The “warmup aug” shows a slightly improved accuracy compared to fusing strong augmentation.
+Therefore, we provide a different result from [9] on the smaller dataset CIFAR-10, i.e., we show that weak augmentation is better than warmup augmentation.
+Even data augmentation with reduced amplitude is still difficult to learn at the beginning of adversarial training.
+Although strong augmentation is harmful to a normal training schedule, we show in Table X that CutMix [29], MixUp [30], and Randaugment [28] increase the accuracy of adversarial training when training with a longer schedule, e.g., 800 epochs of fine-tuning.
+We conjecture that combining data and strong augmentations is helpful but difficult for adversarial training to learn.
+Thus, more epochs are needed to learn meaningful representation.
+Loss and accuracy curves can be found in Appendix -E.
+TABLE X: Data augmentation with longer fine-tuning schedule.
+Arch Epoch Augmentation Natural PGD20
+ViT-B 800 Weak Augmentation 89.90 60.26
++ CutMix [29],MixUp [30] 91.01 60.62
++ Randaugment [28] 90.19 62.75
+TABLE XI: Adversarial accuracy by adaptive attacks. The models are pre-trained for 800 epochs by MIMIR and fine-tuned for 100 epochs by 1-step PGD AT.
 Dataset Model PGD20 PGD-MI100 PGD-fea100
@@
 ViT-B 55.41 55.36 73.51
-Adaptive Attacks.
-We evaluate MIMIR against adaptive adversaries following common practices (Tramer et al.
-2020).
+IV-D Adaptive Attacks
+We evaluate MIMIR against adaptive adversaries following common practices [81].
+Adaptive adversaries possess the capability to devise targeted attacks specifically tailored to exploit the mechanisms of MIMIR, particularly if they have prior knowledge of its architecture and defensive strategies.
+For example, the adversary may attack feature space [82, 83] since MIMIR trains the backbone to extract robust features.
+Here, the backbone refers to the ViT model without the classification layer, i.e., the encoder of MIMIR.
 We provide two adaptive attacks specifically designed against MIMIR.
@@
 where the value of $\lambda$ in MIMIR pre-training is available to adversaries.
-Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following (Sabour et al.
-2016).
+Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following [82].
 In particular, we attack the feature extractor from the backbones after the adversarial fine-tuning.
@@
 $\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{\text{mse}}$ $(f_{e}(x),f_{e}(x+\delta)).$ (22)
-Both PGD-MI and PGD-fea are optimized for 100 steps to make sure the attacking algorithm converges.
+Both PGD-MI and PGD-fea are optimized for 100 steps to ensure the attacking algorithm converges.
 The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$ for CIFAR-10 and Tiny-ImageNet, and $\epsilon=4/255$ for ImageNet-1K.
-Table 6 demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks, showing that MIMIR is robust against adaptive attacks on benchmark datasets.
-Table 7: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
-Architecture Pre-train CIFAR-10 Tiny ImageNet
+Table XI demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks.
+PGD-MI performs slightly better than the standard PGD attack, which means MI is exploitable information for perturbation crafting, but cannot significantly reduce the robustness.
+Furthermore, MIMIR-trained models are converged to a local optimal where the majority of predictions are constantly around the ground truth within the ball function of $\epsilon$.
+This also explains the resilience against both PGD and PGD-MI variants.
+Regarding PGD-fea, it aims to maximize feature-space divergence rather than the distance of output logits.
+However, MIMIR learned robust features that cannot be easily separated, so PGD-fea performs worse than PGD and PGD-MI.
+The collective results demonstrate that MIMIR’s IB framework induces robust learning dynamics that resist both output-space and feature-space attacks.
+Figure 6: Training loss and natural accuracy of ViT-S with three different data augmentations on CIFAR-10.
+Figure 7: The loss landscapes of MIMIR and MAE pre-trained models.
+TABLE XII: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
+Architecture Pre-train CIFAR-10 Tiny-ImageNet ImageNet-1K
 ViT-B MAE 96.79 73.38 82.92
@@
 MIMIR 95.95 71.14 78.69
-Figure 3: The loss landscapes of MIMIR pre-trained models.
-Visualization of the Loss Landscape.
-To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape (Li et al.
-2018) in Figure 3.
+IV-E Fine-tuning with Natural Images
+In Table XII, we show the results of MIMIR when fine-tuning with natural images.
+We compared the performance with MAE [37].
+We fine-tune for 50 epochs for CIFAR-10 and Tiny-ImageNet, and 100 epochs for ImageNet-1K.
+The results in Table XII are reported with 800 pre-training epochs.
+The base learning rate used in Table XII is 0.001.
+The fine-tuning batch size is 512 for CIFAR-10 and Tiny-ImageNet, and 1024 for ImageNet-1K.
+We use weak data augmentation (“weak aug”), which includes random crop and random horizontal flip.
+Not surprisingly, MIMIR outperforms MAE when fine-tuning with natural data.
+This is because MIMIR creates a harder learning task, which is helpful to learn more discriminative representation, as we discussed in Section III-B, Design Intuition.
+According to Tables XII and XIII, MIMIR consistently shows improved performance on natural data.
+Although the models in Table XIII show poor robustness due to fine-tuning on natural data, MIMIR pre-trained ones provide slightly better robustness.
+We want to clarify that poor robustness is expected when fine-tuning with natural data.
+First, it is known that standard training on natural data learns non-robust features [84], which hurts performance under adversarial attacks.
+Second, MIMIR pre-training is implemented using MSE loss plus an MI penalty between natural inputs and adversarial images.
+The adversarial perturbations and MI penalty help MIMIR create a more difficult and discriminative learning task to learn meaningful and robust features.
+This process does not include the classification layer of the final model.
+Therefore, MIMIR needs a fine-tuning process for superior performance on natural data and adversarial inputs.
+In other words, the superior performance of our experiments comes from the combination of MIMIR and the simple fine-tuning process.
+IV-F Visualization of the Loss Landscape
+To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape [85] in Figure 7.
 The loss landscape is the visualization of the loss function as parameters change.
 The basic idea is to plot the loss around the optimal parameters.
+Formally, we consider in the 2D case,
+$f_{l}(\alpha,\beta)=L(\theta^{*}+\alpha\theta_{1}+\beta\theta_{2}),$ (23)
+where $\theta_{1}$ and $\theta_{2}$ are two direction vectors, $\alpha$ and $\beta$ are two arguments of $f_{l}$.
+In practice, we use the parameters of trained models, i.e., $\theta^{*}$.
 The landscapes of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
-Fine-tuning with Natural Images.
-In Table 7, we show the results of MIMIR and the original MAE (He et al.
-2022) with the same hyperparameter.
-We fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K.
-Surprisingly, MIMIR outperforms MAE when fine-tuning with natural data.
-We leave further exploration of this finding to future work.
-5 Conclusions and Future Work
-This paper provides a novel theoretical MI analysis on ViT adversarial training, showing that MI between the adversarial example and its latent representation in ViT-based autoencoders, i.e., $I(x+\delta,z)$, should be constrained by utilizing the MI bounds.
-Based on this finding, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs.
-MIMIR uses adversarial examples as inputs and natural data as the reconstruction target.
-In this way, the information from the adversarial perturbations is decreased by the bottleneck, while the information from natural data is preserved while reconstructing the target.
-Our experimental results show that MIMIR substantially improves adversarial robustness compared to recent related works on various benchmark datasets.
-Appendix A Datasets
-We use three commonly used datasets to evaluate MIMIR: CIFAR-10 (Krizhevsky, Hinton et al.
-2009), Tiny-ImageNet (Le and Yang 2015), and ImageNet-1K (Deng et al.
-2009).
-CIFAR-10 (Krizhevsky, Hinton et al.
-2009) comprises 50,000 images with size $3\times 32\times 32$ in 10 classes.
-ImageNet-1K (Deng et al.
-2009) is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes.
-In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$.
-For completeness, we also include Tiny-ImageNet (Le and Yang 2015) as a medium size dataset between CIFAR-10 (Krizhevsky, Hinton et al.
-2009) and ImageNet-1K (Deng et al.
-2009).
-Tiny-ImageNet (Le and Yang 2015) contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
-Appendix B Decoder Hyperparameters
-We use transformer blocks but fewer layers as the backbone of the decoder.
-For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K.
-Table 8 shows the hyperparameters of decoder architectures.
-For different ViT architectures, we use the transformer blocks of the respective architectures to build the encoder.
-Table 8: Model architectures of the encoder and decoder.
-Model Layers Hidden size MLP ratio Heads
-ViT-T (encoder) 12 192 4 3
-decoder 2 128 4 16
-Appendix C Details of Training Hyperparameters
-In Tables 9 and 10, we provide the default hyperparameters used in our experiments.
-In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K.
-Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient.
-Table 9: Pre-training hyperparameters.
-Config Value
-optimizer AdamW
-base learning rate 1.5e-4
-weight decay 0.05
-optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.95$
-batch size 512(CIFAR-10, Tiny), 2,048 (ImageNet-1K)
-learning rate schedule cosine decay
-warmup epochs 40
-training epochs 800
-augmentation RandomResizedCrop, RandomHorizontalFlip
-Table 10: Fine-tuning hyperparameters.
-Config Value
-optimizer AdamW
-base learning rate 0.5e-2 (CIFAR-10), 1e-3 (ImageNet-1K, Tiny)
-weight decay 0.05
-optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
-layer-wise lr decay 0.65, 1.0
-batch size 128 (CIFAR-10), 256 (Tiny), 1,024 (ImageNet-1K)
-learning rate schedule cosine decay
-warmup epochs 10, 20
-training epochs 100, 300
-augmentation RandomResizedCrop, RandomHorizontalFlip
-drop path 0.1
-Table 11: Different learning rates. Fine-tuned for 50 epochs.
-Dataset Models LR Natural PGD10
-CIFAR-10 ViT-T 5.0e-4 76.30 47.60
-1.0e-3 80.69 49.56
-1.0e-2 85.62 48.78
-5.0e-2 85.12 50.30
-1.0e-1 84.51 50.40
-Hardware. Experiments are performed on a single machine with 4 RTX A6000 (48GB) and 4 RTX A5000 (24GB) GPUs, CUDA 12.0.
-Figure 4: Natural and adversarial accuracy of ViT-S and ViT-B that are fine-tuned with different numbers of pre-train epochs under a 20-step PGD attack.
-Appendix D Training Epochs Evaluation Study
-Pre-training epoch.
-Our experiments so far are based on 800-epoch pre-training.
-In Figure 4, we show the influence of different numbers of epochs.
-We use ViT-T and ViT-S as the models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet.
-We fine-tune each model 50 epochs after MIMIR pre-training.
-Both adversarial and natural accuracy are improved with a longer training schedule.
-On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S.
-On Tiny-ImageNet, the increase is more obvious.
-Therefore, a longer pre-training schedule obviously increases performance.
-Fine-tuning epoch.
-To show that the performance of the final model is largely attributed to MIMIR pre-training, we do a very short fine-tuning for pre-trained models.
-This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training.
-Therefore, we evaluate the pre-training performance.
-In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning.
-Figure 5: Natural and adversarial accuracy of ViT-S adversarially fine-tuned for 5 or 10 epochs on CIFAR-10, Tiny-ImageNet, and ImageNet-1K.
-Table 12: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
-CIFAR-10 Tiny-ImageNet ImageNet-1K
-Architecture #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
-ViT-S 21.34 PGD10 AT 149.3 2.54$\times$4 306.0 3.99$\times$4 2251.7 12.5$\times$4
-FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
-MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
-MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
-ViT-B 85.27 PGD10 AT 361.2 5.39$\times$4 1022.2 8.30$\times$4 5416.7 22.1$\times$4
-FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
-MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
-MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
-ConViT-S 27.05 PGD10 AT 442.5 6.64$\times$4 897.0 12.19$\times$4 6626.5 32.5$\times$4
-FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
-MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
-MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
-Appendix E Efficiency
-Table 12 provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning.
-The time consumption and memory usage are evaluated on 4 A6000 GPUs.
-Compared to 10-step PGD and FastAT, MIMIR shows much less time consumption.
-At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$.
-Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time.
-We also consider using a 10-step PGD at pre-training.
-Table 5 shows that using a one-step algorithm with larger $\alpha=10$ provides better performance.
-In addition, pre-training with MIMIR also costs less memory.
-Note in Table 12 that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning.
-In Table 12, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training (Wong, Rice, and Kolter 2020).
-Therefore, MIMIR is more promising regarding training costs when conducting adversarial training.
-Appendix F Mutual Information and HSIC
-MI measures the mutual dependence between two random variables, $X$ and $Y$. It can be decomposed as:
-$I(X,Y)$ $=H(X)-H(X|Y),$ (20)
-$=H(Y)-H(Y|X),$
-$=H(X)+H(Y)-H(X,Y),$
-where $H(X)$ and $H(Y)$ are the information entropies, $H(X|Y)$ and $H(Y|X)$ are the conditional entropies, and $H(X,Y)$ is the joint entropy of $X$ and $Y$.
-Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$.
-To address this issue, the deterministic information bottleneck (DIB) (Yu, Yu, and Príncipe 2021) uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ (Sanchez Giraldo, Rao, and Principe 2015; Yu et al.
-2019), which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
-Specifically, given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$ (in our setup, $N$ refers to the mini-batch size), we can obtain two Gram (or kernel) matrices $K_{x}$ and $K_{y}$, for variables $X$ and $Y$, respectively, with $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, in which $\kappa_{x}$ and $\kappa_{y}$ are corresponding kernel functions.
-The information entropy of $X$ can be expressed as:
-$H_{\alpha}(X)=\frac{1}{1-\alpha}\log_{2}\left(\text{tr}(\tilde{K_{x}}^{\alpha})\right)=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{N}\lambda_{i}(\tilde{K_{x}})^{\alpha}\right),$ (21)
-where $\tilde{K}$ is the normalized version of $K$, i.e., $\tilde{K}=K/{\text{tr}(K)}$, and $\lambda_{i}(\tilde{K})$ denotes the $i$-th eigenvalue of $\tilde{K}$.
-Further, the joint entropy for $X$ and $Y$ can be expressed as:
-$H_{\alpha}(X,Y)=H_{\alpha}\left(\frac{K_{x}\circ K_{y}}{\text{tr}(K_{x}\circ K_{y})}\right),$ (22)
-where $K_{x}\circ K_{y}$ denotes the Hadamard product between the matrices $K_{x}$ and $K_{y}$.
-Given Eqs. (21) and (22), the matrix-based Rényi’s $\alpha$-order mutual information $I_{\alpha}(X;Y)$ in analogy of Shannon’s MI is given by:
-$I_{\alpha}(X;Y)=H_{\alpha}(X)+H_{\alpha}(Y)-H_{\alpha}(X,Y).$ (23)
-Throughout this paper, we use the radial basis function (RBF) kernel $\kappa(x_{i},x_{j})=\exp(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}})$ with kernel width $\sigma$ to obtain the Gram matrices.
-The Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al.
-2005) is also a kernel-based dependence measure and is usually used as a surrogate of MI.
-Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
-$\begin{split}&\text{HSIC}_{P_{X,Y}}(X,Y)\\ &=||C_{XY}||^{2}\\ &=\mathbb{E}_{xyx^{\prime}y^{\prime}}[\kappa_{x}(x,x^{\prime})\kappa_{y^{\prime}}(y,y^{\prime})]\\ &+\mathbb{E}_{xx^{\prime}}[\kappa_{x}(x,x^{\prime})]E_{yy^{\prime}}[\kappa_{y}(y,y^{\prime})]\\ &-2\mathbb{E}_{xy}[\mathbb{E}_{x^{\prime}}[\kappa_{x}(x,x^{\prime})]\mathbb{E}_{y^{\prime}}[\kappa_{y}(y,y^{\prime})]],\end{split}$ (24)
-where $\kappa_{x}$ and $\kappa_{y}$ are kernel functions, $\mathbb{E}$ is the expectation, $x^{\prime}$ and $y^{\prime}$ are independent copies of $x$ and $y$, respectively.
-Given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$, the empirical estimator of HSIC is given by:
-$\widehat{\text{HSIC}}_{P_{X,Y}}(X,Y)=\frac{1}{N^{2}}\text{tr}\left(K_{x}HH_{y}H\right),$ (25)
-in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathds{1}\mathds{1}^{T}$ is the centering matrix.
+TABLE XIII: Natural and adversarial accuracy of ViT-S that is fine-tuned for 5 or 50 epochs with natural images.
+Fine-tune Dataset Pre-train Natural PGD
+5 epochs CIFAR-10 MIMIR 93.22 0.55
+MAE 93.16 0.01
+Tiny-ImageNet MIMIR 63.65 0.00
+MAE 62.21 0.00
+ImageNet-1K MIMIR 72.45 0.20
+MAE 69.47 0.02
+50 epochs CIFAR-10 MIMIR 95.95 0.28
+MAE 95.95 0.29
+Tiny-ImageNet MIMIR 71.14 0.00
+MAE 70.00 0.00
+ImageNet-1K MIMIR 78.69 0.18
+MAE 77.45 0.05
+V Related Work
+V-A Vision Transformer
+The transformers [86] were first proposed in natural language processing (NLP).
+With the mechanism of global self-attention, transformers can effectively capture the non-local relationships among all text tokens [87, 88, 89].
+A substantial effort is made to apply the transformer and self-attention mechanism in computer vision [3, 90, 5].
+The pioneering work, ViT [3], demonstrated that the pure transformer architecture could achieve competitive performance on various tasks.
+ViT also reveals that transformers lack inductive biases [3].
+For example, locality, two-dimensional neighborhood structure, and translation equivariance are inherent to CNNs but not applicable to ViTs [3].
+Due to this shortcoming, ViTs usually require large-scale training to get competitive performance, such as pre-training on ImageNet-21K [36] and JFT-300M [91].
+To alleviate the ViT need for large datasets, DeiT [92] introduced a teacher-student strategy to distill knowledge from a teacher CNN for a student ViT.
+In the MIM field, MAE [37] uses a masked autoencoder with a lightweight decoder as a visual representation learner.
+Its learning objective is to reconstruct the original image by the decoder while using masked images as input to the autoencoder.
+The advantage is that MAE can randomly discard 75% image patches when pre-training under ImageNet-1K [36], which means more efficient training.
+V-B Adversarial Attacks on ViTs
+The concept of adversarial attacks first appeared in [43], which proposed a formal framework and algorithms against the adversarial spam detection domain.
+Then, the adversarial attacks were popularized by Biggio et al.
+[14] and Szegedy et al.
+[15] in image classification.
+The generation of adversarial examples depends on the model’s gradient or estimated gradient in a black box situation [93].
+Therefore, adversarial attacks can be easily applied to transformers by using the gradient of attention blocks concerning inputs.
+This also raises the question of whether transformers are more robust than CNNs.
+Benz et al.
+[94] found that CNNs are less robust than ViTs due to their shift-invariant property.
+Bhojanapalli et al.
+[95] found that ResNet models are more robust than transformers at the same model size under FGSM attack, but under PGD [21] attack, transformer models show better robustness.
+As ViTs process the input image as a sequence of patches, Gu et al.
+[13] found that ViTs are more robust than CNNs to naturally corrupted patches because the attention mechanism helps ignore naturally corrupted image patches.
+The later work [9] revealed that CNNs could be as robust as ViTs against adversarial attacks if CNNs are trained with proper hyperparameters.
+V-C Adversarial Defense
+PGD [21] adversarial training is considered one of the most effective defenses for CNNs and can withstand adaptively designed attacks [96].
+However, PGD AT is harmful to the accuracy of clean data [21, 26, 2].
+Generalist [1] solves this problem by formulating different training strategies for robust and natural generalization separately.
+DBAT [2] solves the decrease in natural accuracy by adding dummy classes [97] to the classification space.
+Due to the difference between CNNs and ViTs, there have been some recent efforts to explore new adversarial training approaches for ViTs [11, 20, 12].
+Mo et al.
+[11] presented a new adversarial training strategy based on the following observations: 1) pre-training with natural data can provide better robustness after adversarial fine-tuning, 2) gradient clipping is necessary for adversarial training, and 3) using SGD as the optimizer is better than Adam.
+Debenedetti et al.
+[20] also presented an improved training strategy for ViTs by evaluating different combinations of data augmentation policies.
+As adversarial training is time-consuming, AGAT [12] leverages the attention score while training to discard non-critical image patches after every layer.
+Unlike previous works, we provide a different training paradigm by using MIM for adversarial pre-training.
+Our method is efficient as we discard 75% image patches while pre-training.
+Our method is effective as we eliminate the information of adversarial perturbations from two information sources of natural and adversarial inputs.
+We also provide theoretical proof that the information of adversarial perturbations is eliminated.
+V-D Self-Supervised Adversarial Pre-Training
+Self-supervised learning [98, 99, 37, 100] refers to extracting meaningful representation from unlabeled data, which can be used for downstream recognition tasks.
+Self-supervised methods are beneficial for out-of-distribution detection on difficult, near-distribution outliers [27], which leads to using self-supervised training to improve adversarial robustness [101, 27, 102, 103, 104, 105].
+The basic idea is to build a min-max learning object similar to traditional adversarial training.
+For example, Jiang et al.
+[102] considered using two adversarial samples or combining one adversarial sample and one natural sample to learn a consistent representation in contrastive learning.
+In more recent work, You et al.
+[106] proposed NIM De3 to denoise adversarial perturbations.
+However, the motivation of these works relies on complex self-supervised pre-training technologies, making it more difficult to understand the inner mechanisms or provide theoretical results.
+MIMIR not only provides better performance but also provides intuitive insights with theoretical motivation.
+VI Discussion and Limitations
+Following the principle of IB, we can intuitively consider a bottleneck between the encoder and decoder.
+As the reconstruction output is constrained by natural data $x$, the bottleneck will filter out information from adversarial perturbations $\delta$.
+We provide a theoretical guarantee of this bottleneck.
+In Eq.
+(8), we embed this bottleneck as a learning object to further improve the performance, which also confirms the correctness of our theoretical guarantee.
+With the two information sources of $x$ and $\delta$, the model is trained to learn the robust features from $x$ and forget the information of $\delta$ under the constraint of the reconstruction target.
+While MIMIR shows better performance, there are still limitations.
+MIMIR is a pre-training method.
+Adversarial fine-tuning is necessary to build the final robust model.
+Thus, the shortcomings of traditional adversarial training cannot be completely avoided.
+In our experiments, we utilize the simple PGD algorithm for fine-tuning, but one can further improve MIMIR pre-trained models with more advanced approaches.
+In addition, MIMIR follows the design of MAE, and we also utilize the characteristic that ViTs can process variable-length inputs.
+Therefore, the current MIMIR cannot directly handle CNNs.
+While it is not trivial, we apply MIMIR to the latest CNN architecture by sparse convolution from SparK [40].
+However, sparse convolution is not as efficient as dropping patch embeddings.
+We leave these limitations to future work.
+VII Conclusions
+This paper provides a novel theoretical analysis of AT for ViTs through the lens of IB.
+We found that constraining the MI between adversarial perturbations and their latent representations in ViT-based autoencoders, as governed by derived MI bounds, is critical for enhancing model robustness.
+Building upon this theoretical foundation, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs.
+MIMIR operates by processing adversarial examples as inputs while reconstructing their natural data as targets.
+This approach leverages the inherent information bottleneck in autoencoder architectures to achieve two key objectives: (1) progressively eliminating perturbation-related information while (2) preserving the essential features of the original data distribution.
+Our extensive experimental evaluation demonstrates that MIMIR significantly outperforms existing adversarial training methods across multiple benchmark datasets, achieving SOTA results on ImageNet-1K.
+In addition, MIMIR is robust against unforeseen attacks and common corrupted data and can resist adaptive attacks.
+Acknowledgment
+The authors thank the anonymous reviewers for their constructive comments that helped improve this paper and the artifact.
+This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no.
+EINF-10853.
+This work was partially supported by the Horizon Europe programme under the project SHASAI (No.
+101225866).